### v2

# Strategic Analysis of Fair Rank-Minimizing Mechanisms with Agent Refusal Option

Abstract
This study examines strategic issues in fair rank-minimizing mechanisms, which choose an assignment that minimizes the average rank of object types to which agents are assigned and satisfy a fairness property called equal treatment of equals. As one of these fair mechanisms, the uniform rank-minimizing mechanism is considered. We focus on the case where agents can refuse their assignment and obtain the outside option instead. Without the refusal option, truth-telling is not strategically dominated by any strategies if a fair rank-minimizing mechanism is used. However, if agents have the option and the uniform rank-minimizing mechanism is used, then a strategy called an outside option demotion strategy strategically dominates truth-telling. Moreover, we show that adopting this strategy may lead to inefficient assignments. To counter this, we propose a modification of the uniform rank-minimizing mechanism, though it may lead agents to strategically reduce the number of acceptable types.
JEL classification: C78; D47
Keywords: Rank-minimizing mechanism; Strategic manipulation; Random assignment; Outside option; Refusal option
1 Introduction
In this study, we discuss the strategic issues of fair rank-minimizing allocation mechanisms mainly when agents have the option to refuse their assignment determined by one of the mechanisms and obtain their outside option instead. There are several agents and types of object, including a null object type, which corresponds to the outside option. In the mechanism, each agent reveals one preference ranking, which may or may not be their true preference ranking. In rank-minimizing mechanisms, assignments are determined as to minimize the total (or average) revealed rank of the objects to which the agents are assigned.
The rank distributions of the assignments are seriously considered by many real world matching authorities. For example, Featherstone (2020) mentions that Teach for America considers the rank distribution when selecting its assignment. Therefore, the rank-minimizing mechanisms are focused on by several recent studies such as Featherstone (2020), Nikzad (2022), Ortega and Klein (2023) and Troyan (2024). Some of them highlight several advantages of the mechanism with regard to the efficiency of assignment. Featherstone (2020) shows that any rank-minimizing mechanism satisfies another property of efficiency called ordinal efficiency, first introduced by Bogomolnaia and Moulin (2001). Moreover, Ortega and Klein (2023) compare the result of a rank-minimizing mechanism and those of the DA mechanism and the TTC mechanism, and show that the result of a rank-minimizing mechanism improves upon those of others in several terms of efficiency.
In this study, we especially consider the rank-minimizing mechanisms that satisfy a fairness property called equal treatment of equals, which requires that if two agents reveal the same preference ranking, then they are assigned to each object type with the same probability. This fairness property is discussed by Bogomolnaia and Moulin (2001) and many other previous studies on various allocation problems (Thomson 2011). We consider a slightly stronger version of this property. We call a rank-minimizing mechanism satisfying equal treatment of equals a fair rank-minimizing mechanism. To consider these mechanisms, we adopt the probabilistic assignment framework. Among the mechanisms, we first consider a representative one called the uniform rank-minimizing mechanism. Since Nikzad (2022), Ortega and Klein (2023) and Troyan (2024) also mainly consider this mechanism, at least in the literature, this is the most common rank-minimizing mechanism.
We discuss the strategic manipulation incentives of agents when the assignment is determined by a fair rank-minimizing mechanism. Specifically, we consider the case where agents can refuse their assignment, and if they refuse it, then the outside option is assigned to them. First, as a benchmark case, we discuss the situation where agents cannot refuse their assignment. This case is discussed by many previous studies, such as those mentioned above. Featherstone (2020) shows that any rank-minimizing mechanism is not strategy-proof. On the other hand, Troyan (2024) shows that the uniform rank-minimizing mechanism is not obviously manipulable, which is a strategic property introduced by Troyan and Morrill (2020). Moreover, Nikzad (2022) shows that the uniform rank-minimizing mechanism is Bayesian incentive compatible in uniform markets, which he primarily discusses. Related to but logically independent to their results, we show that, in this case, truth-telling is not strategically dominated by any untrue preference under any fair rank-minimizing mechanism. These results imply that in this case, strategic manipulation may not be a critical problem of rank-minimizing mechanisms.
Next, we consider the case where agents can refuse their assignment and instead obtain the outside option. Featherstone (2020) firstly considers this situation, but focuses only on the case where each agent has limited information on both the preferences of other agents and the capacities of all object types except the outside option. In contrast, we assume that each agent has complete information on the capacities of all object types. For example, in a school choice market, information about all schools, including their capacities, is assumed to be disclosed to all students. Additionally, we do not need any assumptions about the information structure of agents regarding the preferences and rationality of other agents because we focus only on strategic dominance relations.
We consider an outside option demotion strategy, which involves revealing a ranking where the rank of the outside option is placed at the bottom, without changing the order of the truly acceptable object types. We show that, if the uniform rank-minimizing mechanism is used and agents have the refusal option, then any outside option demotion strategy weakly strategically dominates truth-telling. Moreover, with some additional conditions, some outside option demotion strategy (strictly) strategically dominates truth-telling.
The reason is as follows: first, for some preference profiles of other agents, an agent succeeds in avoiding the outside option and getting a truly acceptable object type because assigning the outside option to them significantly increases the sum of ranks. Of course, due to this manipulation, the agent faces the risk of being assigned an unacceptable object type. However, in the case where the agents have the refusal option, the agent can refuse any of such types and obtain the outside option instead. Moreover, we show that if an agent adopts an outside option demotion strategy and the uniform rank-minimizing mechanism is used, an inefficient assignment may be realized. This is because, in that case, an unacceptable type may be assigned to the agent who adopts the strategy, and then the agent refuses it, even though the type is desired by other agents. Therefore, in this case, strategic manipulation is a critical problem for rank-minimizing mechanisms.
Featherstone (2020) shows that revealing the full extension, which is one of the outside option demotion strategies, is a weakly better strategy than any other strategies in the case where the agents have little information and the refusal option. The implication of Featherstone (2020) is that rank-efficient mechanisms have strategic issues if the agents have little information and the outside option is somewhat valued by some agents. On the other hand, by our result, this implication is robust for the uniform rank-minimizing mechanism even without the low information assumption. Moreover, we show that revealing the full extension may be strategically dominated by another outside option demotion strategy because, contrary to our model, Featherstone (2020) assumes that the agents have no information on the capacities of object types other than their outside option.
Due to the negative result on the uniform rank-minimizing mechanism introduced above, we modify it so that truth-telling is not strategically dominated by any other strategy. We call this mechanism the modified uniform rank-minimizing mechanism, which is shown to be fair. That is, to ensure that any outside option demotion strategy can fail, in the mechanism, increasing acceptable object types can make the agent less likely to be assigned their acceptable types. However, using the modified uniform rank-minimizing mechanism may create another strategic issue: an agent may have an incentive to reduce the number of truly acceptable object types.
2 Model
We consider the model discussed by Kojima and Manea (2010) and Featherstone (2020). Let $\mathcal{A}$ and $\mathcal{O}$ be finite sets of agents and object types respectively. Each agent must be allocated one copy of an object whose type is an element of a set denoted by $\mathcal{O}$. Let $q_{o}\geq 1$ be the capacity of an object type $o\in\mathcal{O}$. We assume that there is a null object type denoted by $\varnothing\in\mathcal{O}$ satisfying $q_{\varnothing}\geq\left|\mathcal{A}\right|$. That is, the null object type can be assigned to all agents simultaneously. The null object type $\varnothing\in\mathcal{O}$ can be interpreted as the (best) outside option of an agent. In Section 4, we consider the situation where the null object type has the other characteristic. To avoid trivial cases, we assume $q_{o}<\left|\mathcal{A}\right|$ for all $o\in\mathcal{O\setminus}\left\{\varnothing\right\}$.
Let $\mathcal{P}$ be the set of all possible strict rankings (linear orders) of the elements in $\mathcal{O}$. Let $p_{a}=\left(p_{a}^{1},\cdots,p_{a}^{\left|\mathcal{O}\right|}\right)\in\mathcal{P}$ be a preference order revealed by $a$, where $p_{a}^{k}=o$ means that type $o$ object is ranked $k$th best one under $p_{a}$. For notational convenience, we let $R:\mathcal{O\times P}\rightarrow\left\{1,\cdots,\left|\mathcal{O}\right|\right\}$ be a rank function; that is, $p_{a}^{k}=o$ if and only if $R\left(o,p_{a}\right)=k$. Moreover, if $R\left(o,p_{a}\right)<R\left(\varnothing,p_{a}\right)$ (resp. $R\left(o,p_{a}\right)>R\left(\varnothing,p_{a}\right)$), then $o$ is said to be acceptable (resp. unacceptable) for $p_{a}$. Thus, note that $R\left(\varnothing,p_{a}\right)-1$ is the number of acceptable object types for $p_{a}$. Let $r_{a}\in\mathcal{P}$ be the true preference order of agent $a$. Especially, if $R\left(o,r_{a}\right)<R\left(\varnothing,r_{a}\right)$, then $o$ is said to be truly acceptable for $a$.
An assignment is represented as lotteries over types. Let $x_{ao}$ be the probability that agent $a$ assigns (a copy of) type $o$ object. Moreover, let $x$ be a matrix whose $\left(a,o\right)$ entry is equal to $x_{ao}$. We say that a matrix $x$ is an assignment if
$x_{ao}\in\left[0,1\right]\text{, }\mathop{\displaystyle\sum}\limits_{o^{\prime}\in\mathcal{O}}x_{ao^{\prime}}=1\text{,\ and }\mathop{\displaystyle\sum}\limits_{a^{\prime}\in\mathcal{A}}x_{a^{\prime}o}\leq q_{o},$
for all $a\in\mathcal{A}$ and all $o\in\mathcal{O}$. Let $\mathcal{X}$ be the set of all possible assignments. Specifically, let $y\in\mathcal{X}$ be a deterministic assignment if $y_{ao}$ is either 00 or $1$ for all $\left(a,o\right)\in\mathcal{A}\times\mathcal{O}$. Let $\mathcal{Y}\subseteq\mathcal{X}$ be a set of all possible deterministic assignments.
The following result is due to Kojima and Manea (2010), which is a generalization of the Birkhoff–von Neumann theorem.
Remark 1
Every assignment can be written as a convex combination of deterministic assignments.
We consider the preference of agent $a$ with $p_{a}$ with regard to two assignments $x$ and $x^{\prime}$. If
$\mathop{\displaystyle\sum}\limits_{o^{\prime}:\text{ }R\left(o^{\prime},p_{a}\right)\leq k}x_{ao^{\prime}}\geq\mathop{\displaystyle\sum}\limits_{o^{\prime}:\text{ }R\left(o^{\prime},p_{a}\right)\leq k}x_{ao^{\prime}}^{\prime}$ (1)
for all $k=1,2,\cdots,\left|\mathcal{O}\right|$, then agent $a$ with $p_{a}$ is said to weakly prefer (in the sense of first-order stochastic dominance) $x$ to $x^{\prime}$. Moreover, $a$ is said to strictly prefers $x$ to $x^{\prime}$ if $a$ weakly prefers $x$ to $x^{\prime}$ and (1) is satisfied with inequality for some $k=1,2,\cdots,\left|\mathcal{O}\right|$.
Assignment $x$ weakly ordinally dominates assignment $x^{\prime}$ for $p$ if each agent $a$ with $p_{a}$ weakly prefers $x$ to $x^{\prime}$. If $x$ weakly ordinally dominates assignment $x^{\prime}$ for $p$ and (1) is satisfied with inequality for at least one agent–rank pair, then $x$ ordinally dominates assignment $x^{\prime}$ for $p$. An assignment that is not ordinally dominated by any other assignments is said to be ordinally efficient for $p$.
Assignment $x$ is wasteful for $p$ if there are $o,o^{\prime}\in\mathcal{O}$ and $a\in\mathcal{A}$ such that $\mathop{\textstyle\sum}\nolimits_{a^{\prime}\in\mathcal{A}}x_{a^{\prime}o}<q_{o}$, $x_{ao^{\prime}}>0$ and $R\left(o,p_{a}\right)<R\left(o^{\prime},p_{a}\right)$. Any wasteful assignment for $p$ must not be ordinally efficient. [^1] Moreover, a wasteful assignment for $p$ must not be efficient in a weaker efficiency concept such as ex-post efficiency, which is discussed several previous studies such as Bogomolnaia and Moulin (2001). Thus, not being wasteful is a very weak requirement for efficiency.
For an assignment $x\in\mathcal{X}$ and a preference profile $p=\left(p_{a}\right)_{a\in\mathcal{A}}\in\mathcal{P}^{\left|\mathcal{A}\right|}$, let $RV\left(x,p\right)$ be such that
$RV\left(x,p\right)=\mathop{\displaystyle\sum}\limits_{a\in\mathcal{A}}\mathop{\displaystyle\sum}\limits_{o\in\mathcal{O}}R\left(o,p_{a}\right)\times x_{ao},$
which is the sum of expected rank. An assignment $x^{\ast}\in\mathcal{X}$ is rank efficient for $p$ if $RV\left(x^{\ast},p\right)\leq RV\left(x,p\right)$ for all $x\in\mathcal{X}$. Let $\mathcal{X}^{\ast}\left(p\right)$ be the set of rank efficient assignments for $p$.
The following result is shown by Featherstone (2020).
Remark 2
Any rank efficient assignment for $p$ is ordinally efficient for $p$.
In the subsequent section, we focus on the mechanisms whose input is $p\in\mathcal{P}^{\left|\mathcal{A}\right|}$ and output is rank efficient for any $p\in\mathcal{P}^{\left|\mathcal{A}\right|}$. Before that we introduce some technical results.
Lemma 1
Fix any $x^{\ast}\in\mathcal{X\setminus Y}$. Let $y^{1},\cdots,y^{I}\in\mathcal{Y}$ with $I\geq 2$ be such that
$x^{\ast}=\mathop{\displaystyle\sum}\limits_{i=1}^{I}\sigma^{i}y^{i}\in\mathcal{X}\text{ where }\mathop{\displaystyle\sum}\limits_{i=1}^{I}\sigma^{i}=1$
and $\sigma^{i}\in\left[0,1\right]$ for all $i=1,\cdots,I$. If $x^{\ast}$ is rank efficient, then $y^{i}$ is also rank efficient for all $i=1,\cdots,I$.
We provide the proofs of our results in the Appendix.
By Lemma 1, we immediately have the following result.
Corollary 1
There must exist at least one deterministic rank efficient assignment.
Nikzad (2022) and Troyan (2024) consider a deterministic rank efficient assignment; that is, they focus only on a deterministic assignment $y^{\ast}\in\mathcal{Y}$ such that $RV\left(y^{\ast},p\right)\leq RV\left(y,p\right)$ for all $y\in\mathcal{Y}$. However, by Corollary 1, such $y^{\ast}$ must also be rank efficient in our definition. Moreover, this result allows us to show that the mechanism introduced later is well-defined.
Finally, we introduce another technical result that is used to show our main results introduced later.
Lemma 2
Let $x\in\mathcal{X}$ be an assignment that is not wasteful for $p$ and $k$ be the smallest integer satisfying
$\mathop{\displaystyle\sum}\limits_{\bar{o}:R(\bar{o},p_{a})<k}q_{\bar{o}}>\left|\mathcal{A}\right|.$ (2)
Then, $x_{ao}=0$ for all $o$ such that $R(o,p_{a})\geq k$.
The condition (2) means that there are sufficient number of copies of object types that ranked higher than $k$ in the preference ranking of agent $a$. Then, if the assignment is not wasteful, then $a$ must be assigned to an object type that ranked higher than $k$. Typically, in that case, the object types that ranked lower than $\varnothing$ in the preference ranking of agent $a$ are never assigned to $a$, because $\varnothing$ is the object with a sufficient number of copies.
3 Rank-minimizing Mechanism
In this section, we consider the following game. First, all agents simultaneously reveal a preference order in $\mathcal{P}$, which may or may not be their true preference order $r_{a}\in\mathcal{P}$. Let $p_{a}\in\mathcal{P}$ be the preference order revealed by $a\in\mathcal{A}$ and $p=\left(p_{a}\right)_{a\in\mathcal{A}}\in\mathcal{P}^{\left|\mathcal{A}\right|}$. Second, based on $p$, an assignment is determined via a mechanism. We assume that all agents know the capacities of all object types $\left(q_{o}\right)_{o\in\mathcal{O}}$ and the mechanism used to determine the assignment, when they reveal a preference order. Note that in this section, the null object type is just that with sufficient number of copies.
A mechanism used in the second stage is represented by $f:\mathcal{P}^{\left|\mathcal{A}\right|}\rightarrow\mathcal{X}$. First, we consider an efficiency property on mechanisms. We call $f$ a rank-minimizing mechanism if $f(p)$ is rank efficient for all $p$.
Second, we introduce a fairness property. First, we introduce a well-known one. A mechanism $f$ is said to satisfy weak equal treatment of equals (hereafter weak ETE) if $p\in\mathcal{P}^{\left|\mathcal{A}\right|}$ satisfying $p_{a}=p_{a^{\prime}}$ implies $\left(f\left(p\right)\right)_{ao}=\left(f\left(p\right)\right)_{a^{\prime}o}$ for all $o\in\mathcal{O}$. This requires that two agents revealing the same preference must have the same assignment in the ex-ante sense. This property is called just “equal treatment of equals” by several previous studies such as Bogomolnaia and Moulin (2001) and Troyan (2024). In this study, we mainly use a slightly stronger one. A mechanism $f$ is said to satisfy equal treatment of equals (hereafter ETE) if $p\in\mathcal{P}^{\left|\mathcal{A}\right|}$ satisfying $p_{a}^{k^{\prime}}=p_{a^{\prime}}^{k^{\prime}}$ for all $11,\cdots,k-$ 1 implies $\left(f\left(p\right)\right)_{ao}=\left(f\left(p\right)\right)_{a^{\prime}o}$ for all $o\in\mathcal{O}$, where $k$ is the smallest integer satisfying
$\mathop{\displaystyle\sum}\limits_{\bar{o}:R(\bar{o},p_{a})<k}q_{\bar{o}}\geq\left|\mathcal{A}\right|\text{.}$
By Lemma 2, as long as the assignment is not wasteful, any type that is ranked $k$th or lower in the preference ranking revealed by $a$ will not be assigned to $a$, because there are sufficient copies of types ranked higher than $k$th. Thus, if two individuals reveal rankings that are identical for positions above the $k$th rank, those rankings are considered essentially the same, and they should receive the same assignment in an ex-ante sense.
We focus on the efficient and fair mechanisms; that is, we call $f$ a fair rank-minimizing mechanism if $f(p)$ is rank efficient for all $p$ and satisfies ETE. The following mechanism is one of them.
Let $\mathcal{Y}^{\ast}\left(p\right)\subseteq\mathcal{Y}$ be the set of deterministic rank efficient assignments; that is, $y^{\ast}\in\mathcal{Y}^{\ast}\left(p\right),$ $RV\left(y^{\ast},p\right)\leq RV\left(x,p\right)$ for any $x\in\mathcal{X}$. Note that by Lemma 1, $\mathcal{Y}^{\ast}\left(p\right)\neq\emptyset$. Let
$f^{U}\left(p\right)=\mathop{\displaystyle\sum}\limits_{y^{\ast}\in\mathcal{Y}^{\ast}\left(p\right)}\frac{y^{\ast}}{\left|\mathcal{Y}^{\ast}\left(p\right)\right|}\text{ for all }p\in\mathcal{P}^{\left|\mathcal{A}\right|}$
be the uniform rank-minimizing mechanism. By (9),
$RV\left(f^{U}\left(p\right),p\right)=\mathop{\displaystyle\sum}\limits_{y^{\ast}\in\mathcal{Y}^{\ast}\left(p\right)}\frac{1}{\left|\mathcal{Y}^{\ast}\left(p\right)\right|}RV\left(y^{\ast},p\right).$
Thus, $RV\left(f^{U}\left(p\right),p\right)\leq RV\left(x,p\right)$ for any $x\in\mathcal{X}$; that is, $f^{U}\left(p\right)$ must be rank efficient for $p$. Trivially, $f^{U}$ satisfies ETE and thus is a fair rank-minimizing mechanism. Since Nikzad (2022), Ortega and Klein (2023) and Troyan (2024) also discuss the uniform rank-minimizing mechanism, this is one of the most common rank-minimizing mechanisms at least in the literature.
We have the following result.
Proposition 1
The uniform rank-minimizing mechanism satisfies ETE.
We discuss strategic problems of fair rank-minimizing mechanisms. We call the strategy of revealing their true preference truth-telling. Featherstone (2020) shows that any rank-minimizing mechanism does not satisfy strategy-proofness. That is, for any rank-minimizing mechanism $f$, there are $r_{a}\in\mathcal{P}$ and $p\in\mathcal{P}^{\left|\mathcal{A}\right|}$ such that $a$ with $r_{a}$ prefers $f\left(p\right)$ to $f\left(r_{a},p_{-a}\right)$. Therefore, in this paper, we consider a weaker strategic property.
We discuss whether truth-telling ($r_{a}$) can be strategically dominated by an untrue preference $p_{a}\neq r_{a}$. Let $k$ be the smallest integer such that $p_{ak}\neq r_{ak}$. By Lemma 2, if (2) is satisfied for that $k$, then $p_{a}$ and $r_{a}$ are essentially equivalent under a rank-minimizing mechanism. Thus, we focus only on the case where (2) is not satisfied for $k$ defined above; that is,
$\mathop{\displaystyle\sum}\limits_{\bar{o}:R(\bar{o},p_{a})<k}q_{\bar{o}}\left(=\mathop{\displaystyle\sum}\limits_{\bar{o}:R(\bar{o},r_{a})<k}q_{\bar{o}}\right)<\left|\mathcal{A}\right|$ (3)
is satisfied for that $k$. In other word, we consider the case where the number of object types that are more preferable than the $k$th best one in $p_{a}$ is insufficient to distribute to all agents. In this case, for some $p_{-a}\in\mathcal{P}^{\left|\mathcal{A}\right|-1}$ and some fair rank-minimizing mechanism, $a$ is assigned to an object type that is less preferable or equal to the $k$th best one.
Lemma 3
Suppose that $f$ is a fair rank-minimizing mechanism. Let $p_{a}\neq r_{a}$, and $k$ be the smallest integer such that $p_{a}^{k}\neq r_{a}^{k}$. If (3) is satisfied for that $k$, then there is $p_{-a}\in\mathcal{P}^{\left|\mathcal{A}\right|-1}$ such that agent $a$ (with $r_{a}$) strictly prefers $f\left(r_{a},p_{-a}\right)$ to $f\left(p_{a},p_{-a}\right)$.
Next, we have the following result.
Proposition 2
Suppose that a fair rank-minimizing mechanism is used. For any $r_{a}\in\mathcal{P}$, there is no $p_{a}\in\mathcal{P}$ such that revealing $p_{a}$ strategically dominates truth-telling.
This result implies that any fair rank-minimizing mechanism satisfies a weak strategic characteristic; that is, for each untrue preference of an agent $a$, there must exist a profile of preferences of agents other than $a$, such that $a$ is worse off by revealing it instead of $r_{a}$. Thus, if a fair rank-minimizing mechanism is used, any strategic manipulation by an agent carries some level of risk. This characteristic is quite weak, but in the subsequent section, we show that this is not satisfied when the agents can refuse their assignment and instead obtain the null object type.
Note that Proposition 2 holds for any fair rank-minimizing mechanisms but cannot be generalized to the rank-minimizing mechanisms satisfying weak ETE. We show this by introducing the following example.
Example 1
Let $\mathcal{A=}\left\{a_{1},a_{2},a_{3}\right\}$ and $\mathcal{O=}\left\{o_{1},o_{2},o_{3},\varnothing\right\}$. Let $\psi=\left(o_{1},o_{2},\varnothing,o_{3}\right)$ and $\psi^{\prime}=\left(o_{1},o_{2},o_{3},\varnothing\right)$. A rank-minimizing mechanism $f$ is such that
$\left(f\left(\psi^{\prime},\psi,\psi\right)\right)_{a_{1}o_{1}}$ $=$ $\left(f\left(\psi,\psi^{\prime},\psi\right)\right)_{a_{2}o_{1}}=\left(f\left(\psi,\psi,\psi^{\prime}\right)\right)_{a_{3}o_{1}}=0,$ (4)
$f\left(p_{a_{1}},p_{a_{2}},p_{a_{3}}\right)$ $=$ $\mathop{\displaystyle\sum}\limits_{y^{\ast}\in\mathcal{Y}^{\ast}\left(p\right)}\frac{y^{\ast}}{\left|\mathcal{Y}^{\ast}\left(p\right)\right|}$
if $\left(p_{a_{1}},p_{a_{2}},p_{a_{3}}\right)$ is neither $\left(\psi^{\prime},\psi,\psi\right)$ nor $\left(\psi,\psi^{\prime},\psi\right)$ nor $\left(\psi,\psi,\psi^{\prime}\right)$. That is, if the best type and the second best type of the three agents are respectively $o_{1}$ and $o_{2}$ and the third best type of the two of them is $\varnothing$, then the agent revealing $o_{3}$ as the third best one obtains $o_{1}$ with probability 00. Trivially, $f$ satisfies weak ETE. However, whether $f$ satisfies ETE or not is dependent on $q_{o_{1}}$ and $q_{o_{2}}$.
Suppose $r_{a_{1}}=\psi^{\prime}$ and this rank-minimizing mechanism is used. First, we assume $q_{o_{1}}=q_{o_{2}}=1$. In this case, $r_{a_{1}}=\psi^{\prime}$ is not strategically dominated by $\psi$ and any other preference orders, because
$\left(f\left(\psi^{\prime},\psi^{\prime},\psi^{\prime}\right)\right)_{a_{1}o_{1}}$ $=$ $\left(f\left(\psi^{\prime},\psi^{\prime},\psi^{\prime}\right)\right)_{a_{1}o_{2}}=\left(f\left(\psi^{\prime},\psi^{\prime},\psi^{\prime}\right)\right)_{a_{1}o_{3}}=\frac{1}{3},$
$\left(f\left(\psi,\psi^{\prime},\psi^{\prime}\right)\right)_{a_{1}o_{1}}$ $=$ $\left(f\left(\psi,\psi^{\prime},\psi^{\prime}\right)\right)_{a_{1}o_{2}}=\left(f\left(\psi,\psi^{\prime},\psi^{\prime}\right)\right)_{a_{1}\varnothing}=\frac{1}{3}.$
That is, if $p_{a_{2}}=p_{a_{3}}=\psi^{\prime}$, then $a_{1}$ is worse off by revealing $\psi$ instead of $\psi^{\prime}$. Moreover, in this case, $f$ satisfies ETE and thus this is a fair rank-minimizing mechanism.
On the other hand, suppose $q_{o_{1}}=1$ and $q_{o_{2}}\geq 2$. In this case, $a_{1}$ must have either $o_{1}$ or $o_{2}$ if it chooses either $\psi$ or $\psi^{\prime}$ and thus, the unique interest of $a_{1}$ is increasing the probability to obtain $o_{1}$. Thus, in this case, $r_{a_{1}}=\psi^{\prime}$ is strategically dominated by $\psi$. Therefore, even if a fair rank-minimizing mechanism is used, truth-telling may be strategically dominated by some other strategy. Moreover, in this case, $f$ does not satisfy ETE. First, since $q_{o_{1}}+q_{o_{2}}\geq\left|\mathcal{A}\right|$, any agent is assigned to neither $o_{3}$ nor $\varnothing$,∅ , as long as the assignment is not wasteful. Thus, $\psi$ and $\psi^{\prime}$ are essentially the same in this case. However, $\left(f\left(\psi^{\prime},\psi,\psi\right)\right)_{a_{1}o_{1}}=0,$ but $\left(f\left(\psi^{\prime},\psi,\psi\right)\right)_{a_{1}o_{2}}>0$ or $\left(f\left(\psi^{\prime},\psi,\psi\right)\right)_{a_{1}o_{3}}>0$ must be satisfied. Therefore, in this case, $f$ does not satisfy ETE.
4 Refusal option
In this section, we consider the following three-stage game, which is a modification of that in Section 3; that is, the first and second stages in this game are the same as in the previous game. In the third stage, all agents decide whether accept or refuse the object type that is assigned to them. If an agent refuses their assignment, then the agent instead obtains the null object type $\varnothing$ (the outside option). In this game, the null object type has two characteristics: first, there are sufficiently many copies of this type to allocate to all individuals, and second, agents can always obtain it at no cost as long as they abandon their assignment.
This third stage is also considered by Featherstone (2020) but is not explicitly addressed by almost previous studies related to ours. [^2] Doğan and Yenmez (2019) also explicitly consider this stage. However, this stage is implicitly considered by many of them; that is, they consider an ex-post individually rational (or more strictly an ex-post efficient) matching as a requirement and if an assignment is ex-post individually rational, no agent refuses their assignment. See, for example, Bogomolnaia and Moulin (2001) on the property.
Moreover, note that, in this study, all agents are assumed to know the capacities of all object types $\left(q_{o}\right)_{o\in\mathcal{O}}$, when they reveal their preference. On the other hand, Featherstone (2020) considers the situation where any agents do not know the capacity of any object types. Due to this difference, some of our results differ from his.
In the third stage, the obvious optimal (dominant) strategy of an agent $a$ is to accept their assignment denoted by $o$ if $o$ acceptable; that is, if $R\left(o,r_{a}\right)<R\left(\varnothing,r_{a}\right)$ and refuse it if otherwise ($R\left(o,r_{a}\right)\geq R\left(\varnothing,r_{a}\right)$). Based on this optimal strategy in the third stage, an assignment obtained after the second stage denoted by $x$, and the profile of (true) preferences denoted by $r$, we let $g\left(x,r\right)$ be such that for all $a\in\mathcal{A}$,
$g\left(x,r\right)_{ao}$ $=$ $x_{ao}\text{ for all }R\left(o,r_{a}\right)<R\left(\varnothing,r_{a}\right)$
$g\left(x,r\right)_{ao}$ $=$ 0 for all $0\text{ for all }R\left(o,r_{a}\right)>R\left(\varnothing,r_{a}\right)$
$g\left(x,r\right)_{a\varnothing}$ $=$ $\mathop{\displaystyle\sum}\limits_{o:R\left(o,r_{a}\right)\geq R\left(\varnothing,r_{a}\right)}x_{ao}.$
That is, since an agent refuses any of the unacceptable object types when it is assigned, the probability that the null object type is assigned will increase by the probability that any of the unacceptable object types is assigned.
We reconsider strategic problems of a fair rank-minimizing mechanism. For a mechanism $f$, (revealing) $p_{a}^{\prime}$ (for $a$) is said to be strategically dominated by (revealing) $p_{a}$ with the refusal option if $a$ weakly prefers $g\left(f\left(p_{a},p_{-a}\right),r\right)$ to $g\left(f\left(p_{a}^{\prime},p_{-a}\right),r\right)$ for all $p_{-a}\in\mathcal{P}^{\left|\mathcal{A}\right|-1}$ and strictly prefers $g\left(f\left(p_{a},p_{-a}^{\prime}\right),r\right)$ to $g\left(f\left(p_{a}^{\prime},p_{-a}^{\prime}\right),r\right)$ for some $p_{-a}^{\prime}\in\mathcal{P}^{\left|\mathcal{A}\right|-1}$.
We show that, under the uniform rank-minimizing mechanism, truth-telling may be strategically dominated by another strategy when the agents have the refusal option. We let $\mathcal{D}_{a}\subseteq\mathcal{P}$ be a set of preferences such that $d_{a}\in\mathcal{D}_{a}$ if (i) $R\left(o,r_{a}\right)=R\left(o,d_{a}\right)$ for all $o\in\mathcal{O}$ such that $R\left(o,r_{a}\right)<R\left(\varnothing,r_{a}\right)$ and (ii) $R\left(\varnothing,d_{a}\right)=\left|\mathcal{O}\right|$. First, the ranks of the truly acceptable object types are unchanged. Second, the null object type $\varnothing$ is the least preferable one under $d_{a}\in\mathcal{D}_{a}$. Thus, any object type that is less preferable than $\varnothing$ in the true preference has changed to be more preferable than $\varnothing$ in $d_{a}$. We call the strategy of agent $a$ of revealing a preference in $\mathcal{D}_{a}$ an outside option demotion strategy (hereafter ODS) of agent $a$. Note that an ODS (revealing $d_{a}\in\mathcal{D}_{a}$) increases the number of acceptable object types revealed by $a$; that is, $R\left(\varnothing,r_{a}\right)\leq R\left(\varnothing,d_{a}\right)=\left|\mathcal{O}\right|$ is satisfied.
Featherstone (2020) defines the full extension of $r_{a}$denoted by $e_{a}$ in this study such that (i) $R\left(o,r_{a}\right)=R\left(o,e_{a}\right)$ for all $o\in\mathcal{O}$ such that $R\left(o,r_{a}\right)<R\left(\varnothing,e_{a}\right)$, and (ii) $R\left(\varnothing,e_{a}\right)=\left|\mathcal{O}\right|$, and $R\left(o,r_{a}\right)-1=R\left(o,e_{a}\right)$ for all $o\in\mathcal{O}$ with $R\left(o,r_{a}\right)>R\left(\varnothing,e_{a}\right)$; that is, $e_{a}\in\mathcal{D}_{a}$ is such that all non-null object types are placed in their true preference orders.
First, we consider the following example to briefly consider the main result.
Example 2
Let $\mathcal{A=}\left\{a_{1},a_{2},a_{3}\right\}$ and $\mathcal{O=}\left\{o_{1},o_{2},\varnothing\right\}$. Suppose $q_{o_{1}}=q_{o_{2}}=1$ and $r_{a_{1}}=(o_{1},\varnothing,o_{2})$. Then, the unique ODS is revealing $d_{a_{1}}=(o_{1},o_{2},\varnothing)$. First, suppose
$p_{a_{2}}=(o_{1},o_{2},\varnothing)\text{ and }p_{a_{3}}=(o_{2},o_{1},\varnothing).$
Then,
$f^{U}\left(r_{a_{1}},p_{a_{2}},p_{a_{3}}\right)=y=\left(\begin{array}[]{ccc}y_{a_{1}o_{1}}&y_{a_{1}o_{2}}&y_{a_{1}\varnothing}\\ y_{a_{2}o_{1}}&y_{a_{2}o_{2}}&y_{a_{2}\varnothing}\\ y_{a_{3}o_{1}}&y_{a_{3}o_{2}}&y_{a_{3}\varnothing}\end{array}\right)=\left(\begin{array}[]{ccc}0&0&1\\ 1&0&0\\ 0&1&0\end{array}\right).$
Whenever $\left|\mathcal{A}\right|=\left|\mathcal{O}\right|=3$, we write the matrix representing an assignment in this manner. On the other hand,
$f^{U}\left(d_{a_{1}},p_{a_{2}},p_{a_{3}}\right)=\left(\begin{array}[]{ccc}1/2&0&1/2\\ 1/2&0&1/2\\ 0&1&0\end{array}\right)$
Thus, revealing $d_{a_{1}}$ instead of $r_{a_{1}}$ may increase the probability that $a_{1}$ is assigned to their best object $o_{1}$. On the other hand, suppose
$p_{a_{2}}^{\prime}=(o_{1},\varnothing,o_{2})\text{ and }p_{a_{3}}^{\prime}=(o_{1},\varnothing,o_{2}).$
Then,
$f^{U}\left(r_{a_{1}},p_{a_{2}}^{\prime},p_{a_{3}}^{\prime}\right)=\left(\begin{array}[]{ccc}1/3&0&2/3\\ 1/3&0&2/3\\ 1/3&0&2/3\end{array}\right),$
$f^{U}\left(d_{a_{1}},p_{a_{2}}^{\prime},p_{a_{3}}^{\prime}\right)=\left(\begin{array}[]{ccc}1/3&2/3&0\\ 1/3&0&2/3\\ 1/3&0&2/3\end{array}\right).$
Thus, in this case, $a_{1}$ seems to be worse off by revealing $d_{a_{1}}$ instead of $r_{a_{1}}$, because $a_{1}$ prefers $\varnothing$ to $o_{2}$. However, in the model introduced in this section, $a_{1}$ can refuse $o_{2}$ and in that case, $\varnothing$ is assigned. Therefore,
$g\left(f^{U}\left(d_{a_{1}},p_{a_{2}}^{\prime},p_{a_{3}}^{\prime}\right),r\right)=g\left(f^{U}\left(r_{a_{1}},p_{a_{2}}^{\prime},p_{a_{3}}^{\prime}\right),r\right)\text{;}$
that is, $a_{1}$ is not worse off by choosing $d_{a_{1}}$ instead of $r_{a_{1}}$.
Next, let $\hat{p}_{a_{2}}=\hat{p}_{a_{3}}=(o_{1},o_{2},\varnothing)$. Then,
$f^{U}\left(r_{a_{1}},\hat{p}_{a_{2}},\hat{p}_{a_{3}}\right)=\left(\begin{array}[]{ccc}0&0&1\\ 1/2&1/2&0\\ 1/2&1/2&0\end{array}\right),$
$f^{U}\left(d_{a_{1}},\hat{p}_{a_{2}},\hat{p}_{a_{3}}\right)=\left(\begin{array}[]{ccc}1/3&1/3&1/3\\ 1/3&1/3&1/3\\ 1/3&1/3&1/3\end{array}\right).$
In the latter case, $a_{1}$ may refuse the assignment; that is,
$g\left(f^{U}\left(d_{a_{1}},\hat{p}_{a_{2}},\hat{p}_{a_{3}}\right),r\right)=\left(\begin{array}[]{ccc}1/3&0&2/3\\ 1/3&1/3&1/3\\ 1/3&1/3&1/3\end{array}\right),$
which is wasteful, because
$\mathop{\textstyle\sum}\limits_{a\in\left\{a_{1},a_{2},a_{3}\right\}}g\left(f^{U}\left(d_{a_{1}},\hat{p}_{a_{2}},\hat{p}_{a_{3}}\right),r\right)_{ao_{2}}$ $=$ $\frac{2}{3}<1,$
and $\text{and }g\left(f^{U}\left(d_{a_{1}},\hat{p}_{a_{2}},\hat{p}_{a_{3}}\right),r\right)_{a^{\prime}\varnothing}$ $=$ $\frac{1}{3}>0,$
for $a^{\prime}=a_{2},a_{3}$. Moreover, it is ordinally dominated by
$\tilde{x}=\left(\begin{array}[]{ccc}1/3&0&2/3\\ 1/3&1/2&1/6\\ 1/3&1/2&1/6\end{array}\right).$
In the case where $a_{1}$ is assigned to $o_{2}$, either $a_{2}$ or $a_{3}$ is assigned to $\varnothing$ but $a_{1}$ will refuse $o_{2}$. Thus, in that case, it is more efficient that the agent who is assigned to $\varnothing$ is assigned to $o_{2}$ and $a_{1}$ is assigned to $\varnothing$. That is, revealing $d_{a_{1}}$ and refusing some unacceptable assignment may make $a_{1}$ better off but result in a wasteful assignment, because some agents may desire the object type refused by $a_{1}$.
Note that $\left(d_{a_{1}},\hat{p}_{a_{2}},\hat{p}_{a_{3}}\right)$ can be a Nash equilibrium when $r_{a_{2}}=\hat{p}_{a_{2}}$ and $r_{a_{3}}=\hat{p}_{a_{3}}$. Trivially, $a_{1}$ has no incentive to change $d_{a_{1}}$. On the other hand, suppose that both $a_{2}$ and $a_{3}$ prefer the assignment such that each object $o\in\left\{o_{1},o_{2},\varnothing\right\}$ assigns with the probability $1/3$ to the assignment such that $o_{2}$ assigns with probability $1$. This ensures that neither $a_{2}$ nor $a_{3}$ has an incentive to change $\hat{p}_{a_{2}}$ and $\hat{p}_{a_{3}}$ to $(o_{2},o_{1},\varnothing)$ or $(o_{2},\varnothing,o_{1})$. Then, in this case, $\left(d_{a_{1}},\hat{p}_{a_{2}},\hat{p}_{a_{3}}\right)$ is a Nash equilibrium. Thus, a wasteful assignment can occur as a Nash equilibrium of agents’ strategies when the assignment is determined by the uniform rank-minimizing mechanism.
Now, we formally introduce our main results.
Theorem 1
Suppose that the uniform rank-minimizing mechanism used. Then, for all agents, truth-telling is weakly strategically dominated by any ODS with the refusal option.
By this result, an agent is never hurt by using an ODS instead of truth-telling. Next, we show that with some weak condition, there exists some ODS such that an agent may be better off by using it instead of truth-telling.
Theorem 2
Suppose that there is $a\in\mathcal{A}$ whose true preference $r_{a}$ satisfies
$R\left(o,r_{a}\right)<R\left(\varnothing,r_{a}\right)<R\left(o^{\prime},r_{a}\right)\text{ and }\mathop{\textstyle\sum}\limits_{\bar{o}:R\left(\bar{o},r_{a}\right)<R\left(\varnothing,r_{a}\right)}q_{\bar{o}}+q_{o^{\prime}}<\left|\mathcal{A}\right|\text{,}$ (5)
[…]
This result implies that if there is $a\in\mathcal{A}$ satisfying (5) for some $o,o^{\prime}\in\mathcal{O}$, then agent $a$ has a strong incentive to adopt an ODS. This is satisfied only when some agent has at least one unacceptable object. The condition given in (5) is that there is a truly unacceptable type $o^{\prime}$ such that the number of copies of $o^{\prime}$, plus the sum of the copies of the truly acceptable types, is not very large.
We discuss the cases where (5) is satisfied for some $a,o,o^{\prime}$ in the subsequent section. Note that there is $p_{-a}\in\mathcal{P}^{\left|\mathcal{A}\right|-1}$ such that $R\left(\varnothing,p_{a^{\prime}}\right)=\left|\mathcal{O}\right|$ for all $a^{\prime}\in\mathcal{A\setminus}\left\{a\right\}$ and $a$ strictly prefers $g\left(f^{U}\left(d_{a},p_{-a}\right),r\right)$ to $g\left(f^{U}\left(r_{a},p_{-a}\right),r\right)$. Thus, for all $a^{\prime}$, $p_{a^{\prime}}$ may also be an ODS of $a^{\prime}$.
Theorems 1 and 2 do not say that an ODS is a dominant strategy. However, when if there is $a\in\mathcal{A}$ whose true preference satisfies (5) for some $o,o^{\prime}\in\mathcal{O}$, since truth-telling is strategically dominated by another strategy, $a$ has a strong incentive to manipulate its preference ranking, while one of the ODSs may not necessarily be chosen.
As stated earlier, there may exist multiple ODSs. In that case, an ODS may be strategically dominated by another one. We have the following result.
Proposition 3
Suppose that there is $a\in\mathcal{A}$ whose true preference $r_{a}$ satisfies
$R\left(o,r_{a}\right)<R\left(\varnothing,r_{a}\right)<R\left(o^{\prime},r_{a}\right)\text{ and }\mathop{\textstyle\sum}\limits_{\bar{o}:R\left(\bar{o},r_{a}\right)<R\left(\varnothing,r_{a}\right)}q_{\bar{o}}+q_{o^{\prime}}\geq\left|\mathcal{A}\right|,$
for some $o,o^{\prime}\in\mathcal{O}$. Let $d_{a}\in\mathcal{D}_{a}$ be such that $R\left(\varnothing,r_{a}\right)=R\left(o^{\prime},d_{a}\right)$. Then, for all $p_{-a}\in\mathcal{P}^{\left|\mathcal{A}\right|-1}$ and all $\bar{o}\in\mathcal{O}$,
$g\left(f^{U}\left(d_{a},p_{-a}\right),r\right)_{a\bar{o}}=g\left(f^{U}\left(r_{a},p_{-a}\right),r\right)_{a\bar{o}}.$
In this case, truth-telling is not strategically dominated by an ODS of revealing such $d_{a}$. Note that revealing the full extension may be such an ODS, when $o^{\prime}$ defined in Proposition 3 satisfies $R\left(\varnothing,r_{a}\right)+1=R\left(o^{\prime},r_{a}\right)$.
Moreover, by Theorem 1 and Proposition 3, we show that for an agent, revealing the full extension may be strategically dominated by another ODS. First, we provide an example to briefly see this fact.
Example 3
Let $\mathcal{A=}\left\{a_{1},a_{2},a_{3}\right\}$ and $\mathcal{O=}\left\{o_{1},o_{2},o_{3},\varnothing\right\}$. Suppose $q_{o_{1}}=q_{o_{3}}=1,$ $q_{o_{2}}=2$ and $r_{a_{1}}=(o_{1},\varnothing,o_{2},o_{3})$. Then, there are two ODSs. One is $e_{a_{1}}=(o_{1},o_{2},o_{3},\varnothing)$, which is the full extension of $a_{1}$, and the other is $d_{a_{1}}=(o_{1},o_{3},o_{2},\varnothing)$. First, suppose
$\bar{p}_{a_{2}}=(o_{1},o_{3},o_{2},\varnothing)\text{ and }\bar{p}_{a_{3}}=(o_{3},o_{1},o_{2},\varnothing).$
Then,
$f^{U}\left(r_{a_{1}},\bar{p}_{a_{2}},\bar{p}_{a_{3}}\right)=y=\left(\begin{array}[]{cccc}y_{a_{1}o_{1}}&y_{a_{1}o_{2}}&y_{a_{1}o_{3}}&y_{a_{1}\varnothing}\\ y_{a_{2}o_{1}}&y_{a_{2}o_{2}}&y_{a_{2}o_{3}}&y_{a_{2}\varnothing}\\ y_{a_{3}o_{1}}&y_{a_{3}o_{2}}&y_{a_{3}o_{3}}&y_{a_{3}\varnothing}\end{array}\right)=\left(\begin{array}[]{cccc}0&0&0&1\\ 1&0&0&0\\ 0&0&1&0\end{array}\right),$
In this example, we write the assignments in this manner. Moreover,
$g\left(f^{U}\left(d_{a_{1}},\bar{p}_{a_{2}},\bar{p}_{a_{3}}\right),r\right)=\left(\begin{array}[]{cccc}1/2&0&0&1/2\\ 1/2&1/2&0&0\\ 0&0&1&0\end{array}\right)$
Therefore, $a_{1}$ (strictly) prefers $g\left(f^{U}\left(d_{a_{1}},\bar{p}_{a_{2}},\bar{p}_{a_{3}}\right),r\right)$ to $g\left(f^{U}\left(r_{a_{1}},\bar{p}_{a_{2}},\bar{p}_{a_{3}}\right),r\right)$. On the other hand, there is no pair $\left(p_{a_{2}},p_{a_{3}}\right)$ of preference orders of $a_{2}$ and $a_{3}$ such that $a_{1}$ (strictly) prefers $g\left(f^{U}\left(e_{a_{1}},p_{a_{2}},p_{a_{3}}\right),r\right)$ to $g\left(f^{U}\left(r_{a_{1}},p_{a_{2}},p_{a_{3}}\right),r\right)$. This is because, by Proposition 3 and $q_{o_{1}}+q_{o_{2}}=3=\left|\mathcal{A}\right|$, for any $\left(p_{a_{2}},p_{a_{3}}\right)\in\mathcal{P}^{2}$ and any $\bar{o}\in\mathcal{O}$,
$\left(g\left(f^{U}\left(e_{a_{1}},p_{a_{2}},p_{a_{3}}\right),r\right)\right)_{a_{1}\bar{o}}=\left(g\left(f^{U}\left(r_{a_{1}},p_{a_{2}},p_{a_{3}}\right),r\right)\right)_{a_{1}\bar{o}}\text{.}$
We explain why $d_{a_{1}}$ strategically dominates $e_{a_{1}}$. In this example, $q_{o_{1}}+q_{o_{3}}=2<3=q_{o_{1}}+q_{o_{2}}$. That is, the copies of $o_{1}$ and $o_{2}$ are sufficient to allocate all agents, but those of $o_{1}$ and $o_{3}$ are insufficient. Thus, revealing $o_{2}$ as a second best type does not increase the probability that $a_{1}$ is assigned to $o_{1}$. On the other hand, revealing $o_{3}$ as a second best type increases it, when $o_{3}$ is the first best type of another agent. Thus, revealing the full extension may be strategically dominated by another ODS when an unacceptable type (in this case $o_{3}$) is less preferable than the other (in this case $o_{2}$) but the number of copies of the former is less than that of the latter.
Generally, we have the following result.
Lemma 4
Suppose that there is $a\in\mathcal{A}$ whose true preference satisfies
$R\left(o,r_{a}\right)<R\left(\varnothing,r_{a}\right)=R\left(o^{\prime},r_{a}\right)-1<R\left(o^{\prime\prime},r_{a}\right),\text{ }$
$\mathop{\textstyle\sum}\limits_{\bar{o}:R\left(\bar{o},r_{a}\right)<R\left(\varnothing,r_{a}\right)}q_{\bar{o}}+q_{o^{\prime}}\geq\left|\mathcal{A}\right|,\text{ and}$
$\mathop{\textstyle\sum}\limits_{\bar{o}:R\left(\bar{o},r_{a}\right)<R\left(\varnothing,r_{a}\right)}q_{\bar{o}}+q_{o^{\prime\prime}}<\left|\mathcal{A}\right|,$
[…]
$R\left(\varnothing,r_{a}\right)=R\left(o^{\prime},e_{a}\right)=R\left(o^{\prime\prime},d_{a}\right),$
with the refusal option.
Note that Lemma 4 does not contradict a result of Featherstone (2020, Proposition 13). In his model, he assumes that the agents have no information on the capacities of object types. Revealing the full extension can be strategically dominated by another ODS only when the capacities of the unacceptable types are different and agents know them. Thus, the strategic domination in Lemma 4 does not occur in the low-information situation discussed by Featherstone (2020).
Next, we show that if agents adopt an ODS, then the result may be wasteful and thus inefficient.
Proposition 4
Suppose that there is $a\in\mathcal{A}$ whose true preference satisfies (5) for some $o,o^{\prime}\in\mathcal{O}$. Then, there are some ODS of revealing $d_{a}$ and some $p_{-a}\in\mathcal{P}^{\left|\mathcal{A}\right|-1}$ such that $R\left(\varnothing,p_{a^{\prime}}\right)=\left|\mathcal{O}\right|$ for all $a^{\prime}\in\mathcal{A\setminus}\left\{a\right\}$ and $g\left(f^{U}\left(d_{a},p_{-a}\right),r\right)$ is wasteful.
By Theorem 2, for an agent $a$ with $r_{a}$ that satisfy (5) for some $o,o^{\prime}\in\mathcal{O}$, an ODS dominates truth-telling. Moreover, by Proposition 1, in the same case, the strategy may make the assignment wasteful and thus inefficient. In an ODS, an agent reveals actually unacceptable types as acceptable types. Then, the agent may be assigned to an actually unacceptable type and refuse it. Since the type may be desirable for another agent, in that case, the assignment is wasteful.
By Theorems 1 and 2 and Proposition 3, if the uniform rank-minimizing mechanism is used, the agents may have a strong incentive to manipulate their preferences, and these manipulations may result in an inefficient assignment. Thus, next, we consider another fair rank-minimizing mechanism such that, contrary to $f^{U}$, if it is used as the mechanism in the second stage, then for any agent, truth-telling is not strategically dominated by any other strategy.
We let $f^{M}$ be the modified uniform rank-minimizing mechanism such that
$\left(f^{M}\left(p\right)\right)_{ao}=0,\text{ }\left(f^{M}\left(p\right)\right)_{a^{\prime}o}=1,\text{ }\left(f^{M}\left(p\right)\right)_{ap_{a}^{2}}=1\text{ and }\left(f^{M}\left(p\right)\right)_{a^{\prime\prime}\varnothing}=1$
if $p$ satisfies
there is one agent $a$ such that $R\left(o,p_{a}\right)=1$ and $R\left(\varnothing,p_{a}\right)\geq 3,$
there are $q_{o}$ agents (including $a^{\prime}$) with revealed preferences equivalent to those represented by $p_{a}^{\prime}$, such that $R\left(\varnothing,p_{a^{\prime}}\right)=l\geq 2,$ $p_{a}^{l^{\prime}}=p_{a^{\prime}}^{l}$ for all $l^{\prime}=1,\cdots,l-1,$ $R\left(\varnothing,p_{a}\right)>R\left(\varnothing,p_{a^{\prime}}\right),$ and
$\mathop{\displaystyle\sum}\limits_{\bar{o}:R\left(\bar{o},p_{a^{\prime}}\right)<R\left(\varnothing,p_{a^{\prime}}\right)}q_{\bar{o}}<\left|\mathcal{A}\right|,$
if $\left|\mathcal{A}\right|>q_{o}+1$, then the first best type of the other agents ($\left|\mathcal{A}\right|-\left(q_{o}+1\right)$ agents) represented by $a^{\prime\prime}$ is $\varnothing$,
and if otherwise
$f^{M}\left(p\right)=\mathop{\displaystyle\sum}\limits_{y^{\ast}\in\mathcal{Y}^{\ast}\left(p\right)}\frac{y^{\ast}}{\left|\mathcal{Y}^{\ast}\left(p\right)\right|}=f^{U}\left(p\right).$
To understand this mechanism, we reconsider Example 1. First, suppose $q_{o_{1}}=q_{o_{2}}=1$. If $\left(p_{a_{1}},p_{a_{2}},p_{a_{3}}\right)=\left(\psi^{\prime},\psi,\psi\right)$, $\left(\psi,\psi^{\prime},\psi\right)$ or $\left(\psi,\psi,\psi^{\prime}\right)$, then this satisfies (i), (ii) and (iii). Hence the rank-minimizing mechanism $f$ introduced in Example 1 is the modified uniform rank-minimizing mechanism.
On the other hand, suppose $q_{o_{1}}=1$ and $q_{o_{2}}\geq 2$. Even if $\left(p_{a_{1}},p_{a_{2}},p_{a_{3}}\right)=\left(\psi^{\prime},\psi,\psi\right)$, $\left(\psi,\psi^{\prime},\psi\right)$ or $\left(\psi,\psi,\psi^{\prime}\right)$, this does not satisfy (ii), because in this case, $p_{a^{\prime}}=\psi$ and $l=3.$ Then,
$\mathop{\displaystyle\sum}\limits_{\bar{o}:R\left(\bar{o},p_{a^{\prime}}\right)<R\left(\varnothing,p_{a^{\prime}}\right)}q_{\bar{o}}=q_{o_{1}}+q_{o_{2}}=3=\left|\mathcal{A}\right|.$
Therefore, (ii) is not satisfied with $p_{a^{\prime}}=\psi$.
We have the following result.
Proposition 5
The modified uniform rank-minimizing mechanism $f^{M}$ satisfies ETE.
The strategic issues, mentioned in Theorem 2, do not arise under this mechanism $f^{M}$.
First, we explain why the strategic issues mentioned in Example 2 does not arise in the modified uniform rank-minimizing mechanism. If $r_{a}=p_{a^{\prime}}$ where $p_{a^{\prime}}$ is defined in (ii), then $\left(d_{a},p_{-a}\right)$ satisfies (i), (ii) and (iii) where $d_{a}\in\mathcal{D}_{a}$. This implies that for some $p_{-a}$, if agent $a$ adopts an ODS instead of truth-telling and $f^{M}$ is used, then $a$ is worse off.
To see this fact, we revisit Example 2. In this example, $\left(d_{a_{1}},p_{a_{2}}^{\prime},p_{a_{3}}^{\prime}\right)$ satisfies (i), (ii) and (iii), but $\left(r_{a_{1}},p_{a_{2}}^{\prime},p_{a_{3}}^{\prime}\right)$ does not. Thus,
$f^{M}\left(r_{a_{1}},p_{a_{2}}^{\prime},p_{a_{3}}^{\prime}\right)$ $=$ $f^{U}\left(r_{a_{1}},p_{a_{2}}^{\prime},p_{a_{3}}^{\prime}\right)=\left(\begin{array}[]{ccc}1/3&0&2/3\\ 1/3&0&2/3\\ 1/3&0&2/3\end{array}\right),$
$f^{M}\left(d_{a_{1}},p_{a_{2}}^{\prime},p_{a_{3}}^{\prime}\right)$ $=$ $\left(\begin{array}[]{ccc}0&1&0\\ 1/2&0&1/2\\ 1/2&0&1/2\end{array}\right),$
$g\left(f^{M}\left(d_{a_{1}},p_{a_{2}}^{\prime},p_{a_{3}}^{\prime}\right)\right)$ $=$ $\left(\begin{array}[]{ccc}0&0&1\\ 1/2&0&1/2\\ 1/2&0&1/2\end{array}\right).$
This implies that $a_{1}$ may be worse off by revealing $d_{a_{1}}$ instead of $r_{a_{1}}$, even if $a_{1}$ refuses any unacceptable objects. Therefore, $r_{a_{1}}$ is not strategically dominated by $d_{a_{1}}$, if $f^{M}$ is used.
Formally, we have the following fact.
Proposition 6
Suppose that the modified uniform rank-minimizing mechanism is used. For any $r_{a}\in\mathcal{P}$, there is no $p_{a}\in\mathcal{P}$ that strategically dominates $r_{a}$ with (or without) the refusal option.
By adopting $f^{M}$, any ODS can fail, because increasing the number of acceptable objects may lead the agent to be unassigned acceptable objects in their true preference. Thus, in some profiles of other agents’ (revealed) preferences, truth-telling is better than adopting any ODS. However, $f^{M}$ may lead to other strategic problems. That is, if $f^{M}$ is adopted instead of $f^{U}$, an agent may have an incentive to reduce the number of acceptable object types.
Example 4
Let $R\left(\varnothing,r_{a}\right)=k$. Suppose $k\geq 3$; that is, $\varnothing$ in neither the best type nor the second best type for $r_{a}$. Thus, we let $r_{a}^{1}=o$ and $r_{a}^{2}=o^{\prime}$. For $p_{-a}$, there are $q_{o}$ agents represented by $a^{\prime}$ whose revealed preference order is such that $p_{a^{\prime}}^{k^{\prime}}=r_{a^{\prime}}^{k^{\prime}}$ for all $k^{\prime}=1,\cdots,k-2$ and $p_{a^{\prime}}^{k-1}=\varnothing$. We assume
$q_{r_{a}^{1}}+\cdots+q_{r_{a}^{k-2}}<\left|\mathcal{A}\right|\text{.}$
Moreover, for $p_{-a}$ and any other agent represented by $a^{\prime\prime}$, $R\left(\varnothing,r_{a^{\prime\prime}}\right)=1$. In this case, since $\left(r_{a},p_{-a}\right)$ satisfies (i), (ii) and (iii),
$\left(f^{M}\left(r_{a},p_{-a}\right)\right)_{ao^{\prime}}$ $=$ 0, ($0,\text{ }\left(f^{M}\left(r_{a},p_{-a}\right)\right)_{ao^{\prime}}=1,$
$\left(f^{U}\left(r_{a},p_{-a}\right)\right)_{ao}$ $=$ $\frac{q_{o}}{q_{o}+1},$
$\left(f^{U}\left(r_{a},p_{-a}\right)\right)_{ao^{\prime}}$ $=$ $\frac{1}{q_{o}+1}.$
On the other hand, we let $p_{a}=p_{a^{\prime}}$. Since $\left(p_{a},p_{-a}\right)$ does not satisfy all of (i), (ii) and (iii),
$\left(f^{M}\left(p_{a},p_{-a}\right)\right)_{ao}$ $=$ $\left(f^{U}\left(p_{a},p_{-a}\right)\right)_{ao}=\frac{q_{o}}{q_{o}+1},$
$\left(f^{M}\left(p_{a},p_{-a}\right)\right)_{ao^{\prime}}$ $=$ $\left(f^{U}\left(p_{a},p_{-a}\right)\right)_{ao^{\prime}}=\frac{1}{q_{o}+1}.$
Thus, in this case, if $f^{M}$ is used, then $a$ has an incentive to reveal by reducing its truly acceptable types. On the other hand, if $f^{U}$ is used, then $a$ has no such an incentive.
Of course, this does not mean that truth-telling is strategically dominated revealing some reduction of $r_{a}$. However, agents may have such incentives even if they do not have the refusal option.
5 Concluding Remarks
We show that if an assignment is determined by the uniform rank-minimizing mechanism, which is a fair rank-minimizing mechanism, and agents can refuse the assignment and obtain the outside option instead, then the true preference is strategically dominated by an ODS. The adoption of such strategies by agents may result in inefficient assignments; that is, an agent who adopts an ODS refuses their assigned type, although some other agents desire that type.
We consider a public school choice system as an example. Our result implies that a student may have an incentive to reveal truly unacceptable schools as acceptable schools, and they may be assigned to one of them and then refuse the admission even if some other students desire the school. In such a case, the school may re-enroll students, but even in that case, it would cause market disruption.
Our result implies that if all agents prefer many other types over the outside option, then the strategic problem considered in this study is not so severe. For example, in the case of matching systems for schools that are part of compulsory education, the outside option does not need to be considered. Therefore, there is no strategic problem of the kind discussed in this study. On the other hand, the strategic problem may be severe when there are agents who relatively highly value the outside option.
Moreover, if agents who refuse their assignment can be sufficiently penalized, the problem would be resolved. However, considering that some agents may refuse their assignment due to unavoidable circumstances (without any strategic intention), the penalties should not be too severe.
Therefore, if resolving the problem in that way is difficult, it is necessary to consider alternative mechanisms. First, as shown in this study, if we modify the uniform rank-minimizing mechanism, then truth-telling is not strategically dominated by any other strategies, including the outside option demotion ones. However, in this modification, it is necessary to ensure that increasing the number of acceptable types can result in a disadvantage. Thus, this modification may lead to the opposite problem, where agents strategically decrease the number of acceptable types. Second, we consider efficient and fair mechanisms other than the fair rank-minimizing mechanisms, such as the (general) probabilistic serial mechanism introduced by Bogomolnaia and Moulin (2001) and Budish et al. (2013), because it is fair and ordinally efficient, although as stated earlier, this efficiency property is weaker than the rank efficiency.
Appendix
Proof of Lemma 1
First, by Remark 1, any $x^{\ast}\in\mathcal{X\setminus Y}$ can be represented by a convex combination of deterministic assignments $y^{1},\cdots,y^{I}$ with $I\geq 2$. Suppose not; that is, there is $y^{i}$ such that $RV\left(y^{i},p\right)>RV\left(x^{\ast},p\right)$.
$RV\left(x^{\ast},p\right)$ $=$ $\mathop{\displaystyle\sum}\limits_{a\in\mathcal{A}}\mathop{\displaystyle\sum}\limits_{o\in\mathcal{O}}R\left(o,p_{a}\right)\times x_{ao}^{\ast}$ (9)
$=$ $\mathop{\displaystyle\sum}\limits_{a\in\mathcal{A}}\mathop{\displaystyle\sum}\limits_{o\in\mathcal{O}}\left(R\left(o,p_{a}\right)\times\mathop{\displaystyle\sum}\limits_{i=1}^{I}\sigma^{i}y_{ao}^{i}\right)$
$=$ $\mathop{\displaystyle\sum}\limits_{i=1}^{I}\sigma^{i}\mathop{\displaystyle\sum}\limits_{a\in\mathcal{A}}\mathop{\displaystyle\sum}\limits_{o\in\mathcal{O}}R\left(o,p_{a}\right)\times y_{ao}^{i}$
$=$ $\mathop{\displaystyle\sum}\limits_{i=1}^{I}\sigma^{i}RV\left(y^{i},p\right).$
Thus, $RV\left(y^{i},p\right)>RV\left(x^{\ast},p\right)$ implies that there is $y^{j}$ such that $RV\left(y^{j},p\right)<RV\left(x^{\ast},p\right)$, which contradicts that $x^{\ast}$ is a rank efficient assignment.
Proof of Lemma 2
Fix $p\in\mathcal{P}^{\left|\mathcal{A}\right|}$ and $a\in\mathcal{A}$. Moreover, let $k$ be the smallest integer satisfying (2) for $p_{a}$. Suppose not; that is, for an assignment $x\in\mathcal{X}$ that is not wasteful, there is $o\in\mathcal{O}$ satisfying $R(o,p_{a})\geq k$ and $x_{ao}>0$. First, since $\mathop{\textstyle\sum}_{\bar{o}\in\mathcal{O}}x_{a\bar{o}}=1$,
$\mathop{\displaystyle\sum}\limits_{\bar{o}\in\mathcal{O}}\mathop{\displaystyle\sum}\limits_{\bar{a}\in\mathcal{A}}x_{\bar{a}\bar{o}}=\left|\mathcal{A}\right|\text{.}$
Moreover, since $x_{ao}>0$ and (2) are satisfied,
$\mathop{\displaystyle\sum}\limits_{\bar{o}:R(\bar{o},p_{a})<k}\mathop{\displaystyle\sum}\limits_{\bar{a}\in\mathcal{A}}x_{\bar{a}o^{\prime}}$ $<$ $\left|\mathcal{A}\right|\leq\mathop{\displaystyle\sum}\limits_{\bar{o}:R(\bar{o},p_{a})<k}q_{o^{\prime}}\text{, and}$
$\mathop{\displaystyle\sum}\limits_{\bar{o}:R(\bar{o},p_{a})<k}x_{a\bar{o}}$ $<$ $1$
Since $\mathop{\textstyle\sum}_{\bar{a}\in\mathcal{A}}x_{\bar{a}o^{\prime\prime}}\leq q_{o^{\prime\prime}}$ for all $o^{\prime\prime}\in\mathcal{O}$, there is some $o^{\prime}\in\mathcal{O}$ such that $R(o^{\prime},p_{a})\leq k$ and $\mathop{\textstyle\sum}_{\bar{a}\in\mathcal{A}}x_{\bar{a}o^{\prime}}<q_{o^{\prime}}$. Then, $\mathop{\textstyle\sum}_{\bar{a}\in\mathcal{A}}x_{\bar{a}o^{\prime}}<q_{o^{\prime}},$ $x_{ao}>0,$ and $R(o^{\prime},p_{a})\leq k<R(o,p_{a})$ contradict that $x\in\mathcal{X}$ is not wasteful.
Proof of Proposition 1
Let $p_{a}\in\mathcal{P}$ and $k$ be the smallest integer satisfying (2) for $p_{a}$. Moreover, let $p\in\mathcal{P}^{\left|\mathcal{A}\right|}$ satisfy $p_{a}=p_{a^{\prime}}$. Suppose not; that is, $\left(f^{U}\left(p\right)\right)_{ao}>\left(f^{U}\left(p\right)\right)_{a^{\prime}o}$ for some $o\in\mathcal{O}$. Since $\left(f^{U}\left(p\right)\right)_{ao}>0$ and $f^{U}\left(p\right)$ is not wasteful, $R\left(o,p_{a}\right)<k$. Therefore, $R\left(o,p_{a}\right)=R\left(o,p_{a^{\prime}}\right)$.
Now, let
$Y=\left\{y\in\mathcal{Y}^{\ast}\left(p\right)\left|\text{ }y_{ao}=1\right.\right\},Y^{\prime}=\left\{y^{\ast}\in\mathcal{Y}^{\ast}\left(p\right)\left|\text{ }y_{a^{\prime}o}^{\ast}=1\right.\right\}.$
Since $\left(f^{U}\left(p\right)\right)_{ao}>0$, $Y\neq\emptyset$ and thus let $Y=\left\{y^{1},\cdots,y^{\left|Y\right|}\right\}$. Then, for each $i=1,\cdots,\left|Y\right|$, there is $y^{i\prime}$ be such that $y_{ao}^{i\prime}=0,$ $y_{a^{\prime}o}^{i\prime}=1,$ and $y_{a^{\prime\prime}o^{\prime}}^{i\prime}=y_{a^{\prime\prime}o^{\prime}}^{i}$ for all $o^{\prime}\in\mathcal{O}\setminus\{o\}$ and all $a^{\prime\prime}\in\mathcal{A}\setminus\{a,a^{\prime}\}$. Since $R\left(o,p_{a}\right)=R\left(o,p_{a^{\prime}}\right)$, $RV\left(y^{i},p\right)=RV\left(y^{i\prime},p\right)$ and therefore $y^{i\prime}\in Y^{\prime}$. Moreover, since $y^{1\prime},$ $\cdots,y^{\left|Y\right|\prime}$ are distinct, $\left|Y\right|\leq\left|Y^{\prime}\right|$. However,
$\left(f^{U}\left(p\right)\right)_{ao}=\frac{\left|Y\right|}{\left|\mathcal{Y}^{\ast}\left(p\right)\right|}\leq\frac{\left|Y^{\prime}\right|}{\left|\mathcal{Y}^{\ast}\left(p\right)\right|}=\left(f^{U}\left(p\right)\right)_{a^{\prime}o}$
contradicts $\left(f^{U}\left(p\right)\right)_{ao}>\left(f^{U}\left(p\right)\right)_{a^{\prime}o}$.
Proof of Lemma 3
Suppose that (3) is satisfied for $k$ defined in this lemma. Then, since $q_{\varnothing}\geq\left|\mathcal{A}\right|$, $r_{a}^{k^{\prime}}=p_{a}^{k^{\prime}}\neq\varnothing$ for all $k^{\prime}=1,\cdots,k-1$. $1\leq k<\left|\mathcal{O}\right|$.
When $k=1$, we let $R(\varnothing,p_{a^{\prime}})=1$ for all $a^{\prime}\in\mathcal{A\setminus}\left\{a\right\}$. Then,
$\left(f\left(r_{a},p_{-a}\right)\right)_{ar_{a}^{1}}=\left(f\left(p_{a},p_{-a}\right)\right)_{ap_{a}^{1}}=1,$
where $R(r_{a}^{1},r_{a})=1<R(p_{a}^{1},r_{a})$. Thus, $a$ with $r_{a}$ strictly prefers $f\left(r_{a},p_{-a}\right)$ to $f\left(p_{a},p_{-a}\right)$.
Next, we assume $k=2$. Then, $p_{a}^{1}=r_{a}^{1}$ and $p_{a}^{2}\neq r_{a}^{2}$. We assume that under $\left(r_{a},p_{-a}\right)$, the agents are divided into two groups. For the first group, there are $q_{r_{a}^{1}}+1$ agents whose preference is equivalent to $r_{a}$ under $\left(r_{a},p_{-a}\right)$; that is, if $a^{\prime}$ belongs to this group, then $r_{a}=p_{a^{\prime}}$. Note that $a$ itself belongs to this group. For the second group, the best object type is $\varnothing$ under $\left(r_{a},p_{-a}\right)$. Then, since all agents in the second group always assign $\varnothing$ with probability $1$ and $f$ satisfies ETE,
$\left(f\left(r_{a},p_{-a}\right)\right)_{ar_{a}^{1}}$ $=$ $\frac{q_{r_{a}^{1}}}{q_{r_{a}^{1}}+1},$
$\left(f\left(r_{a},p_{-a}\right)\right)_{ar_{a}^{2}}$ $=$ $\frac{1}{q_{r_{a}^{1}}+1}.$
On the other hand,
$\left(f\left(p_{a},p_{-a}\right)\right)_{ap_{a}^{1}}$ $=$ $\frac{q_{r_{a}^{1}}}{q_{r_{a}^{1}}+1},$
$\left(f\left(p_{a},p_{-a}\right)\right)_{ap_{a}^{2}}$ $=$ $\frac{1}{q_{r_{a}^{1}}+1}.$
Since $p_{a}^{1}=r_{a}^{1}$ and $p_{a}^{2}\neq r_{a}^{2}$, $a$ prefers $r_{a}^{2}$ to $p_{a}^{2}$. Therefore, $a$ strictly prefers $f\left(r_{a},p_{-a}\right)$ to $f\left(p_{a},p_{-a}\right)$.
We consider an arbitrary integer $k\geq 3$. Then, $r_{a}^{k^{\prime}}=p_{a}^{k^{\prime}}$ for all $k^{\prime}=1,\cdots,k-1$. We assume that under $\left(r_{a},p_{-a}\right)$ the agents are divided into $k$ groups. For the first group, there are $q_{r_{a}^{1}}+1$ agents whose preference is equivalent to $r_{a}$ under $\left(r_{a},p_{-a}\right)$; that is, if $a^{\prime}$ is in this group, then $r_{a}=p_{a^{\prime}}$. Note that $a$ belongs to this group. Next, we consider $j$th group for each $j=2,3,\cdots,k-1$. In $j$th group, there are just $q_{r_{a}^{j}}$ agents whose best object type is $r_{a}^{j}$ and the second best object is $r_{a}^{j+1},$ the third best $r_{a}^{j+2},$cdots and the $k$th best $r_{a}^{k}$ under $\left(r_{a},p_{-a}\right)$. If $a_{j}$ belongs $j$th group for $j=1,2,\cdots,k-1$, the preference orders of them are as follows:
$p_{a_{1}}$ $=$ $\left(r_{a}^{1},r_{a}^{2},\ldots,r_{a}^{k-2},r_{a}^{k-1},r_{a}^{k},\ldots\right),$ (10)
$p_{a_{2}}$ $=$ $\left(r_{a}^{2},r_{a}^{3},\ldots,r_{a}^{k-1},r_{a}^{1},r_{a}^{k},,\ldots\right),$
$\vdots$
$p_{a_{j}}$ $=$ $\left(r_{a}^{j},r_{a}^{j+1},\ldots,r_{a}^{j-2},r_{a}^{j-1},r_{a}^{k},\ldots\right),$
$\vdots$
$p_{a_{k-1}}$ $=$ $\left(r_{a}^{k-1},r_{a}^{1},\ldots,r_{a}^{k-3},r_{a}^{k-2},r_{a}^{k},\ldots\right).$
Since
$q_{r_{a}^{1}}+q_{r_{a}^{2}}+\cdots+q_{r_{a}^{k-1}}=\mathop{\displaystyle\sum}\limits_{o:R(o,r_{a})<k}q_{o}<\left|\mathcal{A}\right|$
by (3), $\left(r_{a},p_{-a}\right)$ is well-defined so far. Finally, for all the remaining agents (belonging to the $k$th group), the best object type is $\varnothing$ under $\left(r_{a},p_{-a}\right)$.
In this case, at least one agent cannot be assigned to their best object. Thus, in a rank-minimizing assignment, all agents in $j$th group for all $j=2,3,\cdots,k-1$ is assigned to their best object type $r_{a}^{j}$, and $q_{r_{a}^{1}}$ agents in the first group is also assigned to their best object type $r_{a}^{1},$ and moreover, just one agent in the first group is also assigned to $r_{a}^{k}$. Moreover, an agent who belongs to the $k$th group is assigned to their best object type $\varnothing$.
Since all agents excepting those belonging to the first group are assigned to their best alternative with probability $1$ and $f^{U}$ satisfies ETE,
$\left(f\left(r_{a},p_{-a}\right)\right)_{ar_{a}^{1}}$ $=$ $\frac{q_{r_{a}^{1}}}{q_{r_{a}^{1}}+1},$
$\left(f\left(r_{a},p_{-a}\right)\right)_{ar_{a}^{k}}$ $=$ $\frac{1}{q_{r_{a}^{1}}+1}.$
On the other hand,
$\left(f\left(p_{a},p_{-a}\right)\right)_{ap_{a}^{1}}$ $=$ $\frac{q_{r_{a}^{1}}}{q_{r_{a}^{1}}+1},$
$\left(f\left(p_{a},p_{-a}\right)\right)_{ap_{a}^{k}}$ $=$ $\frac{1}{q_{r_{a}^{1}}+1}.$
Since $p_{a}^{k^{\prime}}=r_{a}^{k^{\prime}}$ for all $k^{\prime}=1,\cdots,k-1$ and $p_{a}^{k}\neq r_{a}^{k}$, $a$ prefers $r_{a}^{k}$ to $p_{a}^{k}$ under $r_{a}$. Therefore, $a$ strictly prefers $f^{U}\left(r_{a},p_{-a}\right)$ to $f^{U}\left(p_{a},p_{-a}\right)$ under $r_{a}$. Thus, we have the first result.
We show the second result. Suppose that (2) is satisfied for $p_{a}$ and $k$ defined in the lemma that is,
$\mathop{\displaystyle\sum}\limits_{o:R(o,p_{a})<k}q_{o}=\mathop{\displaystyle\sum}\limits_{o:R(o,r_{a})<k}q_{o}\geq\left|\mathcal{A}\right|.$
By Lemma 1 and Remark 2, $f^{U}\left(p\right)_{ao}=f^{U}\left(r_{a},p_{-a}\right)_{ao}=0$ for all $o$ such that $R(o,p_{a})\geq k$. Therefore, for any $y^{\ast}\in\mathcal{Y}^{\ast}\left(r_{a},p_{-a}\right)$, $y_{ao}^{\ast}=0$; that is, for any assignment $x$ such that $x_{ao}>0$, $RV\left(y^{\ast},p\right)<RV\left(x,p\right)$. Therefore, $\mathcal{Y}^{\ast}\left(r_{a},p_{-a}\right)=\mathcal{Y}^{\ast}\left(p\right)$. Since $f^{U}$ is used, we have the second result.
Proof of Proposition 2
Let $f$ be a fair mechanism, $p_{a}\neq r_{a}$ and $k$ be the smallest integer such that $p_{ak}\neq r_{ak}$, which is also defined in Lemma 2. First, suppose (3) holds for that $k$. By the first result of Lemma 2, there is $p_{-a}\in\mathcal{P}^{\left|\mathcal{A}\right|-1}$ such that $a$ strictly prefers $f\left(r_{a},p_{-a}\right)$ to $f\left(p_{a},p_{-a}\right)$. Therefore, in this case, $r_{a}$ is not strategically dominate by $p_{a}$.
Second, suppose (3) does not hold for that $k$. In this case, since $r_{a}$ and $p_{a}$ are essentially identical and $f$ satisfies ETE, $f\left(r_{a},p_{-a}\right)=f\left(p_{a},p_{-a}\right)$. Therefore, $p_{a}$ does not strategically dominate $r_{a}$.
Proof of Theorem 1
Fix any $\left(r_{a},p_{-a}\right)\in\mathcal{P}^{\left|\mathcal{A}\right|}$ and any $d_{a}\in\mathcal{D}_{a}$. We show that $a$ weakly prefers $g\left(f^{U}\left(d_{a},p_{-a}\right),r\right)$ to $g\left(f^{U}\left(r_{a},p_{-a}\right),r\right)$. First, suppose that there is no $y\in\mathcal{Y}^{\ast}\left(r_{a},p_{-a}\right)$ such that $y_{ao}=1$ for some $o\in\mathcal{O}$ with $R\left(o,r_{a}\right)<R\left(\varnothing,r_{a}\right)$. Then, $g\left(f^{U}\left(r_{a},p_{-a}\right),r\right)_{a\varnothing}=1.$ Thus, for any assignment $x$, $a$ weakly prefers $g\left(x,r\right)$ to $g\left(f^{U}\left(r_{a},p_{-a}\right),r\right)$.
Second, suppose that there is some $y\in\mathcal{Y}^{\ast}\left(r_{a},p_{-a}\right)$ such that $y_{ao}=1$ with $R\left(o,r_{a}\right)<R\left(\varnothing,r_{a}\right)$. In this case, we have the following two results.
Claim 1
Let $y$ be such that $y_{ao^{\prime\prime}}=1$ where $R\left(o^{\prime\prime},r_{a}\right)<R\left(\varnothing,r_{a}\right)$. Then, $y\in\mathcal{Y}^{\ast}\left(r_{a},p_{-a}\right)$ if and only if $y\in\mathcal{Y}^{\ast}\left(d_{a},p_{-a}\right)$.
Proof. By the construction,
$RV\left(y,\left(r_{a},p_{-a}\right)\right)=RV\left(y,\left(d_{a},p_{-a}\right)\right).$
First, we show the only-if-part. Suppose $y\in\mathcal{Y}^{\ast}\left(r_{a},p_{-a}\right)$ such that $y_{ao^{\prime\prime}}=1$ where $R\left(o^{\prime\prime},r_{a}\right)<R\left(\varnothing,r_{a}\right)$. Toward a contradiction, suppose $y\notin\mathcal{Y}^{\ast}\left(d_{a},p_{-a}\right)$. We first consider the case where there is $y^{\prime}\in\mathcal{Y}^{\ast}\left(d_{a},p_{-a}\right)$ that satisfies $y_{ao^{\prime}}^{\prime}=1$ for $o^{\prime}$ such that $R\left(o^{\prime},r_{a}\right)<R\left(\varnothing,r_{a}\right)$. By the construction of $d_{a}$,
$RV\left(y^{\prime},\left(r_{a},p_{-a}\right)\right)$ $=$ $RV\left(y^{\prime},\left(d_{a},p_{-a}\right)\right)<$
$RV\left(y,\left(d_{a},p_{-a}\right)\right)$ $=$ $RV\left(y,\left(r_{a},p_{-a}\right)\right),$
which contradicts $y\in\mathcal{Y}^{\ast}\left(r_{a},p_{-a}\right)$.
We second consider the case where there is no such a deterministic assignment. Since $\mathcal{Y}^{\ast}\left(d_{a},p_{-a}\right)\neq\emptyset$ (by Corollary 1), there is $y^{\prime}\in\mathcal{Y}^{\ast}\left(d_{a},p_{-a}\right)$ such that $y_{ao^{\prime}}^{\prime}=1$ for $o^{\prime\prime}$ such that $R\left(o^{\prime\prime},r_{a}\right)\geq R\left(\varnothing,r_{a}\right)$. In this case,
$RV\left(y,\left(r_{a},p_{-a}\right)\right)=RV\left(y,\left(d_{a},p_{-a}\right)\right)>RV\left(y^{\prime},\left(d_{a},p_{-a}\right)\right).$
Now, we consider $y^{\prime\prime}$ such that $y_{a\varnothing}^{\prime\prime}=1,$ $y_{ao}^{\prime\prime}=0$ for all $o\in\mathcal{O}\setminus\left\{\varnothing\right\}$ and $y_{a^{\prime}o}^{\prime\prime}=y_{a^{\prime}o}$ for all $a^{\prime}\in\mathcal{A}\setminus\left\{a\right\}$ and $o\in\mathcal{O}$. Since $q_{\varnothing}\geq\left|\mathcal{A}\right|$, $y^{\prime\prime}$ is an assignment. Moreover, since $R\left(o^{\prime\prime},r_{a}\right)\geq R\left(\varnothing,r_{a}\right)$,
$RV\left(y^{\prime},\left(d_{a},p_{-a}\right)\right)\geq RV\left(y^{\prime\prime},\left(r_{a},p_{-a}\right)\right)\text{.}$
Then,
$RV\left(y^{\ast},\left(r_{a},p_{-a}\right)\right)>RV\left(y^{\prime\prime},\left(r_{a},p_{-a}\right)\right)$
contradicts $y\in\mathcal{Y}^{\ast}\left(r_{a},p_{-a}\right)$.
Second, we show the if-part. Suppose $y\in\mathcal{Y}^{\ast}\left(d_{a},p_{-a}\right)$. We show $y\in\mathcal{Y}^{\ast}\left(r_{a},p_{-a}\right)$. Suppose not; that is, there is $y^{\prime}\in\mathcal{Y}$ such that $RV\left(y^{\prime},\left(r_{a},p_{-a}\right)\right)<RV\left(y,\left(r_{a},p_{-a}\right)\right)$. Then, $y_{a\bar{o}}^{\prime}=1$ for $R\left(\bar{o},r_{a}\right)\leq R\left(\varnothing,r_{a}\right)$.
First, we assume $R\left(\bar{o},r_{a}\right)<R\left(\varnothing,r_{a}\right)$. Then, $RV\left(y^{\prime},\left(r_{a},p_{-a}\right)\right)=RV\left(y^{\prime},\left(d_{a},p_{-a}\right)\right)$. However,
$RV\left(y^{\prime},\left(d_{a},p_{-a}\right)\right)=RV\left(y^{\prime},\left(r_{a},p_{-a}\right)\right)<RV\left(y,\left(r_{a},p_{-a}\right)\right)=RV\left(y,\left(d_{a},p_{-a}\right)\right),$
contradicting $y\in\mathcal{Y}^{\ast}\left(d_{a},p_{-a}\right)$.
Second, we assume $\bar{o}=\varnothing$. Then, since
$\mathop{\textstyle\sum}\limits_{\bar{o}:R\left(\bar{o},r_{a}\right)<R\left(\varnothing,r_{a}\right)}q_{\bar{o}}+q_{o^{\prime}}\geq\left|\mathcal{A}\right|,$ (11)
there is $\hat{o}$ such that $R\left(\hat{o},d_{a}\right)\leq R\left(o^{\prime},d_{a}\right)$. Thus, for $y^{\prime\prime}\in\mathcal{Y}$ such that $y_{a\hat{o}}^{\prime\prime}=1$ and $y_{a^{\prime}\tilde{o}}^{\prime\prime}=y_{a^{\prime}\tilde{o}}^{\prime}$ for all $a\in\mathcal{A}$ and all $\tilde{o}\in\mathcal{O}\setminus\left\{\hat{o}\right\}$, $RV\left(y^{\prime\prime},\left(d_{a},p_{-a}\right)\right)\leq RV\left(y^{\prime},\left(r_{a},p_{-a}\right)\right)$. However,
$RV\left(y^{\prime\prime},\left(d_{a},p_{-a}\right)\right)\leq RV\left(y^{\prime},\left(r_{a},p_{-a}\right)\right)<RV\left(y,\left(r_{a},p_{-a}\right)\right)=RV\left(y,\left(d_{a},p_{-a}\right)\right),$
contradicting $y\in\mathcal{Y}^{\ast}\left(d_{a},p_{-a}\right)$.
Therefore, $y\in\mathcal{Y}^{\ast}\left(d_{a},p_{-a}\right)$, then $y\in\mathcal{Y}^{\ast}\left(r_{a},p_{-a}\right)$. Q.E.D.
Claim 2
Suppose that $y^{\ast}\in\mathcal{Y}^{\ast}\left(r_{a},p_{-a}\right)$ such that $y_{ao}^{\ast}=1$ with $R\left(o,r_{a}\right)<R\left(\varnothing,r_{a}\right)$. Then, $\left|\mathcal{Y}^{\ast}\left(d_{a},p_{-a}\right)\right|\leq\left|\mathcal{Y}^{\ast}\left(r_{a},p_{-a}\right)\right|$.
Proof. First, we show that if $y\in\mathcal{Y}^{\ast}\left(d_{a},p_{-a}\right)$, then either $y\in\mathcal{Y}^{\ast}\left(r_{a},p_{-a}\right)$ or $y^{\prime}\in\mathcal{Y}^{\ast}\left(r_{a},p_{-a}\right),$ such that $y_{a\varnothing}^{\prime}=1$ and $y_{a^{\prime}o}^{\prime}=y_{a^{\prime}o}$ for all $a^{\prime}\in\mathcal{A}\setminus\left\{a\right\}$ and $o\in\mathcal{O}$.
Suppose that $y\in\mathcal{Y}^{\ast}\left(d_{a},p_{-a}\right)$ and $y\notin\mathcal{Y}^{\ast}\left(r_{a},p_{-a}\right)$. We have
$RV\left(y,\left(d_{a},p_{-a}\right)\right)\leq RV\left(y^{\ast},\left(d_{a},p_{-a}\right)\right)=RV\left(y^{\ast},\left(r_{a},p_{-a}\right)\right)<RV\left(y,\left(r_{a},p_{-a}\right)\right).$
Then, by the construction of $d_{a}$, $y_{ao^{\prime}}=1\,$such that $R\left(o^{\prime},r_{a}\right)\geq R\left(\varnothing,r_{a}\right)$. Since $q_{\varnothing}\geq\left|\mathcal{A}\right|$, $y^{\prime}$ is an assignment such that $y_{a\varnothing}^{\prime}=1,$ $y_{ao}^{\prime}=0$ for all $o\in\mathcal{O}\setminus\left\{\varnothing\right\}$ and $y_{a^{\prime}o}^{\prime}=y_{a^{\prime}o}$ for all $a^{\prime}\in\mathcal{A}\setminus\left\{a\right\}$ and $o\in\mathcal{O}$. Since $R\left(o^{\prime},d_{a}\right)\geq R\left(\varnothing,r_{a}\right)$ (by the construction of $d_{a}$), we have
$RV\left(y^{\prime},\left(r_{a},p_{-a}\right)\right)\leq RV\left(y,\left(d_{a},p_{-a}\right)\right)\leq RV\left(y^{\ast},\left(d_{a},p_{-a}\right)\right)\text{.}$
Since $y^{\ast}\in\mathcal{Y}^{\ast}\left(r_{a},p_{-a}\right)$, $y^{\prime}\in\mathcal{Y}^{\ast}\left(r_{a},p_{-a}\right)$. Therefore, if $y\in\mathcal{Y}^{\ast}\left(d_{a},p_{-a}\right)$, then either $y\in\mathcal{Y}^{\ast}\left(r_{a},p_{-a}\right)$ or $y^{\prime}\in\mathcal{Y}^{\ast}\left(r_{a},p_{-a}\right)$ defined above.
Finally, we assume $\left|\mathcal{Y}^{\ast}\left(d_{a},p_{-a}\right)\right|\geq 2$ and arbitrarily let $y_{1},y_{2}\in\mathcal{Y}^{\ast}\left(d_{a},p_{-a}\right)$ such that $y_{1}\neq y_{2}$. For $i=1,2$, let $y_{i}^{\prime}$ be such that $\left(y_{i}^{\prime}\right)_{a\varnothing}=1$ and $\left(y_{i}^{\prime}\right)_{a^{\prime}o}=\left(y_{i}\right)_{a^{\prime}o}$ for all $a^{\prime}\in\mathcal{A}\setminus\left\{a\right\}$ and $o\in\mathcal{O}$. We show $y_{1}^{\prime}\neq y_{2}^{\prime}$. Suppose not; that is, $y_{1}^{\prime}=y_{2}^{\prime}$. Then, $\left(y_{1}\right)_{a^{\prime}o}=\left(y_{2}\right)_{a^{\prime}o}$ for all $a^{\prime}\in\mathcal{A}\setminus\left\{a\right\}$ and $o\in\mathcal{O}$. Since $y_{1}\neq y_{2}$, we have $o_{1}\neq o_{2}$ where $\left(y_{1}\right)_{ao_{1}}=1$ and $\left(y_{2}\right)_{ao_{2}}=1$. Without loss of generality, we assume $R\left(o_{1},d_{a}\right)<R\left(o_{2},d_{a}\right)$. However, $R\left(o_{1},d_{a}\right)<R\left(o_{2},d_{a}\right)$ and $\left(y_{1}\right)_{a^{\prime}o}=\left(y_{2}\right)_{a^{\prime}o}$ for all $a^{\prime}\in\mathcal{A}\setminus\left\{a\right\}$ and $o\in\mathcal{O}$ imply
$RV\left(y_{1},\left(d_{a},p_{-a}\right)\right)<RV\left(y_{2},\left(d_{a},p_{-a}\right)\right)$
contradicting that $y_{2}\in\mathcal{Y}^{\ast}\left(d_{a},p_{-a}\right)$. Therefore, $y_{1}^{\prime}\neq y_{2}^{\prime}$.
By these results, if $\left\{y_{1},y_{2},\cdots,y_{n}\right\}=\mathcal{Y}^{\ast}\left(d_{a},p_{-a}\right),$ then $\left\{y_{1}^{\prime\prime},y_{2}^{\prime\prime},\cdots,y_{n}^{\prime\prime}\right\}\subseteq\mathcal{Y}^{\ast}\left(r_{a},p_{-a}\right)$ where $y_{i}^{\prime\prime}$ is either $y_{i}$ or $y_{i}^{\prime}$ constructed above, because $y_{1}^{\prime},y_{2}^{\prime},\cdots,y_{n}^{\prime}$ are distinct. Q.E.D.
Now, we show that $a$ weakly prefers $g\left(f^{U}\left(d_{a},p_{-a}\right),r\right)$ to $g\left(f^{U}\left(r_{a},p_{-a}\right),r\right)$. By Claim 2,
$\left|\mathcal{Y}^{\ast}\left(d_{a},p_{-a}\right)\right|\geq\left|\mathcal{Y}^{\ast}\left(r_{a},p_{-a}\right)\right|\text{.}$
Moreover, any $y^{\ast}\in\mathcal{Y}^{\ast}\left(r_{a},p_{-a}\right)$ such that $y_{ao}^{\ast}=1$ with $R\left(o,r_{a}\right)<R\left(\varnothing,r_{a}\right)$ is also an element of $\mathcal{Y}^{\ast}\left(d_{a},p_{-a}\right)$.
Thus, for any $o\in\mathcal{O}$ such that $R\left(o,r_{a}\right)<R\left(\varnothing,r_{a}\right)$,
$\left(g\left(f^{U}\left(d_{a},p_{-a}\right),r\right)\right)_{ao}$ $=$ $\left(f^{U}\left(d_{a},p_{-a}\right)\right)_{ao}\geq$
$\left(f^{U}\left(r_{a},p_{-a}\right)\right)_{ao}$ $=$ $\left(g\left(f^{U}\left(r_{a},p_{-a}\right),r\right)\right)_{ao}\text{.}$
Since $g\left(f^{U}\left(d_{a},p_{-a}\right)\right)_{ao^{\prime}}=0$ for all $R\left(o^{\prime},r_{a}\right)>R\left(\varnothing,r_{a}\right)$,
$\left(g\left(f^{U}\left(d_{a},p_{-a}\right),r\right)\right)_{a\varnothing}\leq\left(g\left(f^{U}\left(r_{a},p_{-a}\right),r\right)\right)_{a\varnothing}.$
Thus, $a$ weakly prefers $g\left(f\left(d_{a},p_{-a}\right),r\right)$ to $g\left(f\left(r_{a},p_{-a}\right),r\right)$.
Proof of Theorem 2
[…]
First, we show that for all $y\in\mathcal{Y}^{\ast}\left(r_{a},p_{-a}\right)$, $y_{a\varnothing}=0$. Suppose not; that is, there is $y\in\mathcal{Y}^{\ast}\left(r_{a},p_{-a}\right)$, $y_{a\hat{o}}=1$ where $R\left(\hat{o},r_{a}\right)<R\left(\varnothing,r_{a}\right)$. Then, by (5), there is $a^{\prime}\in\mathcal{A}$ such that $y_{a^{\prime}o^{\prime\prime}}=1$ with $R\left(o^{\prime\prime},p_{a^{\prime}}\right)>R\left(o^{\prime},p_{a^{\prime}}\right)\left(=R\left(\varnothing,r_{a}\right)\right),$ because $p_{a^{\prime}}=d_{a}$ for all $a^{\prime}\in\mathcal{A\setminus}\left\{a\right\}$. Now, let $y^{\prime}\in\mathcal{Y}$ such that $y_{a^{\prime}\hat{o}}^{\prime}=1,$ $y_{a\varnothing}^{\prime}=1,$ and $y_{a^{\prime\prime}\bar{o}}^{\prime}=y_{a^{\prime\prime}\bar{o}}$ for all $a^{\prime\prime}\in\mathcal{A\setminus}\left\{a,a^{\prime}\right\}$ and all $\bar{o}\in\mathcal{O}$. Then, since
$R\left(\hat{o},r_{a^{\prime}}\right)=R\left(\hat{o},p_{a^{\prime}}\right)\text{ and }R\left(o^{\prime\prime},p_{a^{\prime}}\right)>R\left(\varnothing,r_{a}\right),$
$RV\left(y^{\prime},\left(r_{a},p_{-a}\right)\right)<RV\left(y,\left(r_{a},p_{-a}\right)\right),$
contradicting $y\in\mathcal{Y}^{\ast}\left(r_{a},p_{-a}\right)$. Therefore, for all $y\in\mathcal{Y}^{\ast}\left(r_{a},p_{-a}\right)$, $y_{a\varnothing}=0$.
By this result,
$\left(f^{U}\left(r_{a},p_{-a}\right),r\right)_{a\varnothing}=\left(g\left(f^{U}\left(d_{a},p_{-a}\right),r\right)\right)_{a\varnothing}=1$
On the other hand, since $f^{U}$ satisfies ETE,
$\left(f^{U}\left(d_{a},p_{-a}\right),r\right)_{a\hat{o}}=\left(g\left(f^{U}\left(d_{a},p_{-a}\right),r\right)\right)_{a\hat{o}}>0.$
Therefore, $a$ strictly prefers $g\left(f^{U}\left(d_{a},p_{-a}\right),r\right)$ to $g\left(f^{U}\left(r_{a},p_{-a}\right),r\right)$.
Proof of Proposition 3
Fix $p_{-a}\in\mathcal{P}^{\left|\mathcal{A}\right|-1}$ arbitrarily. Let $r_{a}\in\mathcal{P}$ and $o,o^{\prime}\in\mathcal{O}$ satisfy the condition in Proposition 3. Moreover, we let $d_{a}\in\mathcal{D}_{a}$ be such that $R\left(\varnothing,r_{a}\right)=R\left(o^{\prime},d_{a}\right)$.
We show the following fact.
Claim 3
Let $y$ and $y^{\prime}$ be such that $y_{a\varnothing}=y_{ao^{\prime}}^{\prime}=1$ and $y_{a^{\prime}\bar{o}}=y_{a^{\prime}\bar{o}}^{\prime}$ for all $a\in\mathcal{A}$ and all $\bar{o}\in\mathcal{O}\setminus\left\{\varnothing,o^{\prime}\right\}$. Then, $y\in\mathcal{Y}^{\ast}\left(r_{a},p_{-a}\right)$ if and only if $y^{\prime}\in\mathcal{Y}^{\ast}\left(d_{a},p_{-a}\right)$.
Proof. By the construction,
$RV\left(y,\left(r_{a},p_{-a}\right)\right)=RV\left(y^{\prime},\left(d_{a},p_{-a}\right)\right)\text{.}$
First, suppose $y\in\mathcal{Y}^{\ast}\left(r_{a},p_{-a}\right)$. Suppose not; that is, there is $y^{\prime\prime}\in\mathcal{Y}$ such that $RV\left(y^{\prime\prime},\left(d_{a},p_{-a}\right)\right)<RV\left(y^{\prime},\left(d_{a},p_{-a}\right)\right)$. Then, since (11) and Lemma 2 are satisfied, $o\in\mathcal{O}$ such that $y_{ao}^{\prime\prime}=1$ and $R\left(o,d_{a}\right)\leq R\left(o^{\prime},d_{a}\right)$. Let
$o^{\prime\prime\prime}=o\text{ if }o\neq o^{\prime},\text{ }o^{\prime\prime\prime}=\varnothing\text{ if }o=o^{\prime}\text{.}$
Moreover, let $y^{\prime\prime\prime}$ be such that $y_{ao^{\prime\prime\prime}}^{\prime\prime\prime}=1$ and $y_{a^{\prime}o}^{\prime\prime\prime}=y_{a^{\prime}o}^{\prime\prime}$ for any other pairs $(a^{\prime},o)\neq\left(a,o^{\prime\prime\prime}\right)$. Then,
$RV\left(y^{\prime\prime\prime},\left(r_{a},p_{-a}\right)\right)\leq RV\left(y^{\prime\prime},\left(d_{a},p_{-a}\right)\right)<RV\left(y^{\prime},\left(d_{a},p_{-a}\right)\right)=RV\left(y,\left(r_{a},p_{-a}\right)\right)$
contradicting $y\in\mathcal{Y}^{\ast}\left(r_{a},p_{-a}\right)$.
Second, suppose $y^{\prime}\in\mathcal{Y}^{\ast}\left(d_{a},p_{-a}\right)$. Suppose not; that is, there is $y^{\prime\prime}\in\mathcal{Y}$ such that $RV\left(y^{\prime\prime},\left(r_{a},p_{-a}\right)\right)<RV\left(y,\left(r_{a},p_{-a}\right)\right)$. Then, since (11) and Lemma 2 are satisfied, $o\in\mathcal{O}$ such that $y_{ao}^{\prime\prime}=1$ and $R\left(o,r_{a}\right)\leq R\left(\varnothing,r_{a}\right)$. Let
$o^{\prime\prime\prime}=o\text{ if }o\neq\varnothing,\text{ }o^{\prime\prime\prime}=\bar{o}\text{ if }o=\varnothing\text{,}$
where $\bar{o}$ is such that $\sum y_{a\bar{o}}^{\prime\prime}<q_{\bar{o}}$ and $R\left(\bar{o},d_{a}\right)\leq R\left(o^{\prime},d_{a}\right)$. There exist such an $\bar{o}$, because (11) is satisfied. Then, let $y^{\prime\prime\prime}$ be such that $y_{ao^{\prime\prime\prime}}^{\prime\prime\prime}=1$ and $y_{a^{\prime}o}^{\prime\prime\prime}=y_{a^{\prime}o}^{\prime\prime}$ for any other pairs $(a^{\prime},o)\neq\left(a,o^{\prime\prime\prime}\right)$. Then,
$RV\left(y^{\prime\prime\prime},\left(d_{a},p_{-a}\right)\right)\leq RV\left(y^{\prime\prime},\left(r_{a},p_{-a}\right)\right)<RV\left(y^{\prime},\left(r_{a},p_{-a}\right)\right)=RV\left(y,\left(d_{a},p_{-a}\right)\right)$
contradicting $y\in\mathcal{Y}^{\ast}\left(d_{a},p_{-a}\right)$.
Therefore, $y\in\mathcal{Y}^{\ast}\left(r_{a},p_{-a}\right)$ if and only if $y^{\prime}\in\mathcal{Y}^{\ast}\left(d_{a},p_{-a}\right)$. Q.E.D.
By Claims 1 and 3, we have Proposition 3.
Proof of Lemma 4
We assume that the assumptions of Lemma 4 is satisfied. First, by Proposition 3, for all $p_{-a}\in\mathcal{P}^{\left|\mathcal{A}\right|-1}$ and all $\bar{o}\in\mathcal{O}$,
$g\left(f^{U}\left(e_{a},p_{-a}\right),r\right)_{a\bar{o}}=g\left(f^{U}\left(r_{a},p_{-a}\right),r\right)_{a\bar{o}}.$
Second, by Theorem 2, $r_{a}$ is strategically dominated by $d_{a}$ with the refusal option. Therefore, $e_{a}$ is also strategically dominated by $d_{a}$ with the refusal option.
Proof of Proposition 4
We use the example introduced in the proof of Theorem 2. Suppose that there are $a\in\mathcal{A}$, $o,o^{\prime}\in\mathcal{O}$ satisfying (5). Let $d_{a}$ be such that $R\left(\varnothing,r_{a}\right)=R\left(o^{\prime},d_{a}\right)$ and $p_{a^{\prime}}=d_{a}$ for all $a^{\prime}\in\mathcal{A\setminus}\left\{a\right\}$. Then, since (5) is satisfied and $f^{U}$ satisfies ETE, there is $y\in\mathcal{Y}^{\ast}\left(d_{a},p_{-a}\right)$, $y_{ao^{\prime}}=1$ and $y_{a^{\prime}o^{\prime\prime}}=1$ for some $a^{\prime}\in\mathcal{A\setminus}\left\{a\right\}$ and $o^{\prime\prime}\in\mathcal{O}$ such that $R\left(o^{\prime},d_{a}\right)<R\left(o^{\prime\prime},d_{a}\right)$. Then, if $y$ is realized, $a$ refuses $o^{\prime}$ but $a^{\prime}$ prefers $o^{\prime}$ to $o^{\prime\prime}$. Therefore, $g\left(f^{U}\left(d_{a},p_{-a}\right),r\right)$ is wasteful.
Proof of Proposition 5
First, suppose that $p$ satisfies all of Conditions (i), (ii) and (iii). Then, for any two agents $\hat{a}$ and $\bar{a}$ belonging to different groups, $p_{\hat{a}}$ and $p_{\bar{a}}$ are not essentially the same. Moreover, for any two agents $\hat{a}$ and $\bar{a}$ belonging to the second group or third group, $\left(f^{M}\left(p\right)\right)_{\hat{a}o^{\prime}}=\left(f^{M}\left(p\right)\right)_{\bar{a}o^{\prime}}$ for all $o^{\prime}\in\mathcal{O}$.
Second if $p$ does not satisfies all of Conditions (i), (ii) and (iii), then $f^{M}\left(p\right)=f^{U}\left(p\right)$. By Proposition 2, $f^{M}$ satisfies ETE.
Proof of Proposition 6
First, if $r_{a}^{1}=\varnothing,$ then $\left(f^{M}\left(r_{a},p_{-a}\right)\right)_{a\varnothing}=1$ for all $p_{-a}\in\mathcal{P}^{\left|\mathcal{A}\right|-1}$. Thus, in this case, there is no $p_{a}\in\mathcal{P}$ that strategically dominate $r_{a}$. Thus, we assume $r_{a}^{1}\neq\varnothing$. We consider $p_{a}\left(\neq r_{a}\right)$. Let $k$ be the smallest integer such that $p_{a}^{k}\neq r_{a}^{k}$.
First, suppose that (3) is satisfied for that $k$ and $r_{a}^{k}\neq\varnothing$. Since (3) is satisfied for $k$, $r_{a}^{k^{\prime}}\neq\varnothing$ for all $k^{\prime}=1,\cdots,k-1$; that is, $R(\varnothing,r_{a})>k$ and $R(\varnothing,p_{a})\geq k$. In this case, since $f^{M}$ satisfies ETE (by Proposition 5), we revisit the examples in the first result of Lemma 3. When $k=1$ or $2$, trivially, we can use the example as is, because
$\left(g\left(f^{M}\left(r_{a},p_{-a}\right),r\right)\right)_{ao}$ $=$ $\left(f^{M}\left(r_{a},p_{-a}\right),r\right)_{ao},$
$\left(g\left(f^{M}\left(p_{a},p_{-a}\right),r\right)\right)_{ao}$ $=$ $\left(f^{M}\left(p_{a},p_{-a}\right),r\right)_{ao}.$
for all $o\in\mathcal{O}$.
We consider $k=3.$ Then, since $R(\varnothing,r_{a})>k$ and $R(\varnothing,p_{a})\geq k$,
$\left(g\left(f^{M}\left(r_{a},p_{-a}\right),r\right)\right)_{ar_{a}^{1}}$ $=$ $\frac{q_{r_{a}^{1}}}{q_{r_{a}^{1}}+1},\text{ }\left(g\left(f^{M}\left(r_{a},p_{-a}\right),r\right)\right)_{ar_{a}^{k}}=\frac{1}{q_{r_{a}^{1}}+1}.$
$\left(g\left(f^{M}\left(r_{a},p_{-a}\right),r\right)\right)_{ap_{a}^{1}}$ $=$ $\frac{q_{r_{a}^{1}}}{q_{r_{a}^{1}}+1},$
and moreover, either
$\left(g\left(f^{M}\left(p_{a},p_{-a}\right),r\right)\right)_{ap_{a}^{k}}$ $=$ $\frac{1}{q_{r_{a}^{1}}+1}\text{ or}$
$\left(g\left(f^{M}\left(p_{a},p_{-a}\right),r\right)\right)_{a\varnothing}$ $=$ $\frac{1}{q_{r_{a}^{1}}+1}\text{.}$
Since $R(\varnothing,r_{a})>k$ and $R(p_{a}^{k},r_{a})>k$, in either case, $a$ prefers $g\left(f^{M}\left(r_{a},p_{-a}\right),r\right)$ to $g\left(f^{M}\left(p_{a},p_{-a}\right),r\right)$.
Second, suppose (3) is satisfied and $r_{a}^{k}=\varnothing$. Then, $k\geq 2$. Moreover, in this case, $p_{a}^{k^{\prime}}\neq\varnothing$ for all $k^{\prime}=1,\cdots,k$. Therefore, $R(\varnothing,p_{a})>R(\varnothing,r_{a})=k$. Let $p_{-a}$ be such that there are $q_{o}$ students represented by $a^{\prime}$ whose revealed preference is equivalent to $r_{a}(=p_{a^{\prime}})$ and the first best type of the other agents represented by $a^{\prime\prime}$ is $\varnothing$. Then, $\left(p_{a},p_{-a}\right)$ satisfies (i), (ii) and (iii) with $l=k$, because $k$ satisfies (3). Moreover, $\left(r_{a},p_{-a}\right)$ does not satisfy all of them. Therefore, in this case, $g\left(f^{M}\left(r_{a},p_{-a}\right)\right)$ is more preferable for $a$ than $g\left(f^{M}\left(p_{a},p_{-a}\right)\right)$.
Third, suppose that $k$ does not satisfy (3); that is,
$\mathop{\displaystyle\sum}\limits_{\bar{o}:R(\bar{o},p_{a})<k}q_{\bar{o}}=\mathop{\displaystyle\sum}\limits_{\bar{o}:R(\bar{o},r_{a})<k}q_{\bar{o}}\geq\left|\mathcal{A}\right|$
In this case, $r_{a}$ and $p_{a}$ are essentially the same. Since $f^{M}$ is a fair rank-minimizing mechanism (by Proposition 5),
$g\left(f^{M}\left(r_{a},p_{-a}\right),r\right)=g\left(f^{M}\left(p_{a},p_{-a}\right),r\right)$
for all $p_{-a}\in\mathcal{P}^{\left|\mathcal{A}\right|-1}$.
Therefore, for any $r_{a}\in\mathcal{P}$, there is no $p_{a}\in\mathcal{P}$ that strategically dominate $r_{a}$.

### v3

@@
 Let $\mathcal{A}$ and $\mathcal{O}$ be finite sets of agents and object types respectively.
+We assume $\left|\mathcal{A}\right|\geq 2$ and $\left|\mathcal{O}\right|\geq 2$.
 Each agent must be allocated one copy of an object whose type is an element of a set denoted by $\mathcal{O}$.
@@
 In Section 4, we consider the situation where the null object type has the other characteristic.
-To avoid trivial cases, we assume $q_{o}<\left|\mathcal{A}\right|$ for all $o\in\mathcal{O\setminus}\left\{\varnothing\right\}$.
+To avoid trivial cases, we assume $q_{o}\in\left[1,\left|\mathcal{A}\right|\right)$ for all $o\in\mathcal{O\setminus}\left\{\varnothing\right\}$.
 Let $\mathcal{P}$ be the set of all possible strict rankings (linear orders) of the elements in $\mathcal{O}$.
-Let $p_{a}=\left(p_{a}^{1},\cdots,p_{a}^{\left|\mathcal{O}\right|}\right)\in\mathcal{P}$ be a preference order revealed by $a$, where $p_{a}^{k}=o$ means that type $o$ object is ranked $k$th best one under $p_{a}$.
+Let $p_{a}=\left(p_{a}^{1},\cdots,p_{a}^{\left|\mathcal{O}\right|}\right)\in\mathcal{P}^{\left|\mathcal{A}\right|}$ be a preference order revealed by $a$, where $p_{a}^{k}=o$ means that type $o$ object is ranked $k$th best one under $p_{a}$.
 For notational convenience, we let $R:\mathcal{O\times P}\rightarrow\left\{1,\cdots,\left|\mathcal{O}\right|\right\}$ be a rank function; that is, $p_{a}^{k}=o$ if and only if $R\left(o,p_{a}\right)=k$.
@@
 unacceptable) for $p_{a}$.
-Thus, note that $R\left(\varnothing,p_{a}\right)-1$ is the number of acceptable object types for $p_{a}$.
+Thus, $R\left(\varnothing,p_{a}\right)-1$ represents the number of acceptable object types for $p_{a}$.
 Let $r_{a}\in\mathcal{P}$ be the true preference order of agent $a$.
@@
 We say that a matrix $x$ is an assignment if
-$x_{ao}\in\left[0,1\right]\text{, }\mathop{\displaystyle\sum}\limits_{o^{\prime}\in\mathcal{O}}x_{ao^{\prime}}=1\text{,\ and }\mathop{\displaystyle\sum}\limits_{a^{\prime}\in\mathcal{A}}x_{a^{\prime}o}\leq q_{o},$
+$x_{ao}\in\left[0,1\right]\text{, }\sum\limits_{o^{\prime}\in\mathcal{O}}x_{ao^{\prime}}=1\text{,\ and }\sum\limits_{a^{\prime}\in\mathcal{A}}x_{a^{\prime}o}\leq q_{o},$
 for all $a\in\mathcal{A}$ and all $o\in\mathcal{O}$.
@@
 We consider the preference of agent $a$ with $p_{a}$ with regard to two assignments $x$ and $x^{\prime}$. If
-$\mathop{\displaystyle\sum}\limits_{o^{\prime}:\text{ }R\left(o^{\prime},p_{a}\right)\leq k}x_{ao^{\prime}}\geq\mathop{\displaystyle\sum}\limits_{o^{\prime}:\text{ }R\left(o^{\prime},p_{a}\right)\leq k}x_{ao^{\prime}}^{\prime}$ (1)
+$\sum\limits_{o^{\prime}:\text{ }R\left(o^{\prime},p_{a}\right)\leq k}x_{ao^{\prime}}\geq\sum\limits_{o^{\prime}:\text{ }R\left(o^{\prime},p_{a}\right)\leq k}x_{ao^{\prime}}^{\prime}$ (1)
 for all $k=1,2,\cdots,\left|\mathcal{O}\right|$, then agent $a$ with $p_{a}$ is said to weakly prefer (in the sense of first-order stochastic dominance) $x$ to $x^{\prime}$.
-Moreover, $a$ is said to strictly prefers $x$ to $x^{\prime}$ if $a$ weakly prefers $x$ to $x^{\prime}$ and (1) is satisfied with inequality for some $k=1,2,\cdots,\left|\mathcal{O}\right|$.
+Moreover, $a$ is said to strictly prefers $x$ to $x^{\prime}$ if $a$ weakly prefers $x$ to $x^{\prime}$ and (1) is satisfied with inequality for some $k=1,2,\cdots,\left|\mathcal{O}\right|-1$.
 Assignment $x$ weakly ordinally dominates assignment $x^{\prime}$ for $p$ if each agent $a$ with $p_{a}$ weakly prefers $x$ to $x^{\prime}$.
@@
 An assignment that is not ordinally dominated by any other assignments is said to be ordinally efficient for $p$.
-Assignment $x$ is wasteful for $p$ if there are $o,o^{\prime}\in\mathcal{O}$ and $a\in\mathcal{A}$ such that $\mathop{\textstyle\sum}\nolimits_{a^{\prime}\in\mathcal{A}}x_{a^{\prime}o}<q_{o}$, $x_{ao^{\prime}}>0$ and $R\left(o,p_{a}\right)<R\left(o^{\prime},p_{a}\right)$.
+Assignment $x$ is wasteful for $p$ if there are $o,o^{\prime}\in\mathcal{O}$ and $a\in\mathcal{A}$ such that $\sum\nolimits_{a^{\prime}\in\mathcal{A}}x_{a^{\prime}o}<q_{o}$, $x_{ao^{\prime}}>0$ and $R\left(o,p_{a}\right)<R\left(o^{\prime},p_{a}\right)$.
 Any wasteful assignment for $p$ must not be ordinally efficient.
@@
 For an assignment $x\in\mathcal{X}$ and a preference profile $p=\left(p_{a}\right)_{a\in\mathcal{A}}\in\mathcal{P}^{\left|\mathcal{A}\right|}$, let $RV\left(x,p\right)$ be such that
-$RV\left(x,p\right)=\mathop{\displaystyle\sum}\limits_{a\in\mathcal{A}}\mathop{\displaystyle\sum}\limits_{o\in\mathcal{O}}R\left(o,p_{a}\right)\times x_{ao},$
+$RV\left(x,p\right)=\sum\limits_{a\in\mathcal{A}}\sum\limits_{o\in\mathcal{O}}R\left(o,p_{a}\right)\times x_{ao},$
 which is the sum of expected rank.
@@
 Fix any $x^{\ast}\in\mathcal{X\setminus Y}$. Let $y^{1},\cdots,y^{I}\in\mathcal{Y}$ with $I\geq 2$ be such that
-$x^{\ast}=\mathop{\displaystyle\sum}\limits_{i=1}^{I}\sigma^{i}y^{i}\in\mathcal{X}\text{ where }\mathop{\displaystyle\sum}\limits_{i=1}^{I}\sigma^{i}=1$
+$x^{\ast}=\sum\limits_{i=1}^{I}\sigma^{i}y^{i}\in\mathcal{X}\text{ where }\sum\limits_{i=1}^{I}\sigma^{i}=1$
 and $\sigma^{i}\in\left[0,1\right]$ for all $i=1,\cdots,I$. If $x^{\ast}$ is rank efficient, then $y^{i}$ is also rank efficient for all $i=1,\cdots,I$.
@@
 There must exist at least one deterministic rank efficient assignment.
-Nikzad (2022) and Troyan (2024) consider a deterministic rank efficient assignment; that is, they focus only on a deterministic assignment $y^{\ast}\in\mathcal{Y}$ such that $RV\left(y^{\ast},p\right)\leq RV\left(y,p\right)$ for all $y\in\mathcal{Y}$.
-However, by Corollary 1, such $y^{\ast}$ must also be rank efficient in our definition.
+Nikzad (2022) and Troyan (2024) consider a deterministic rank efficient assignment; that is, they focus on a deterministic assignment $y^{\ast}\in\mathcal{Y}$ such that $RV\left(y^{\ast},p\right)\leq RV\left(y,p\right)$ for all $y\in\mathcal{Y}$.
+However, by Corollary 1, such $y^{\ast}$ must also be rank efficient in our definition; that is, $RV\left(y^{\ast},p\right)\leq RV\left(x,p\right)$ for all $x\in\mathcal{X}$.
 Moreover, this result allows us to show that the mechanism introduced later is well-defined.
-Finally, we introduce another technical result that is used to show our main results introduced later.
+Finally, we introduce another technical result that is used to show our main results introduced later. Let
+$\bar{k}\left(p_{a}\right)=\min\left\{k\in\left\{1,\cdots,\left|\mathcal{O}\right|\right\}\text{ }\left|\text{ }\sum\limits_{\bar{o}:R(\bar{o},p_{a})\leq k}q_{\bar{o}}\geq\left|\mathcal{A}\right|\right.\right\},$
+which is well-defined because $\sum\nolimits_{\bar{o}\in\mathcal{O}}q_{\bar{o}}>\left|\mathcal{A}\right|$.
+This means that for a given $p_{a}$, there are sufficient number of copies of object types that ranked higher than or equal to $\bar{k}\left(p_{a}\right)$.
+Trivially, if $p_{a}^{k}=\varnothing$, then $\bar{k}\left(p_{a}\right)\leq k$.
+We have the following result.
 Lemma 2
-Let $x\in\mathcal{X}$ be an assignment that is not wasteful for $p$ and $k$ be the smallest integer satisfying
-$\mathop{\displaystyle\sum}\limits_{\bar{o}:R(\bar{o},p_{a})<k}q_{\bar{o}}>\left|\mathcal{A}\right|.$ (2)
-Then, $x_{ao}=0$ for all $o$ such that $R(o,p_{a})\geq k$.
-The condition (2) means that there are sufficient number of copies of object types that ranked higher than $k$ in the preference ranking of agent $a$.
-Then, if the assignment is not wasteful, then $a$ must be assigned to an object type that ranked higher than $k$.
+Let $x\in\mathcal{X}$ be an assignment that is not wasteful for $p$. Then, $x_{ao}=0$ for all $o$ such that $R(o,p_{a})>\bar{k}\left(p_{a}\right)$.
+Then, if an assignment is not wasteful, then $a$ must be assigned to an object type that ranked higher than or equals to $\bar{k}\left(p_{a}\right)$.
 Typically, in that case, the object types that ranked lower than $\varnothing$ in the preference ranking of agent $a$ are never assigned to $a$, because $\varnothing$ is the object with a sufficient number of copies.
@@
 This requires that two agents revealing the same preference must have the same assignment in the ex-ante sense.
-This property is called just “equal treatment of equals” by several previous studies such as Bogomolnaia and Moulin (2001) and Troyan (2024).
+This property is called just “equal treatment of equals” by many previous studies such as Bogomolnaia and Moulin (2001) and Troyan (2024).
 In this study, we mainly use a slightly stronger one.
-A mechanism $f$ is said to satisfy equal treatment of equals (hereafter ETE) if $p\in\mathcal{P}^{\left|\mathcal{A}\right|}$ satisfying $p_{a}^{k^{\prime}}=p_{a^{\prime}}^{k^{\prime}}$ for all $11,\cdots,k-$ 1 implies $\left(f\left(p\right)\right)_{ao}=\left(f\left(p\right)\right)_{a^{\prime}o}$ for all $o\in\mathcal{O}$, where $k$ is the smallest integer satisfying
-$\mathop{\displaystyle\sum}\limits_{\bar{o}:R(\bar{o},p_{a})<k}q_{\bar{o}}\geq\left|\mathcal{A}\right|\text{.}$
-By Lemma 2, as long as the assignment is not wasteful, any type that is ranked $k$th or lower in the preference ranking revealed by $a$ will not be assigned to $a$, because there are sufficient copies of types ranked higher than $k$th.
-Thus, if two individuals reveal rankings that are identical for positions above the $k$th rank, those rankings are considered essentially the same, and they should receive the same assignment in an ex-ante sense.
+A mechanism $f$ is said to satisfy equal treatment of equals (hereafter ETE) if $p\in\mathcal{P}^{\left|\mathcal{A}\right|}$ satisfying $p_{a}^{k^{\prime}}=p_{a^{\prime}}^{k^{\prime}}$ for all $1,\cdots,\bar{k}\left(p_{a}\right)$ implies $\left(f\left(p\right)\right)_{ao}=\left(f\left(p\right)\right)_{a^{\prime}o}$ for all $o\in\mathcal{O}$ [^2] Han (2024) also considers a similar property, called symmetricity, which is slightly stronger than equal treatment of equals.
+Unlike in our model, they take into account the priorities of objects and define agents as equal only if they are equal with respect to the priorities of all objects.
+By Lemma 2, as long as the assignment is not wasteful, any type that is ranked lower than $\bar{k}\left(p_{a}\right)$th in the preference ranking revealed by $a$ will not be assigned to $a$, because there are sufficient copies of types ranked higher than or equal to $\bar{k}\left(p_{a}\right)$th.
+Thus, if two individuals reveal rankings that are identical for positions at or above $\bar{k}\left(p_{a}\right)$th rank, those rankings are considered essentially the same, and they should receive the same assignment in an ex-ante sense.
 We focus on the efficient and fair mechanisms; that is, we call $f$ a fair rank-minimizing mechanism if $f(p)$ is rank efficient for all $p$ and satisfies ETE. The following mechanism is one of them.
 Let $\mathcal{Y}^{\ast}\left(p\right)\subseteq\mathcal{Y}$ be the set of deterministic rank efficient assignments; that is, $y^{\ast}\in\mathcal{Y}^{\ast}\left(p\right),$ $RV\left(y^{\ast},p\right)\leq RV\left(x,p\right)$ for any $x\in\mathcal{X}$.
-Note that by Lemma 1, $\mathcal{Y}^{\ast}\left(p\right)\neq\emptyset$.
+By Corollary 1, $\mathcal{Y}^{\ast}\left(p\right)\neq\emptyset$.
 Let
-$f^{U}\left(p\right)=\mathop{\displaystyle\sum}\limits_{y^{\ast}\in\mathcal{Y}^{\ast}\left(p\right)}\frac{y^{\ast}}{\left|\mathcal{Y}^{\ast}\left(p\right)\right|}\text{ for all }p\in\mathcal{P}^{\left|\mathcal{A}\right|}$
+$f^{U}\left(p\right)=\sum\limits_{y^{\ast}\in\mathcal{Y}^{\ast}\left(p\right)}\frac{y^{\ast}}{\left|\mathcal{Y}^{\ast}\left(p\right)\right|}\text{ for all }p\in\mathcal{P}^{\left|\mathcal{A}\right|}$
 be the uniform rank-minimizing mechanism. By (7),
-$RV\left(f^{U}\left(p\right),p\right)=\mathop{\displaystyle\sum}\limits_{y^{\ast}\in\mathcal{Y}^{\ast}\left(p\right)}\frac{1}{\left|\mathcal{Y}^{\ast}\left(p\right)\right|}RV\left(y^{\ast},p\right).$
+$RV\left(f^{U}\left(p\right),p\right)=\sum\limits_{y^{\ast}\in\mathcal{Y}^{\ast}\left(p\right)}\frac{1}{\left|\mathcal{Y}^{\ast}\left(p\right)\right|}RV\left(y^{\ast},p\right).$
 Thus, $RV\left(f^{U}\left(p\right),p\right)\leq RV\left(x,p\right)$ for any $x\in\mathcal{X}$; that is, $f^{U}\left(p\right)$ must be rank efficient for $p$.
-Trivially, $f^{U}$ satisfies ETE and thus is a fair rank-minimizing mechanism.
-Since Nikzad (2022), Ortega and Klein (2023) and Troyan (2024) also discuss the uniform rank-minimizing mechanism, this is one of the most common rank-minimizing mechanisms at least in the literature.
+Since Nikzad (2022), Ortega and Klein (2023) and Troyan (2024) also discuss the uniform rank-minimizing mechanism, this is one of the most common rank-minimizing mechanisms in the literature.
 We have the following result.
@@
 The uniform rank-minimizing mechanism satisfies ETE.
+Therefore, the uniform rank-minimizing mechanism is a fair rank-minimizing one.
 We discuss strategic problems of fair rank-minimizing mechanisms.
@@
 Featherstone (2020) shows that any rank-minimizing mechanism does not satisfy strategy-proofness.
-That is, for any rank-minimizing mechanism $f$, there are $r_{a}\in\mathcal{P}$ and $p\in\mathcal{P}^{\left|\mathcal{A}\right|}$ such that $a$ with $r_{a}$ prefers $f\left(p\right)$ to $f\left(r_{a},p_{-a}\right)$.
 Therefore, in this paper, we consider a weaker strategic property.
@@
 Let $k$ be the smallest integer such that $p_{ak}\neq r_{ak}$.
-By Lemma 2, if (2) is satisfied for that $k$, then $p_{a}$ and $r_{a}$ are essentially equivalent under a rank-minimizing mechanism.
-Thus, we focus only on the case where (2) is not satisfied for $k$ defined above; that is,
-$\mathop{\displaystyle\sum}\limits_{\bar{o}:R(\bar{o},p_{a})<k}q_{\bar{o}}\left(=\mathop{\displaystyle\sum}\limits_{\bar{o}:R(\bar{o},r_{a})<k}q_{\bar{o}}\right)<\left|\mathcal{A}\right|$ (3)
+By Lemma 2, if $k>\bar{k}\left(p_{a}\right)$, then $p_{a}$ and $r_{a}$ are essentially equivalent under a fair rank-minimizing mechanism, because any result of a rank-minimizing mechanism is not wasteful.
+Thus, we focus only on the case where $k\leq\bar{k}\left(p_{a}\right)$ defined above; that is,
+$\sum\limits_{\bar{o}:R(\bar{o},p_{a})<k}q_{\bar{o}}\left(=\sum\limits_{\bar{o}:R(\bar{o},r_{a})<k}q_{\bar{o}}\right)<\left|\mathcal{A}\right|$
 is satisfied for that $k$.
-In other word, we consider the case where the number of object types that are more preferable than the $k$th best one in $p_{a}$ is insufficient to distribute to all agents.
+In other word, we consider the case where the number of object types that are more preferable than the $k-1$th best one in $p_{a}$ is insufficient to distribute to all agents.
 In this case, for some $p_{-a}\in\mathcal{P}^{\left|\mathcal{A}\right|-1}$ and some fair rank-minimizing mechanism, $a$ is assigned to an object type that is less preferable or equal to the $k$th best one.
@@
 Let $p_{a}\neq r_{a}$, and $k$ be the smallest integer such that $p_{a}^{k}\neq r_{a}^{k}$.
-If (3) is satisfied for that $k$, then there is $p_{-a}\in\mathcal{P}^{\left|\mathcal{A}\right|-1}$ such that agent $a$ (with $r_{a}$) strictly prefers $f\left(r_{a},p_{-a}\right)$ to $f\left(p_{a},p_{-a}\right)$.
-Next, we have the following result.
+If $k\leq\bar{k}\left(p_{a}\right)$, then there is $p_{-a}\in\mathcal{P}^{\left|\mathcal{A}\right|-1}$ such that agent $a$ (with $r_{a}$) strictly prefers $f\left(r_{a},p_{-a}\right)$ to $f\left(p_{a},p_{-a}\right)$.
+By this lemma, we have the following result.
 Proposition 2
 Suppose that a fair rank-minimizing mechanism is used. For any $r_{a}\in\mathcal{P}$, there is no $p_{a}\in\mathcal{P}$ such that revealing $p_{a}$ strategically dominates truth-telling.
-This result implies that any fair rank-minimizing mechanism satisfies a weak strategic characteristic; that is, for each untrue preference of an agent $a$, there must exist a profile of preferences of agents other than $a$, such that $a$ is worse off by revealing it instead of $r_{a}$.
-Thus, if a fair rank-minimizing mechanism is used, any strategic manipulation by an agent carries some level of risk.
+This result implies that any fair rank-minimizing mechanism satisfies a weak strategic characteristic; that is, for each untrue preference $p_{a}$, either $\left(f\left(r_{a},p_{-a}\right)\right)_{ao}=\left(f\left(p_{a},p_{-a}\right)\right)_{ao}$ for all $o\in\mathcal{O}$ and all $p_{-a}\in\mathcal{P}^{\left|\mathcal{A}\right|-1}$ or there must exist $p_{-a}$ such that $a$ is worse off by revealing it instead of $r_{a}$.
+Thus, if a fair rank-minimizing mechanism is used, any (essential) strategic manipulation by an agent carries some level of risk.
 This characteristic is quite weak, but in the subsequent section, we show that this is not satisfied when the agents can refuse their assignment and instead obtain the null object type.
@@
 Example 1
-Let $\mathcal{A=}\left\{a_{1},a_{2},a_{3}\right\}$ and $\mathcal{O=}\left\{o_{1},o_{2},o_{3},\varnothing\right\}$.
+Let $\mathcal{A=}\left\{a_{1},a_{2},a_{3}\right\}$, $\mathcal{O=}\left\{o_{1},o_{2},o_{3},\varnothing\right\}$ and $q_{o_{1}}=1$.
 Let $\psi=\left(o_{1},o_{2},\varnothing,o_{3}\right)$ and $\psi^{\prime}=\left(o_{1},o_{2},o_{3},\varnothing\right)$.
@@
 $\left(f\left(\psi^{\prime},\psi,\psi\right)\right)_{a_{1}o_{1}}$ $=$ $\left(f\left(\psi,\psi^{\prime},\psi\right)\right)_{a_{2}o_{1}}=\left(f\left(\psi,\psi,\psi^{\prime}\right)\right)_{a_{3}o_{1}}=0,$ (2)
-$f\left(p_{a_{1}},p_{a_{2}},p_{a_{3}}\right)$ $=$ $\mathop{\displaystyle\sum}\limits_{y^{\ast}\in\mathcal{Y}^{\ast}\left(p\right)}\frac{y^{\ast}}{\left|\mathcal{Y}^{\ast}\left(p\right)\right|}$
+$f\left(p_{a_{1}},p_{a_{2}},p_{a_{3}}\right)$ $=$ $\sum\limits_{y^{\ast}\in\mathcal{Y}^{\ast}\left(p\right)}\frac{y^{\ast}}{\left|\mathcal{Y}^{\ast}\left(p\right)\right|}$
 if $\left(p_{a_{1}},p_{a_{2}},p_{a_{3}}\right)$ is neither $\left(\psi^{\prime},\psi,\psi\right)$ nor $\left(\psi,\psi^{\prime},\psi\right)$ nor $\left(\psi,\psi,\psi^{\prime}\right)$.
@@
 Trivially, $f$ satisfies weak ETE.
-However, whether $f$ satisfies ETE or not is dependent on $q_{o_{1}}$ and $q_{o_{2}}$.
-Suppose $r_{a_{1}}=\psi^{\prime}$ and this rank-minimizing mechanism is used.
-First, we assume $q_{o_{1}}=q_{o_{2}}=1$.
-In this case, $r_{a_{1}}=\psi^{\prime}$ is not strategically dominated by $\psi$ and any other preference orders, because
+However, whether $f$ satisfies ETE or not is dependent on $q_{o_{2}}$.
+We assume that $r_{a_{1}}=\psi^{\prime}$ and this rank-minimizing mechanism is used.
+First, suppose $q_{o_{2}}=1$.
+In this case, $\bar{k}\left(\psi\right)=\bar{k}\left(\psi^{\prime}\right)=3$.
+Then, $r_{a_{1}}=\psi^{\prime}$ is not strategically dominated by $\psi$ and any other preference orders, because
 $\left(f\left(\psi^{\prime},\psi^{\prime},\psi^{\prime}\right)\right)_{a_{1}o_{1}}$ $=$ $\left(f\left(\psi^{\prime},\psi^{\prime},\psi^{\prime}\right)\right)_{a_{1}o_{2}}=\left(f\left(\psi^{\prime},\psi^{\prime},\psi^{\prime}\right)\right)_{a_{1}o_{3}}=\frac{1}{3},$
@@
 Moreover, in this case, $f$ satisfies ETE and thus this is a fair rank-minimizing mechanism.
-On the other hand, suppose $q_{o_{1}}=1$ and $q_{o_{2}}\geq 2$.
-In this case, $a_{1}$ must have either $o_{1}$ or $o_{2}$ if it chooses either $\psi$ or $\psi^{\prime}$ and thus, the unique interest of $a_{1}$ is increasing the probability to obtain $o_{1}$.
+Second, suppose $q_{o_{2}}\geq 2$.
+In this case, $\bar{k}\left(\psi\right)=\bar{k}\left(\psi^{\prime}\right)=2$.
+Then, $a_{1}$ must have either $o_{1}$ or $o_{2}$ if it chooses either $\psi$ or $\psi^{\prime}$ and thus, the unique interest of $a_{1}$ is increasing the probability to obtain $o_{1}$.
 Thus, in this case, $r_{a_{1}}=\psi^{\prime}$ is strategically dominated by $\psi$.
@@
 $g\left(x,r\right)_{ao}$ $=$ 0 for all $0\text{ for all }R\left(o,r_{a}\right)>R\left(\varnothing,r_{a}\right)$
-$g\left(x,r\right)_{a\varnothing}$ $=$ $\mathop{\displaystyle\sum}\limits_{o:R\left(o,r_{a}\right)\geq R\left(\varnothing,r_{a}\right)}x_{ao}.$
+$g\left(x,r\right)_{a\varnothing}$ $=$ $\sum\limits_{o:R\left(o,r_{a}\right)\geq R\left(\varnothing,r_{a}\right)}x_{ao}.$
 That is, since an agent refuses any of the unacceptable object types when it is assigned, the probability that the null object type is assigned will increase by the probability that any of the unacceptable object types is assigned.
@@
 For a mechanism $f$, (revealing) $p_{a}^{\prime}$ (for $a$) is said to be strategically dominated by (revealing) $p_{a}$ with the refusal option if $a$ weakly prefers $g\left(f\left(p_{a},p_{-a}\right),r\right)$ to $g\left(f\left(p_{a}^{\prime},p_{-a}\right),r\right)$ for all $p_{-a}\in\mathcal{P}^{\left|\mathcal{A}\right|-1}$ and strictly prefers $g\left(f\left(p_{a},p_{-a}^{\prime}\right),r\right)$ to $g\left(f\left(p_{a}^{\prime},p_{-a}^{\prime}\right),r\right)$ for some $p_{-a}^{\prime}\in\mathcal{P}^{\left|\mathcal{A}\right|-1}$.
-We show that, under the uniform rank-minimizing mechanism, truth-telling may be strategically dominated by another strategy when the agents have the refusal option.
+We show that, under the uniform rank-minimizing mechanism, truth-telling is strategically dominated by another strategy when the agents have the refusal option.
 We let $\mathcal{D}_{a}\subseteq\mathcal{P}$ be a set of preferences such that $d_{a}\in\mathcal{D}_{a}$ if (i) $R\left(o,r_{a}\right)=R\left(o,d_{a}\right)$ for all $o\in\mathcal{O}$ such that $R\left(o,r_{a}\right)<R\left(\varnothing,r_{a}\right)$ and (ii) $R\left(\varnothing,d_{a}\right)=\left|\mathcal{O}\right|$.
@@
 Note that an ODS (revealing $d_{a}\in\mathcal{D}_{a}$) increases the number of acceptable object types revealed by $a$; that is, $R\left(\varnothing,r_{a}\right)\leq R\left(\varnothing,d_{a}\right)=\left|\mathcal{O}\right|$ is satisfied.
-Featherstone (2020) defines the full extension of $r_{a}$denoted by $e_{a}$ in this study such that (i) $R\left(o,r_{a}\right)=R\left(o,e_{a}\right)$ for all $o\in\mathcal{O}$ such that $R\left(o,r_{a}\right)<R\left(\varnothing,e_{a}\right)$, and (ii) $R\left(\varnothing,e_{a}\right)=\left|\mathcal{O}\right|$, and $R\left(o,r_{a}\right)-1=R\left(o,e_{a}\right)$ for all $o\in\mathcal{O}$ with $R\left(o,r_{a}\right)>R\left(\varnothing,e_{a}\right)$; that is, $e_{a}\in\mathcal{D}_{a}$ is such that all non-null object types are placed in their true preference orders.
+Featherstone (2020) defines the full extension of $r_{a}$denoted by $e_{a}$ in this study such that (i) $R\left(o,r_{a}\right)=R\left(o,e_{a}\right)$ for all $o\in\mathcal{O}$ such that $R\left(o,r_{a}\right)<R\left(\varnothing,r_{a}\right)$, and (ii) $R\left(\varnothing,e_{a}\right)=\left|\mathcal{O}\right|$, and $R\left(o,r_{a}\right)-1=R\left(o,e_{a}\right)$ for all $o\in\mathcal{O}$ such that $R\left(o,r_{a}\right)>R\left(\varnothing,r_{a}\right)$; that is, $e_{a}\in\mathcal{D}_{a}$ is such that all non-null object types are placed in their true preference orders.
 First, we consider the following example to briefly consider the main result.
@@
 which is wasteful, because
-$\mathop{\textstyle\sum}\limits_{a\in\left\{a_{1},a_{2},a_{3}\right\}}g\left(f^{U}\left(d_{a_{1}},\hat{p}_{a_{2}},\hat{p}_{a_{3}}\right),r\right)_{ao_{2}}$ $=$ $\frac{2}{3}<1,$
+$\sum\limits_{a\in\left\{a_{1},a_{2},a_{3}\right\}}g\left(f^{U}\left(d_{a_{1}},\hat{p}_{a_{2}},\hat{p}_{a_{3}}\right),r\right)_{ao_{2}}$ $=$ $\frac{2}{3}<1,$
 and $\text{and }g\left(f^{U}\left(d_{a_{1}},\hat{p}_{a_{2}},\hat{p}_{a_{3}}\right),r\right)_{a^{\prime}\varnothing}$ $=$ $\frac{1}{3}>0,$
-for $a^{\prime}=a_{2},a_{3}$. Moreover, it is ordinally dominated by
+for $a^{\prime}=a_{2},a_{3}$. Thus, it is ordinally dominated by
 $\tilde{x}=\left(\begin{array}[]{ccc}1/3&0&2/3\\ 1/3&1/2&1/6\\ 1/3&1/2&1/6\end{array}\right).$
@@
 By this result, an agent is never hurt by using an ODS instead of truth-telling.
-Next, we show that with some weak condition, there exists some ODS such that an agent may be better off by using it instead of truth-telling.
+Next, we show that with some conditions, there exists some ODS such that an agent may be better off by using it instead of truth-telling.
 Theorem 2
 Suppose that there is $a\in\mathcal{A}$ whose true preference $r_{a}$ satisfies
-$R\left(o,r_{a}\right)<R\left(\varnothing,r_{a}\right)<R\left(o^{\prime},r_{a}\right)\text{ and }\mathop{\textstyle\sum}\limits_{\bar{o}:R\left(\bar{o},r_{a}\right)<R\left(\varnothing,r_{a}\right)}q_{\bar{o}}+q_{o^{\prime}}<\left|\mathcal{A}\right|\text{,}$ (5)
+$R\left(o,r_{a}\right)<R\left(\varnothing,r_{a}\right)<R\left(o^{\prime},r_{a}\right)\text{ and }\sum\limits_{\bar{o}:R\left(\bar{o},r_{a}\right)<R\left(\varnothing,r_{a}\right)}q_{\bar{o}}+q_{o^{\prime}}<\left|\mathcal{A}\right|\text{,}$ (3)
 […]
@@
 This is satisfied only when some agent has at least one unacceptable object.
-The condition given in (5) is that there is a truly unacceptable type $o^{\prime}$ such that the number of copies of $o^{\prime}$, plus the sum of the copies of the truly acceptable types, is not very large.
+The condition given in (3) is that there is a truly unacceptable type $o^{\prime}$ such that the number of copies of $o^{\prime}$, plus the sum of the copies of the truly acceptable types, is sufficiently small.
 We discuss the cases where (3) is satisfied for some $a,o,o^{\prime}$ in the subsequent section.
@@
 Suppose that there is $a\in\mathcal{A}$ whose true preference $r_{a}$ satisfies
-$R\left(o,r_{a}\right)<R\left(\varnothing,r_{a}\right)<R\left(o^{\prime},r_{a}\right)\text{ and }\mathop{\textstyle\sum}\limits_{\bar{o}:R\left(\bar{o},r_{a}\right)<R\left(\varnothing,r_{a}\right)}q_{\bar{o}}+q_{o^{\prime}}\geq\left|\mathcal{A}\right|,$
+$R\left(o,r_{a}\right)<R\left(\varnothing,r_{a}\right)<R\left(o^{\prime},r_{a}\right)\text{ and }\sum\limits_{\bar{o}:R\left(\bar{o},r_{a}\right)<R\left(\varnothing,r_{a}\right)}q_{\bar{o}}+q_{o^{\prime}}\geq\left|\mathcal{A}\right|,$
 for some $o,o^{\prime}\in\mathcal{O}$.
@@
 $R\left(o,r_{a}\right)<R\left(\varnothing,r_{a}\right)=R\left(o^{\prime},r_{a}\right)-1<R\left(o^{\prime\prime},r_{a}\right),\text{ }$
-$\mathop{\textstyle\sum}\limits_{\bar{o}:R\left(\bar{o},r_{a}\right)<R\left(\varnothing,r_{a}\right)}q_{\bar{o}}+q_{o^{\prime}}\geq\left|\mathcal{A}\right|,\text{ and}$
-$\mathop{\textstyle\sum}\limits_{\bar{o}:R\left(\bar{o},r_{a}\right)<R\left(\varnothing,r_{a}\right)}q_{\bar{o}}+q_{o^{\prime\prime}}<\left|\mathcal{A}\right|,$
+$\sum\limits_{\bar{o}:R\left(\bar{o},r_{a}\right)<R\left(\varnothing,r_{a}\right)}q_{\bar{o}}+q_{o^{\prime}}\geq\left|\mathcal{A}\right|,\text{ and}$
+$\sum\limits_{\bar{o}:R\left(\bar{o},r_{a}\right)<R\left(\varnothing,r_{a}\right)}q_{\bar{o}}+q_{o^{\prime\prime}}<\left|\mathcal{A}\right|,$
 […]
@@
 Thus, next, we consider another fair rank-minimizing mechanism such that, contrary to $f^{U}$, if it is used as the mechanism in the second stage, then for any agent, truth-telling is not strategically dominated by any other strategy.
-We let $f^{M}$ be the modified uniform rank-minimizing mechanism such that
-$\left(f^{M}\left(p\right)\right)_{ao}=0,\text{ }\left(f^{M}\left(p\right)\right)_{a^{\prime}o}=1,\text{ }\left(f^{M}\left(p\right)\right)_{ap_{a}^{2}}=1\text{ and }\left(f^{M}\left(p\right)\right)_{a^{\prime\prime}\varnothing}=1$
-if $p$ satisfies
-there is one agent $a$ such that $R\left(o,p_{a}\right)=1$ and $R\left(\varnothing,p_{a}\right)\geq 3,$
-there are $q_{o}$ agents (including $a^{\prime}$) with revealed preferences equivalent to those represented by $p_{a}^{\prime}$, such that $R\left(\varnothing,p_{a^{\prime}}\right)=l\geq 2,$ $p_{a}^{l^{\prime}}=p_{a^{\prime}}^{l}$ for all $l^{\prime}=1,\cdots,l-1,$ $R\left(\varnothing,p_{a}\right)>R\left(\varnothing,p_{a^{\prime}}\right),$ and
-$\mathop{\displaystyle\sum}\limits_{\bar{o}:R\left(\bar{o},p_{a^{\prime}}\right)<R\left(\varnothing,p_{a^{\prime}}\right)}q_{\bar{o}}<\left|\mathcal{A}\right|,$
-if $\left|\mathcal{A}\right|>q_{o}+1$, then the first best type of the other agents ($\left|\mathcal{A}\right|-\left(q_{o}+1\right)$ agents) represented by $a^{\prime\prime}$ is $\varnothing$,
-and if otherwise
-$f^{M}\left(p\right)=\mathop{\displaystyle\sum}\limits_{y^{\ast}\in\mathcal{Y}^{\ast}\left(p\right)}\frac{y^{\ast}}{\left|\mathcal{Y}^{\ast}\left(p\right)\right|}=f^{U}\left(p\right).$
+We let $f^{M}$ be the modified uniform rank-minimizing mechanism such that if $p$ satisfies
+for agent $a,$ $R\left(o,p_{a}\right)=1$ and $R\left(\varnothing,p_{a}\right)\geq 3,$
+there are $q_{o}$ agents represented by $a^{\prime}$ revealing $p_{a^{\prime}}$, such that $R\left(\varnothing,p_{a^{\prime}}\right)=l\geq 2,$ $p_{a}^{l^{\prime}}=p_{a^{\prime}}^{l}$ for all $l^{\prime}=1,\cdots,l-1,$ $R\left(\varnothing,p_{a}\right)>l,$ and $\bar{k}(p_{a^{\prime}})=l$; that is,
+$\sum\limits_{\bar{o}:R\left(\bar{o},p_{a^{\prime}}\right)<R\left(\varnothing,p_{a^{\prime}}\right)}q_{\bar{o}}<\left|\mathcal{A}\right|,$
+if $\left|\mathcal{A}\right|>q_{o}+1$, then the first best type of the other agents ($\left|\mathcal{A}\right|-\left(q_{o}+1\right)$ agents) represented by $a^{\prime\prime}$ is $\varnothing$, then
+$\left(f^{M}\left(p\right)\right)_{ao}=0,\text{ }\left(f^{M}\left(p\right)\right)_{a^{\prime}o}=1,\text{ }\left(f^{M}\left(p\right)\right)_{ap_{a}^{2}}=1\text{ and }\left(f^{M}\left(p\right)\right)_{a^{\prime\prime}\varnothing}=1,$
+and if $p$ does not satisfy all of them, then
+$f^{M}\left(p\right)=\sum\limits_{y^{\ast}\in\mathcal{Y}^{\ast}\left(p\right)}\frac{y^{\ast}}{\left|\mathcal{Y}^{\ast}\left(p\right)\right|}=f^{U}\left(p\right).$
 To understand this mechanism, we reconsider Example 1.
@@
 If $\left(p_{a_{1}},p_{a_{2}},p_{a_{3}}\right)=\left(\psi^{\prime},\psi,\psi\right)$, $\left(\psi,\psi^{\prime},\psi\right)$ or $\left(\psi,\psi,\psi^{\prime}\right)$, then this satisfies (i), (ii) and (iii).
-Hence the rank-minimizing mechanism $f$ introduced in Example 1 is the modified uniform rank-minimizing mechanism.
+In these cases, a student who chooses $\psi^{\prime}$ is corresponding to $a,$ another student is corresponding to $a^{\prime}$, and $o_{1}$ is corresponding to $o$.
+The rank-minimizing mechanism $f$ introduced in Example 1 is the modified uniform rank-minimizing mechanism, because a student who chooses $\psi^{\prime}$ obtains $o_{1}$ with probability 00 and another student obtains $o_{1}$ with probability $1$
 On the other hand, suppose $q_{o_{1}}=1$ and $q_{o_{2}}\geq 2$.
 Even if $\left(p_{a_{1}},p_{a_{2}},p_{a_{3}}\right)=\left(\psi^{\prime},\psi,\psi\right)$, $\left(\psi,\psi^{\prime},\psi\right)$ or $\left(\psi,\psi,\psi^{\prime}\right)$, this does not satisfy (ii), because in this case, $p_{a^{\prime}}=\psi$ and $l=3.$ Then,
-$\mathop{\displaystyle\sum}\limits_{\bar{o}:R\left(\bar{o},p_{a^{\prime}}\right)<R\left(\varnothing,p_{a^{\prime}}\right)}q_{\bar{o}}=q_{o_{1}}+q_{o_{2}}=3=\left|\mathcal{A}\right|.$
+$\sum\limits_{\bar{o}:R\left(\bar{o},p_{a^{\prime}}\right)<R\left(\varnothing,p_{a^{\prime}}\right)}q_{\bar{o}}=q_{o_{1}}+q_{o_{2}}=3=\left|\mathcal{A}\right|.$
 Therefore, (ii) is not satisfied with $p_{a^{\prime}}=\psi$.
@@
 Let $R\left(\varnothing,r_{a}\right)=k$.
-Suppose $k\geq 3$; that is, $\varnothing$ in neither the best type nor the second best type for $r_{a}$.
+Suppose $k\geq 3$; that is, $\varnothing$ is neither the best type nor the second best type for $r_{a}$.
 Thus, we let $r_{a}^{1}=o$ and $r_{a}^{2}=o^{\prime}$.
@@
 In this case, since $\left(r_{a},p_{-a}\right)$ satisfies (i), (ii) and (iii),
-$\left(f^{M}\left(r_{a},p_{-a}\right)\right)_{ao^{\prime}}$ $=$ 0, ($0,\text{ }\left(f^{M}\left(r_{a},p_{-a}\right)\right)_{ao^{\prime}}=1,$
+$\left(f^{M}\left(r_{a},p_{-a}\right)\right)_{ao}$ $=$ 0, ($0,\text{ }\left(f^{M}\left(r_{a},p_{-a}\right)\right)_{ao^{\prime}}=1,$
 $\left(f^{U}\left(r_{a},p_{-a}\right)\right)_{ao}$ $=$ $\frac{q_{o}}{q_{o}+1},$
@@
 $\left(f^{M}\left(p_{a},p_{-a}\right)\right)_{ao^{\prime}}$ $=$ $\left(f^{U}\left(p_{a},p_{-a}\right)\right)_{ao^{\prime}}=\frac{1}{q_{o}+1}.$
-Thus, in this case, if $f^{M}$ is used, then $a$ has an incentive to reveal by reducing its truly acceptable types. On the other hand, if $f^{U}$ is used, then $a$ has no such an incentive.
+Thus, in this case, if $f^{M}$ is used, then $a$ has an incentive to choose $p_{a}$.
+In other words, in this case, $a$ has an incentive to reveal by reducing its truly acceptable types.
+On the other hand, if $f^{U}$ is used, then $a$ has no such an incentive.
 Of course, this does not mean that truth-telling is strategically dominated revealing some reduction of $r_{a}$. However, agents may have such incentives even if they do not have the refusal option.
 5 Concluding Remarks
-We show that if an assignment is determined by the uniform rank-minimizing mechanism, which is a fair rank-minimizing mechanism, and agents can refuse the assignment and obtain the outside option instead, then the true preference is strategically dominated by an ODS.
+We show that if an assignment is determined by the uniform rank-minimizing mechanism, which is a fair one, and agents can refuse the assignment and obtain the outside option instead, then the true preference is strategically dominated by an ODS.
 The adoption of such strategies by agents may result in inefficient assignments; that is, an agent who adopts an ODS refuses their assigned type, although some other agents desire that type.
@@
 Suppose not; that is, there is $y^{i}$ such that $RV\left(y^{i},p\right)>RV\left(x^{\ast},p\right)$.
-$RV\left(x^{\ast},p\right)$ $=$ $\mathop{\displaystyle\sum}\limits_{a\in\mathcal{A}}\mathop{\displaystyle\sum}\limits_{o\in\mathcal{O}}R\left(o,p_{a}\right)\times x_{ao}^{\ast}$ (9)
-$=$ $\mathop{\displaystyle\sum}\limits_{a\in\mathcal{A}}\mathop{\displaystyle\sum}\limits_{o\in\mathcal{O}}\left(R\left(o,p_{a}\right)\times\mathop{\displaystyle\sum}\limits_{i=1}^{I}\sigma^{i}y_{ao}^{i}\right)$
-$=$ $\mathop{\displaystyle\sum}\limits_{i=1}^{I}\sigma^{i}\mathop{\displaystyle\sum}\limits_{a\in\mathcal{A}}\mathop{\displaystyle\sum}\limits_{o\in\mathcal{O}}R\left(o,p_{a}\right)\times y_{ao}^{i}$
-$=$ $\mathop{\displaystyle\sum}\limits_{i=1}^{I}\sigma^{i}RV\left(y^{i},p\right).$
+$RV\left(x^{\ast},p\right)$ $=$ $\sum\limits_{a\in\mathcal{A}}\sum\limits_{o\in\mathcal{O}}R\left(o,p_{a}\right)\times x_{ao}^{\ast}$ (7)
+$=$ $\sum\limits_{a\in\mathcal{A}}\sum\limits_{o\in\mathcal{O}}\left(R\left(o,p_{a}\right)\times\sum\limits_{i=1}^{I}\sigma^{i}y_{ao}^{i}\right)$
+$=$ $\sum\limits_{i=1}^{I}\sigma^{i}\sum\limits_{a\in\mathcal{A}}\sum\limits_{o\in\mathcal{O}}R\left(o,p_{a}\right)\times y_{ao}^{i}$
+$=$ $\sum\limits_{i=1}^{I}\sigma^{i}RV\left(y^{i},p\right).$
 Thus, $RV\left(y^{i},p\right)>RV\left(x^{\ast},p\right)$ implies that there is $y^{j}$ such that $RV\left(y^{j},p\right)<RV\left(x^{\ast},p\right)$, which contradicts that $x^{\ast}$ is a rank efficient assignment.
@@
 Fix $p\in\mathcal{P}^{\left|\mathcal{A}\right|}$ and $a\in\mathcal{A}$.
-Moreover, let $k$ be the smallest integer satisfying (2) for $p_{a}$.
-Suppose not; that is, for an assignment $x\in\mathcal{X}$ that is not wasteful, there is $o\in\mathcal{O}$ satisfying $R(o,p_{a})\geq k$ and $x_{ao}>0$.
-First, since $\mathop{\textstyle\sum}_{\bar{o}\in\mathcal{O}}x_{a\bar{o}}=1$,
-$\mathop{\displaystyle\sum}\limits_{\bar{o}\in\mathcal{O}}\mathop{\displaystyle\sum}\limits_{\bar{a}\in\mathcal{A}}x_{\bar{a}\bar{o}}=\left|\mathcal{A}\right|\text{.}$
-Moreover, since $x_{ao}>0$ and (2) are satisfied,
-$\mathop{\displaystyle\sum}\limits_{\bar{o}:R(\bar{o},p_{a})<k}\mathop{\displaystyle\sum}\limits_{\bar{a}\in\mathcal{A}}x_{\bar{a}o^{\prime}}$ $<$ $\left|\mathcal{A}\right|\leq\mathop{\displaystyle\sum}\limits_{\bar{o}:R(\bar{o},p_{a})<k}q_{o^{\prime}}\text{, and}$
-$\mathop{\displaystyle\sum}\limits_{\bar{o}:R(\bar{o},p_{a})<k}x_{a\bar{o}}$ $<$ $1$
-Since $\mathop{\textstyle\sum}_{\bar{a}\in\mathcal{A}}x_{\bar{a}o^{\prime\prime}}\leq q_{o^{\prime\prime}}$ for all $o^{\prime\prime}\in\mathcal{O}$, there is some $o^{\prime}\in\mathcal{O}$ such that $R(o^{\prime},p_{a})\leq k$ and $\mathop{\textstyle\sum}_{\bar{a}\in\mathcal{A}}x_{\bar{a}o^{\prime}}<q_{o^{\prime}}$.
-Then, $\mathop{\textstyle\sum}_{\bar{a}\in\mathcal{A}}x_{\bar{a}o^{\prime}}<q_{o^{\prime}},$ $x_{ao}>0,$ and $R(o^{\prime},p_{a})\leq k<R(o,p_{a})$ contradict that $x\in\mathcal{X}$ is not wasteful.
+Suppose not; that is, for an assignment $x\in\mathcal{X}$ that is not wasteful for $p$, there is $o\in\mathcal{O}$ satisfying $R(o,p_{a})>\bar{k}\left(p_{a}\right)$ and $x_{ao}>0$.
+First, since $\sum_{\bar{o}\in\mathcal{O}}x_{a\bar{o}}=1$,
+$\sum\limits_{\bar{o}\in\mathcal{O}}\sum\limits_{\bar{a}\in\mathcal{A}}x_{\bar{a}\bar{o}}=\left|\mathcal{A}\right|\text{.}$
+Moreover, since $x_{ao}>0$,
+$\sum\limits_{\bar{o}:R(\bar{o},p_{a})\leq\bar{k}\left(p_{a}\right)}\sum\limits_{\bar{a}\in\mathcal{A}}x_{\bar{a}o^{\prime}}$ $<$ $\left|\mathcal{A}\right|\leq\sum\limits_{\bar{o}:R(\bar{o},p_{a})\leq\bar{k}\left(p_{a}\right)}q_{o^{\prime}}\text{, and}$
+$\sum\limits_{\bar{o}:R(\bar{o},p_{a})\leq\bar{k}\left(p_{a}\right)}x_{a\bar{o}}$ $<$ $1$
+Since $\sum_{\bar{a}\in\mathcal{A}}x_{\bar{a}o^{\prime\prime}}\leq q_{o^{\prime\prime}}$ for all $o^{\prime\prime}\in\mathcal{O}$, there is some $o^{\prime}\in\mathcal{O}$ such that $R(o^{\prime},p_{a})\leq\bar{k}\left(p_{a}\right)$ and $\sum_{\bar{a}\in\mathcal{A}}x_{\bar{a}o^{\prime}}<q_{o^{\prime}}$.
+Then, $\sum_{\bar{a}\in\mathcal{A}}x_{\bar{a}o^{\prime}}<q_{o^{\prime}},$ $x_{ao}>0,$ and $R(o^{\prime},p_{a})\leq\bar{k}\left(p_{a}\right)<R(o,p_{a})$ contradict that $x\in\mathcal{X}$ is not wasteful for $p$.
 Proof of Proposition 1
-Let $p_{a}\in\mathcal{P}$ and $k$ be the smallest integer satisfying (2) for $p_{a}$.
-Moreover, let $p\in\mathcal{P}^{\left|\mathcal{A}\right|}$ satisfy $p_{a}=p_{a^{\prime}}$.
+Let $p\in\mathcal{P}^{\left|\mathcal{A}\right|}$ satisfy $p_{a}^{k^{\prime}}=p_{a^{\prime}}^{k^{\prime}}$ for all $1,\cdots,\bar{k}\left(p_{a}\right)$.
 Suppose not; that is, $\left(f^{U}\left(p\right)\right)_{ao}>\left(f^{U}\left(p\right)\right)_{a^{\prime}o}$ for some $o\in\mathcal{O}$.
-Since $\left(f^{U}\left(p\right)\right)_{ao}>0$ and $f^{U}\left(p\right)$ is not wasteful, $R\left(o,p_{a}\right)<k$.
+Since $\left(f^{U}\left(p\right)\right)_{ao}>0$ and $f^{U}\left(p\right)$ is not wasteful, $R\left(o,p_{a}\right)\leq\bar{k}\left(p_{a}\right)$, by Lemma 2.
 Therefore, $R\left(o,p_{a}\right)=R\left(o,p_{a^{\prime}}\right)$.
@@
 Proof of Lemma 3
-Suppose that (3) is satisfied for $k$ defined in this lemma.
-Then, since $q_{\varnothing}\geq\left|\mathcal{A}\right|$, $r_{a}^{k^{\prime}}=p_{a}^{k^{\prime}}\neq\varnothing$ for all $k^{\prime}=1,\cdots,k-1$.
-$1\leq k<\left|\mathcal{O}\right|$.
+Suppose $k\leq\bar{k}\left(p_{a}\right)$. Then, since $q_{\varnothing}\geq\left|\mathcal{A}\right|$, $r_{a}^{k^{\prime}}=p_{a}^{k^{\prime}}\neq\varnothing$ for all $k^{\prime}=1,\cdots,k-1$.
 When $k=1$, we let $R(\varnothing,p_{a^{\prime}})=1$ for all $a^{\prime}\in\mathcal{A\setminus}\left\{a\right\}$. Then,
@@
 $\left(f\left(p_{a},p_{-a}\right)\right)_{ap_{a}^{2}}$ $=$ $\frac{1}{q_{r_{a}^{1}}+1}.$
-Since $p_{a}^{1}=r_{a}^{1}$ and $p_{a}^{2}\neq r_{a}^{2}$, $a$ prefers $r_{a}^{2}$ to $p_{a}^{2}$. Therefore, $a$ strictly prefers $f\left(r_{a},p_{-a}\right)$ to $f\left(p_{a},p_{-a}\right)$.
+Since $p_{a}^{1}=r_{a}^{1}$ and $p_{a}^{2}\neq r_{a}^{2}$, $a$ strictly prefers $f\left(r_{a},p_{-a}\right)$ to $f\left(p_{a},p_{-a}\right)$.
 We consider an arbitrary integer $k\geq 3$.
@@
 $p_{a_{k-1}}$ $=$ $\left(r_{a}^{k-1},r_{a}^{1},\ldots,r_{a}^{k-3},r_{a}^{k-2},r_{a}^{k},\ldots\right).$
-Since
-$q_{r_{a}^{1}}+q_{r_{a}^{2}}+\cdots+q_{r_{a}^{k-1}}=\mathop{\displaystyle\sum}\limits_{o:R(o,r_{a})<k}q_{o}<\left|\mathcal{A}\right|$
-by (3), $\left(r_{a},p_{-a}\right)$ is well-defined so far.
+Since $k\leq\bar{k}\left(p_{a}\right)$,
+$q_{r_{a}^{1}}+q_{r_{a}^{2}}+\cdots+q_{r_{a}^{k-1}}=\sum\limits_{o:R(o,r_{a})<k}q_{o}<\left|\mathcal{A}\right|.$
+and thus $\left(r_{a},p_{-a}\right)$ is well-defined so far.
 Finally, for all the remaining agents (belonging to the $k$th group), the best object type is $\varnothing$ under $\left(r_{a},p_{-a}\right)$.
@@
 $\left(f\left(p_{a},p_{-a}\right)\right)_{ap_{a}^{k}}$ $=$ $\frac{1}{q_{r_{a}^{1}}+1}.$
-Since $p_{a}^{k^{\prime}}=r_{a}^{k^{\prime}}$ for all $k^{\prime}=1,\cdots,k-1$ and $p_{a}^{k}\neq r_{a}^{k}$, $a$ prefers $r_{a}^{k}$ to $p_{a}^{k}$ under $r_{a}$.
-Therefore, $a$ strictly prefers $f^{U}\left(r_{a},p_{-a}\right)$ to $f^{U}\left(p_{a},p_{-a}\right)$ under $r_{a}$.
+Since $p_{a}^{k^{\prime}}=r_{a}^{k^{\prime}}$ for all $k^{\prime}=1,\cdots,k-1$ and $p_{a}^{k}\neq r_{a}^{k}$, $a$ strictly prefers $f^{U}\left(r_{a},p_{-a}\right)$ to $f^{U}\left(p_{a},p_{-a}\right)$.
 Thus, we have the first result.
-We show the second result. Suppose that (2) is satisfied for $p_{a}$ and $k$ defined in the lemma that is,
-$\mathop{\displaystyle\sum}\limits_{o:R(o,p_{a})<k}q_{o}=\mathop{\displaystyle\sum}\limits_{o:R(o,r_{a})<k}q_{o}\geq\left|\mathcal{A}\right|.$
-By Lemma 1 and Remark 2, $f^{U}\left(p\right)_{ao}=f^{U}\left(r_{a},p_{-a}\right)_{ao}=0$ for all $o$ such that $R(o,p_{a})\geq k$.
-Therefore, for any $y^{\ast}\in\mathcal{Y}^{\ast}\left(r_{a},p_{-a}\right)$, $y_{ao}^{\ast}=0$; that is, for any assignment $x$ such that $x_{ao}>0$, $RV\left(y^{\ast},p\right)<RV\left(x,p\right)$.
-Therefore, $\mathcal{Y}^{\ast}\left(r_{a},p_{-a}\right)=\mathcal{Y}^{\ast}\left(p\right)$.
-Since $f^{U}$ is used, we have the second result.
 Proof of Proposition 2
-Let $f$ be a fair mechanism, $p_{a}\neq r_{a}$ and $k$ be the smallest integer such that $p_{ak}\neq r_{ak}$, which is also defined in Lemma 2.
-First, suppose (3) holds for that $k$.
-By the first result of Lemma 2, there is $p_{-a}\in\mathcal{P}^{\left|\mathcal{A}\right|-1}$ such that $a$ strictly prefers $f\left(r_{a},p_{-a}\right)$ to $f\left(p_{a},p_{-a}\right)$.
+Let $f$ be a fair mechanism, $p_{a}\neq r_{a}$ and $k$ be the smallest integer such that $p_{ak}\neq r_{ak}$.
+First, suppose $k\leq\bar{k}\left(p_{a}\right)$.
+By Lemma 2, there is $p_{-a}\in\mathcal{P}^{\left|\mathcal{A}\right|-1}$ such that $a$ strictly prefers $f\left(r_{a},p_{-a}\right)$ to $f\left(p_{a},p_{-a}\right)$.
 Therefore, in this case, $r_{a}$ is not strategically dominate by $p_{a}$.
-Second, suppose (3) does not hold for that $k$.
-In this case, since $r_{a}$ and $p_{a}$ are essentially identical and $f$ satisfies ETE, $f\left(r_{a},p_{-a}\right)=f\left(p_{a},p_{-a}\right)$.
+Second, suppose $k>\bar{k}\left(p_{a}\right)$.
+In this case, since $f$ satisfies ETE, $f\left(r_{a},p_{-a}\right)=f\left(p_{a},p_{-a}\right)$.
 Therefore, $p_{a}$ does not strategically dominate $r_{a}$.
@@
 Claim 1
-Let $y$ be such that $y_{ao^{\prime\prime}}=1$ where $R\left(o^{\prime\prime},r_{a}\right)<R\left(\varnothing,r_{a}\right)$.
-Then, $y\in\mathcal{Y}^{\ast}\left(r_{a},p_{-a}\right)$ if and only if $y\in\mathcal{Y}^{\ast}\left(d_{a},p_{-a}\right)$.
+If $y\in\mathcal{Y}^{\ast}\left(r_{a},p_{-a}\right)$ such that $y_{ao}=1$ where $R\left(o,r_{a}\right)<R\left(\varnothing,r_{a}\right)$, then $y\in\mathcal{Y}^{\ast}\left(d_{a},p_{-a}\right)$.
 Proof. By the construction,
 $RV\left(y,\left(r_{a},p_{-a}\right)\right)=RV\left(y,\left(d_{a},p_{-a}\right)\right).$
-First, we show the only-if-part.
-Suppose $y\in\mathcal{Y}^{\ast}\left(r_{a},p_{-a}\right)$ such that $y_{ao^{\prime\prime}}=1$ where $R\left(o^{\prime\prime},r_{a}\right)<R\left(\varnothing,r_{a}\right)$.
+Suppose $y\in\mathcal{Y}^{\ast}\left(r_{a},p_{-a}\right)$ such that $y_{ao}=1$ where $R\left(o,r_{a}\right)<R\left(\varnothing,r_{a}\right)$.
 Toward a contradiction, suppose $y\notin\mathcal{Y}^{\ast}\left(d_{a},p_{-a}\right)$.
@@
 We second consider the case where there is no such a deterministic assignment.
-Since $\mathcal{Y}^{\ast}\left(d_{a},p_{-a}\right)\neq\emptyset$ (by Corollary 1), there is $y^{\prime}\in\mathcal{Y}^{\ast}\left(d_{a},p_{-a}\right)$ such that $y_{ao^{\prime}}^{\prime}=1$ for $o^{\prime\prime}$ such that $R\left(o^{\prime\prime},r_{a}\right)\geq R\left(\varnothing,r_{a}\right)$.
+Since $\mathcal{Y}^{\ast}\left(d_{a},p_{-a}\right)\neq\emptyset$ (by Corollary 1), there is $y^{\prime}\in\mathcal{Y}^{\ast}\left(d_{a},p_{-a}\right)$ such that $y_{ao^{\prime}}^{\prime}=1$ for $o$ such that $R\left(o,r_{a}\right)\geq R\left(\varnothing,r_{a}\right)$.
 In this case,
 $RV\left(y,\left(r_{a},p_{-a}\right)\right)=RV\left(y,\left(d_{a},p_{-a}\right)\right)>RV\left(y^{\prime},\left(d_{a},p_{-a}\right)\right).$
-Now, we consider $y^{\prime\prime}$ such that $y_{a\varnothing}^{\prime\prime}=1,$ $y_{ao}^{\prime\prime}=0$ for all $o\in\mathcal{O}\setminus\left\{\varnothing\right\}$ and $y_{a^{\prime}o}^{\prime\prime}=y_{a^{\prime}o}$ for all $a^{\prime}\in\mathcal{A}\setminus\left\{a\right\}$ and $o\in\mathcal{O}$.
+Now, we consider $y^{\prime\prime}$ such that $y_{a\varnothing}^{\prime\prime}=1,$ $y_{a\bar{o}}^{\prime\prime}=0$ for all $\bar{o}\in\mathcal{O}\setminus\left\{\varnothing\right\}$ and $y_{a^{\prime}\bar{o}}^{\prime\prime}=y_{a^{\prime}\bar{o}}$ for all $a^{\prime}\in\mathcal{A}\setminus\left\{a\right\}$ and all $\bar{o}\in\mathcal{O}$.
 Since $q_{\varnothing}\geq\left|\mathcal{A}\right|$, $y^{\prime\prime}$ is an assignment.
-Moreover, since $R\left(o^{\prime\prime},r_{a}\right)\geq R\left(\varnothing,r_{a}\right)$,
+Moreover, since $R\left(o,r_{a}\right)\geq R\left(\varnothing,r_{a}\right)$,
 $RV\left(y^{\prime},\left(d_{a},p_{-a}\right)\right)\geq RV\left(y^{\prime\prime},\left(r_{a},p_{-a}\right)\right)\text{.}$
 Then,
-$RV\left(y^{\ast},\left(r_{a},p_{-a}\right)\right)>RV\left(y^{\prime\prime},\left(r_{a},p_{-a}\right)\right)$
-contradicts $y\in\mathcal{Y}^{\ast}\left(r_{a},p_{-a}\right)$.
-Second, we show the if-part.
-Suppose $y\in\mathcal{Y}^{\ast}\left(d_{a},p_{-a}\right)$.
-We show $y\in\mathcal{Y}^{\ast}\left(r_{a},p_{-a}\right)$.
-Suppose not; that is, there is $y^{\prime}\in\mathcal{Y}$ such that $RV\left(y^{\prime},\left(r_{a},p_{-a}\right)\right)<RV\left(y,\left(r_{a},p_{-a}\right)\right)$.
-Then, $y_{a\bar{o}}^{\prime}=1$ for $R\left(\bar{o},r_{a}\right)\leq R\left(\varnothing,r_{a}\right)$.
-First, we assume $R\left(\bar{o},r_{a}\right)<R\left(\varnothing,r_{a}\right)$.
-Then, $RV\left(y^{\prime},\left(r_{a},p_{-a}\right)\right)=RV\left(y^{\prime},\left(d_{a},p_{-a}\right)\right)$.
-However,
-$RV\left(y^{\prime},\left(d_{a},p_{-a}\right)\right)=RV\left(y^{\prime},\left(r_{a},p_{-a}\right)\right)<RV\left(y,\left(r_{a},p_{-a}\right)\right)=RV\left(y,\left(d_{a},p_{-a}\right)\right),$
-contradicting $y\in\mathcal{Y}^{\ast}\left(d_{a},p_{-a}\right)$.
-Second, we assume $\bar{o}=\varnothing$. Then, since
-$\mathop{\textstyle\sum}\limits_{\bar{o}:R\left(\bar{o},r_{a}\right)<R\left(\varnothing,r_{a}\right)}q_{\bar{o}}+q_{o^{\prime}}\geq\left|\mathcal{A}\right|,$ (11)
-there is $\hat{o}$ such that $R\left(\hat{o},d_{a}\right)\leq R\left(o^{\prime},d_{a}\right)$.
-Thus, for $y^{\prime\prime}\in\mathcal{Y}$ such that $y_{a\hat{o}}^{\prime\prime}=1$ and $y_{a^{\prime}\tilde{o}}^{\prime\prime}=y_{a^{\prime}\tilde{o}}^{\prime}$ for all $a\in\mathcal{A}$ and all $\tilde{o}\in\mathcal{O}\setminus\left\{\hat{o}\right\}$, $RV\left(y^{\prime\prime},\left(d_{a},p_{-a}\right)\right)\leq RV\left(y^{\prime},\left(r_{a},p_{-a}\right)\right)$.
-However,
-$RV\left(y^{\prime\prime},\left(d_{a},p_{-a}\right)\right)\leq RV\left(y^{\prime},\left(r_{a},p_{-a}\right)\right)<RV\left(y,\left(r_{a},p_{-a}\right)\right)=RV\left(y,\left(d_{a},p_{-a}\right)\right),$
-contradicting $y\in\mathcal{Y}^{\ast}\left(d_{a},p_{-a}\right)$.
-Therefore, $y\in\mathcal{Y}^{\ast}\left(d_{a},p_{-a}\right)$, then $y\in\mathcal{Y}^{\ast}\left(r_{a},p_{-a}\right)$. Q.E.D.
+$RV\left(y,\left(r_{a},p_{-a}\right)\right)>RV\left(y^{\prime\prime},\left(r_{a},p_{-a}\right)\right)$
+contradicts $y\in\mathcal{Y}^{\ast}\left(r_{a},p_{-a}\right)$. Q.E.D.
 Claim 2
-Suppose that $y^{\ast}\in\mathcal{Y}^{\ast}\left(r_{a},p_{-a}\right)$ such that $y_{ao}^{\ast}=1$ with $R\left(o,r_{a}\right)<R\left(\varnothing,r_{a}\right)$.
+Suppose that there is $y^{\ast}\in\mathcal{Y}^{\ast}\left(r_{a},p_{-a}\right)$ such that $y_{ao}^{\ast}=1$ and $R\left(o,r_{a}\right)<R\left(\varnothing,r_{a}\right)$.
 Then, $\left|\mathcal{Y}^{\ast}\left(d_{a},p_{-a}\right)\right|\leq\left|\mathcal{Y}^{\ast}\left(r_{a},p_{-a}\right)\right|$.
 Proof.
-First, we show that if $y\in\mathcal{Y}^{\ast}\left(d_{a},p_{-a}\right)$, then either $y\in\mathcal{Y}^{\ast}\left(r_{a},p_{-a}\right)$ or $y^{\prime}\in\mathcal{Y}^{\ast}\left(r_{a},p_{-a}\right),$ such that $y_{a\varnothing}^{\prime}=1$ and $y_{a^{\prime}o}^{\prime}=y_{a^{\prime}o}$ for all $a^{\prime}\in\mathcal{A}\setminus\left\{a\right\}$ and $o\in\mathcal{O}$.
+Suppose that there is $y^{\ast}\in\mathcal{Y}^{\ast}\left(r_{a},p_{-a}\right)$ such that $y_{ao}^{\ast}=1$ and $R\left(o,r_{a}\right)<R\left(\varnothing,r_{a}\right)$.
+Then, by Claim 1, $y^{\ast}\in\mathcal{Y}^{\ast}\left(d_{a},p_{-a}\right)$.
+Next, we show that if $y\in\mathcal{Y}^{\ast}\left(d_{a},p_{-a}\right)$, then either $y\in\mathcal{Y}^{\ast}\left(r_{a},p_{-a}\right)$ or $y^{\prime}\in\mathcal{Y}^{\ast}\left(r_{a},p_{-a}\right),$ such that $y_{a\varnothing}^{\prime}=1$ and $y_{a^{\prime}o}^{\prime}=y_{a^{\prime}o}$ for all $a^{\prime}\in\mathcal{A}\setminus\left\{a\right\}$ and $o\in\mathcal{O}$.
+Since $q_{\varnothing}\geq\left|\mathcal{A}\right|$, $y^{\prime}\in\mathcal{Y}$.
 Suppose that $y\in\mathcal{Y}^{\ast}\left(d_{a},p_{-a}\right)$ and $y\notin\mathcal{Y}^{\ast}\left(r_{a},p_{-a}\right)$. We have
-$RV\left(y,\left(d_{a},p_{-a}\right)\right)\leq RV\left(y^{\ast},\left(d_{a},p_{-a}\right)\right)=RV\left(y^{\ast},\left(r_{a},p_{-a}\right)\right)<RV\left(y,\left(r_{a},p_{-a}\right)\right).$
-Then, by the construction of $d_{a}$, $y_{ao^{\prime}}=1\,$such that $R\left(o^{\prime},r_{a}\right)\geq R\left(\varnothing,r_{a}\right)$.
-Since $q_{\varnothing}\geq\left|\mathcal{A}\right|$, $y^{\prime}$ is an assignment such that $y_{a\varnothing}^{\prime}=1,$ $y_{ao}^{\prime}=0$ for all $o\in\mathcal{O}\setminus\left\{\varnothing\right\}$ and $y_{a^{\prime}o}^{\prime}=y_{a^{\prime}o}$ for all $a^{\prime}\in\mathcal{A}\setminus\left\{a\right\}$ and $o\in\mathcal{O}$.
-Since $R\left(o^{\prime},d_{a}\right)\geq R\left(\varnothing,r_{a}\right)$ (by the construction of $d_{a}$), we have
+$RV\left(y,\left(d_{a},p_{-a}\right)\right)=RV\left(y^{\ast},\left(d_{a},p_{-a}\right)\right)=RV\left(y^{\ast},\left(r_{a},p_{-a}\right)\right)<RV\left(y,\left(r_{a},p_{-a}\right)\right).$
+Then, by the construction of $d_{a}$, $y_{ao^{\prime}}=1\,$such that $R\left(o^{\prime},r_{a}\right)\geq R\left(\varnothing,r_{a}\right)$ and $R\left(o^{\prime},d_{a}\right)\geq R\left(\varnothing,r_{a}\right)$.
+Then, we have
 $RV\left(y^{\prime},\left(r_{a},p_{-a}\right)\right)\leq RV\left(y,\left(d_{a},p_{-a}\right)\right)\leq RV\left(y^{\ast},\left(d_{a},p_{-a}\right)\right)\text{.}$
@@
 contradicting that $y_{2}\in\mathcal{Y}^{\ast}\left(d_{a},p_{-a}\right)$. Therefore, $y_{1}^{\prime}\neq y_{2}^{\prime}$.
-By these results, if $\left\{y_{1},y_{2},\cdots,y_{n}\right\}=\mathcal{Y}^{\ast}\left(d_{a},p_{-a}\right),$ then $\left\{y_{1}^{\prime\prime},y_{2}^{\prime\prime},\cdots,y_{n}^{\prime\prime}\right\}\subseteq\mathcal{Y}^{\ast}\left(r_{a},p_{-a}\right)$ where $y_{i}^{\prime\prime}$ is either $y_{i}$ or $y_{i}^{\prime}$ constructed above, because $y_{1}^{\prime},y_{2}^{\prime},\cdots,y_{n}^{\prime}$ are distinct.
+By these results, if $\left\{y_{1},y_{2},\cdots,y_{n}\right\}=\mathcal{Y}^{\ast}\left(d_{a},p_{-a}\right),$ then $\left\{y_{1}^{\prime\prime},y_{2}^{\prime\prime},\cdots,y_{n}^{\prime\prime}\right\}\subseteq\mathcal{Y}^{\ast}\left(r_{a},p_{-a}\right)$ where $y_{i}^{\prime\prime}$ is either $y_{i}$ or $y_{i}^{\prime}$ constructed above.
+Since $y_{1}^{\prime},y_{2}^{\prime},\cdots,y_{n}^{\prime}$ are distinct, $y_{1}^{\prime\prime},y_{2}^{\prime\prime},\cdots,y_{n}^{\prime\prime}$ are also distinct and thus $\left|\mathcal{Y}^{\ast}\left(d_{a},p_{-a}\right)\right|\leq\left|\mathcal{Y}^{\ast}\left(r_{a},p_{-a}\right)\right|$.
 Q.E.D.
 Now, we show that $a$ weakly prefers $g\left(f^{U}\left(d_{a},p_{-a}\right),r\right)$ to $g\left(f^{U}\left(r_{a},p_{-a}\right),r\right)$. By Claim 2,
-$\left|\mathcal{Y}^{\ast}\left(d_{a},p_{-a}\right)\right|\geq\left|\mathcal{Y}^{\ast}\left(r_{a},p_{-a}\right)\right|\text{.}$
-Moreover, any $y^{\ast}\in\mathcal{Y}^{\ast}\left(r_{a},p_{-a}\right)$ such that $y_{ao}^{\ast}=1$ with $R\left(o,r_{a}\right)<R\left(\varnothing,r_{a}\right)$ is also an element of $\mathcal{Y}^{\ast}\left(d_{a},p_{-a}\right)$.
+$\left|\mathcal{Y}^{\ast}\left(d_{a},p_{-a}\right)\right|\leq\left|\mathcal{Y}^{\ast}\left(r_{a},p_{-a}\right)\right|\text{.}$
+Moreover, by Claim 1, any $y^{\ast}\in\mathcal{Y}^{\ast}\left(r_{a},p_{-a}\right)$ such that $y_{ao}^{\ast}=1$ with $R\left(o,r_{a}\right)<R\left(\varnothing,r_{a}\right)$ is also an element of $\mathcal{Y}^{\ast}\left(d_{a},p_{-a}\right)$.
 Thus, for any $o\in\mathcal{O}$ such that $R\left(o,r_{a}\right)<R\left(\varnothing,r_{a}\right)$,
@@
 Proof of Proposition 3
+Suppose that there is $a\in\mathcal{A}$ whose true preference $r_{a}$ satisfies
+$R\left(o,r_{a}\right)$ $<$ $R\left(\varnothing,r_{a}\right)<R\left(o^{\prime},r_{a}\right)$
+$\sum\limits_{\bar{o}:R\left(\bar{o},r_{a}\right)<R\left(\varnothing,r_{a}\right)}q_{\bar{o}}+q_{o^{\prime}}$$\geq$ $\left|\mathcal{A}\right|,$ (9)
+for some $o,o^{\prime}\in\mathcal{O}$.
 Fix $p_{-a}\in\mathcal{P}^{\left|\mathcal{A}\right|-1}$ arbitrarily.
-Let $r_{a}\in\mathcal{P}$ and $o,o^{\prime}\in\mathcal{O}$ satisfy the condition in Proposition 3.
 Moreover, we let $d_{a}\in\mathcal{D}_{a}$ be such that $R\left(\varnothing,r_{a}\right)=R\left(o^{\prime},d_{a}\right)$.
@@
 Claim 3
-Let $y$ and $y^{\prime}$ be such that $y_{a\varnothing}=y_{ao^{\prime}}^{\prime}=1$ and $y_{a^{\prime}\bar{o}}=y_{a^{\prime}\bar{o}}^{\prime}$ for all $a\in\mathcal{A}$ and all $\bar{o}\in\mathcal{O}\setminus\left\{\varnothing,o^{\prime}\right\}$.
+Let $y$ and $y^{\prime}$ be such that $y_{a\varnothing}=y_{ao^{\prime}}^{\prime}=1$ and $y_{a^{\prime}\bar{o}}=y_{a^{\prime}\bar{o}}^{\prime}$ for all $a^{\prime}\in\mathcal{A}$ and all $\bar{o}\in\mathcal{O}$.
 Then, $y\in\mathcal{Y}^{\ast}\left(r_{a},p_{-a}\right)$ if and only if $y^{\prime}\in\mathcal{Y}^{\ast}\left(d_{a},p_{-a}\right)$.
@@
 First, suppose $y\in\mathcal{Y}^{\ast}\left(r_{a},p_{-a}\right)$.
+Then, by (9), $y^{\prime}$ is an allocation.
 Suppose not; that is, there is $y^{\prime\prime}\in\mathcal{Y}$ such that $RV\left(y^{\prime\prime},\left(d_{a},p_{-a}\right)\right)<RV\left(y^{\prime},\left(d_{a},p_{-a}\right)\right)$.
-Then, since (11) and Lemma 2 are satisfied, $o\in\mathcal{O}$ such that $y_{ao}^{\prime\prime}=1$ and $R\left(o,d_{a}\right)\leq R\left(o^{\prime},d_{a}\right)$.
+Then, by (9), $o^{\prime\prime}$ such that $y_{ao^{\prime\prime}}^{\prime\prime}=1$ satisfies $R\left(o,d_{a}\right)\leq R\left(o^{\prime},d_{a}\right)$.
 Let
-$o^{\prime\prime\prime}=o\text{ if }o\neq o^{\prime},\text{ }o^{\prime\prime\prime}=\varnothing\text{ if }o=o^{\prime}\text{.}$
-Moreover, let $y^{\prime\prime\prime}$ be such that $y_{ao^{\prime\prime\prime}}^{\prime\prime\prime}=1$ and $y_{a^{\prime}o}^{\prime\prime\prime}=y_{a^{\prime}o}^{\prime\prime}$ for any other pairs $(a^{\prime},o)\neq\left(a,o^{\prime\prime\prime}\right)$.
+$o^{\prime\prime\prime}=o^{\prime\prime}\text{ if }o\neq o^{\prime},\text{ }o^{\prime\prime\prime}=\varnothing\text{ if }o^{\prime\prime}=o^{\prime}\text{.}$
+Moreover, let $y^{\prime\prime\prime}$ be such that $y_{ao^{\prime\prime\prime}}^{\prime\prime\prime}=1$ and $y_{a^{\prime}\bar{o}}^{\prime\prime\prime}=y_{a^{\prime}\bar{o}}^{\prime\prime}$ for all $a^{\prime}\in\mathcal{A}$ and all $\bar{o}\in\mathcal{O}$.
 Then,
@@
 Suppose not; that is, there is $y^{\prime\prime}\in\mathcal{Y}$ such that $RV\left(y^{\prime\prime},\left(r_{a},p_{-a}\right)\right)<RV\left(y,\left(r_{a},p_{-a}\right)\right)$.
-Then, since (11) and Lemma 2 are satisfied, $o\in\mathcal{O}$ such that $y_{ao}^{\prime\prime}=1$ and $R\left(o,r_{a}\right)\leq R\left(\varnothing,r_{a}\right)$.
+Then, by (9), $o^{\prime\prime}\in\mathcal{O}$ such that $y_{ao^{\prime\prime}}^{\prime\prime}=1$ and $R\left(o^{\prime\prime},r_{a}\right)\leq R\left(\varnothing,r_{a}\right)$.
 Let
-$o^{\prime\prime\prime}=o\text{ if }o\neq\varnothing,\text{ }o^{\prime\prime\prime}=\bar{o}\text{ if }o=\varnothing\text{,}$
-where $\bar{o}$ is such that $\sum y_{a\bar{o}}^{\prime\prime}<q_{\bar{o}}$ and $R\left(\bar{o},d_{a}\right)\leq R\left(o^{\prime},d_{a}\right)$.
+$o^{\prime\prime\prime}=o^{\prime\prime}\text{ if }o^{\prime\prime}\neq\varnothing,\text{ }o^{\prime\prime\prime}=\bar{o}\text{ if }o^{\prime\prime}=\varnothing\text{,}$
+where $\bar{o}$ satisfies $\sum y_{a\bar{o}}^{\prime\prime}<q_{\bar{o}}$ and $R\left(\bar{o},d_{a}\right)\leq R\left(o^{\prime},d_{a}\right)$.
 There exist such an $\bar{o}$, because (9) is satisfied.
-Then, let $y^{\prime\prime\prime}$ be such that $y_{ao^{\prime\prime\prime}}^{\prime\prime\prime}=1$ and $y_{a^{\prime}o}^{\prime\prime\prime}=y_{a^{\prime}o}^{\prime\prime}$ for any other pairs $(a^{\prime},o)\neq\left(a,o^{\prime\prime\prime}\right)$.
+Then, let $y^{\prime\prime\prime}$ be such that $y_{ao^{\prime\prime\prime}}^{\prime\prime\prime}=1$ and $y_{a^{\prime}\bar{o}}^{\prime\prime\prime}=y_{a^{\prime}\bar{o}}^{\prime\prime}$ for all $a^{\prime}\in\mathcal{A}$ and all $\bar{o}\in\mathcal{O}$.
 Then,
@@
 First, suppose that $p$ satisfies all of Conditions (i), (ii) and (iii).
-Then, for any two agents $\hat{a}$ and $\bar{a}$ belonging to different groups, $p_{\hat{a}}$ and $p_{\bar{a}}$ are not essentially the same.
-Moreover, for any two agents $\hat{a}$ and $\bar{a}$ belonging to the second group or third group, $\left(f^{M}\left(p\right)\right)_{\hat{a}o^{\prime}}=\left(f^{M}\left(p\right)\right)_{\bar{a}o^{\prime}}$ for all $o^{\prime}\in\mathcal{O}$.
+Then, for any two agents $a_{1}$ and $a_{2}$ belonging to different groups, $p_{a_{1}}$ and $p_{a_{2}}$ are not essentially the same.
+Moreover, for any two agents $a_{1}$ and $a_{2}$ belonging to the same group (either the second group or the third group), $\left(f^{M}\left(p\right)\right)_{a_{1}o^{\prime}}=\left(f^{M}\left(p\right)\right)_{a_{2}o^{\prime}}$ for all $o^{\prime}\in\mathcal{O}$.
 Second if $p$ does not satisfies all of Conditions (i), (ii) and (iii), then $f^{M}\left(p\right)=f^{U}\left(p\right)$. By Proposition 2, $f^{M}$ satisfies ETE.
@@
 Let $k$ be the smallest integer such that $p_{a}^{k}\neq r_{a}^{k}$.
-First, suppose that (3) is satisfied for that $k$ and $r_{a}^{k}\neq\varnothing$.
-Since (3) is satisfied for $k$, $r_{a}^{k^{\prime}}\neq\varnothing$ for all $k^{\prime}=1,\cdots,k-1$; that is, $R(\varnothing,r_{a})>k$ and $R(\varnothing,p_{a})\geq k$.
+First, suppose $k\leq\bar{k}\left(p_{a}\right)$ and $r_{a}^{k}\neq\varnothing$.
+Since $k\leq\bar{k}\left(p_{a}\right)$, $r_{a}^{k^{\prime}}\neq\varnothing$ for all $k^{\prime}=1,\cdots,k-1$; that is, $R(\varnothing,r_{a})>k$ and $R(\varnothing,p_{a})\geq k$.
 In this case, since $f^{M}$ satisfies ETE (by Proposition 5), we revisit the examples in the first result of Lemma 3.
@@
 Since $R(\varnothing,r_{a})>k$ and $R(p_{a}^{k},r_{a})>k$, in either case, $a$ prefers $g\left(f^{M}\left(r_{a},p_{-a}\right),r\right)$ to $g\left(f^{M}\left(p_{a},p_{-a}\right),r\right)$.
-Second, suppose (3) is satisfied and $r_{a}^{k}=\varnothing$.
-Then, $k\geq 2$.
+Second, suppose $k\leq\bar{k}\left(p_{a}\right)$ and $r_{a}^{k}=\varnothing$.
+Then, since we assume $r_{a}^{1}\neq\varnothing$, $k\geq 2$.
 Moreover, in this case, $p_{a}^{k^{\prime}}\neq\varnothing$ for all $k^{\prime}=1,\cdots,k$.
@@
 Let $p_{-a}$ be such that there are $q_{o}$ students represented by $a^{\prime}$ whose revealed preference is equivalent to $r_{a}(=p_{a^{\prime}})$ and the first best type of the other agents represented by $a^{\prime\prime}$ is $\varnothing$.
-Then, $\left(p_{a},p_{-a}\right)$ satisfies (i), (ii) and (iii) with $l=k$, because $k$ satisfies (3).
+Then, $\left(p_{a},p_{-a}\right)$ satisfies (i), (ii) and (iii) with $l=k$, because $k\leq\bar{k}\left(p_{a}\right)$.
 Moreover, $\left(r_{a},p_{-a}\right)$ does not satisfy all of them.
 Therefore, in this case, $g\left(f^{M}\left(r_{a},p_{-a}\right)\right)$ is more preferable for $a$ than $g\left(f^{M}\left(p_{a},p_{-a}\right)\right)$.
-Third, suppose that $k$ does not satisfy (3); that is,
-$\mathop{\displaystyle\sum}\limits_{\bar{o}:R(\bar{o},p_{a})<k}q_{\bar{o}}=\mathop{\displaystyle\sum}\limits_{\bar{o}:R(\bar{o},r_{a})<k}q_{\bar{o}}\geq\left|\mathcal{A}\right|$
-In this case, $r_{a}$ and $p_{a}$ are essentially the same. Since $f^{M}$ is a fair rank-minimizing mechanism (by Proposition 5),
+Third, suppose $k>\bar{k}\left(p_{a}\right)$. In this case, $r_{a}$ and $p_{a}$ are essentially the same. Since $f^{M}$ satisfies ETE (by Proposition 5),
 $g\left(f^{M}\left(r_{a},p_{-a}\right),r\right)=g\left(f^{M}\left(p_{a},p_{-a}\right),r\right)$

### v4

@@
 In rank-minimizing mechanisms, assignments are determined as to minimize the total (or average) revealed rank of the objects to which the agents are assigned.
-The rank distributions of the assignments are seriously considered by many real world matching authorities.
-For example, Featherstone (2020) mentions that Teach for America considers the rank distribution when selecting its assignment.
-Therefore, the rank-minimizing mechanisms are focused on by several recent studies such as Featherstone (2020), Nikzad (2022), Ortega and Klein (2023) and Troyan (2024).
-Some of them highlight several advantages of the mechanism with regard to the efficiency of assignment.
-Featherstone (2020) shows that any rank-minimizing mechanism satisfies another property of efficiency called ordinal efficiency, first introduced by Bogomolnaia and Moulin (2001).
-Moreover, Ortega and Klein (2023) compare the result of a rank-minimizing mechanism and those of the DA mechanism and the TTC mechanism, and show that the result of a rank-minimizing mechanism improves upon those of others in several terms of efficiency.
+The rank distributions of assignments are a critical consideration for many real-world matching authorities.
+For instance, Featherstone (2020) notes that Teach For America takes rank distribution into account when making assignment decisions.
+Consequently, rank-minimizing mechanisms are discussed in recent studies, including Featherstone (2020), Nikzad (2022), Ortega and Klein (2023), Feizi (2024), and Troyan (2024).
+Several of these studies highlight the advantages of rank-minimizing mechanisms, particularly regarding the efficiency of assignments.
+By the results of Featherstone (2020) and Feizi (2024), any rank-minimizing mechanism satisfies other efficiency properties, such as ordinal efficiency, which is first introduced by Bogomolnaia and Moulin (2001).
+Moreover, Ortega and Klein (2023) compare the outcomes of rank-minimizing mechanisms with those of the DA mechanism and the TTC mechanism, showing that rank-minimizing mechanisms improve upon the others in several aspects of efficiency.
 In this study, we especially consider the rank-minimizing mechanisms that satisfy a fairness property called equal treatment of equals, which requires that if two agents reveal the same preference ranking, then they are assigned to each object type with the same probability.
@@
 These results imply that in this case, strategic manipulation may not be a critical problem of rank-minimizing mechanisms.
-Next, we consider the case where agents can refuse their assignment and instead obtain the outside option.
-Featherstone (2020) firstly considers this situation, but focuses only on the case where each agent has limited information on both the preferences of other agents and the capacities of all object types except the outside option.
+Next, as the main analysis, we consider the case where agents have the refusal options.
+[^1] Feigenbaum et al.
+(2020) and Afacan (2022) state that in some real school choice markets, such as public school markets in Boston, New York City, and Turkey, students actually refuse their assignments.
+For example, about 10% of the students withdrew from the New York City public high school placement.
+Featherstone (2020) theoretically considers the situation, but focuses only on the case where each agent has limited information on both the preferences of other agents and the capacities of all object types except the outside option.
 In contrast, we assume that each agent has complete information on the capacities of all object types.
-For example, in a school choice market, information about all schools, including their capacities, is assumed to be disclosed to all students.
-Additionally, we do not need any assumptions about the information structure of agents regarding the preferences and rationality of other agents because we focus only on strategic dominance relations.
+For example, we consider a school choice market where information about all schools, including their capacities, is disclosed to all students.
+We do not need any assumptions about the information structure of agents regarding the preferences and rationality of other agents, because we focus only on strategic dominance relations.
 We consider an outside option demotion strategy, which involves revealing a ranking where the rank of the outside option is placed at the bottom, without changing the order of the truly acceptable object types.
@@
 We call this mechanism the modified uniform rank-minimizing mechanism, which is shown to be fair.
-That is, to ensure that any outside option demotion strategy can fail, in the mechanism, increasing acceptable object types can make the agent less likely to be assigned their acceptable types.
+That is, to ensure that any outside option demotion strategy can fail in some case, in the mechanism, increasing acceptable object types can make the agent less likely to be assigned their acceptable types.
 However, using the modified uniform rank-minimizing mechanism may create another strategic issue: an agent may have an incentive to reduce the number of truly acceptable object types.
@@
 That is, the null object type can be assigned to all agents simultaneously.
-The null object type $\varnothing\in\mathcal{O}$ can be interpreted as the (best) outside option of an agent.
+The null object type $\varnothing\in\mathcal{O}$ can be interpreted as the first-best alternative object of an agent outside of this market.
 In Section 4, we consider the situation where the null object type has the other characteristic.
@@
 Let $\mathcal{P}$ be the set of all possible strict rankings (linear orders) of the elements in $\mathcal{O}$.
-Let $p_{a}=\left(p_{a}^{1},\cdots,p_{a}^{\left|\mathcal{O}\right|}\right)\in\mathcal{P}^{\left|\mathcal{A}\right|}$ be a preference order revealed by $a$, where $p_{a}^{k}=o$ means that type $o$ object is ranked $k$th best one under $p_{a}$.
+Let $p_{a}=\left(p_{a}^{1},\cdots,p_{a}^{\left|\mathcal{O}\right|}\right)\in\mathcal{P}^{\left|\mathcal{A}\right|}$ be a preference order revealed by $a$, where $p_{a}^{k}=o$ means that type $o$ object is ranked $k$th-best one under $p_{a}$.
 For notational convenience, we let $R:\mathcal{O\times P}\rightarrow\left\{1,\cdots,\left|\mathcal{O}\right|\right\}$ be a rank function; that is, $p_{a}^{k}=o$ if and only if $R\left(o,p_{a}\right)=k$.
@@
 for all $k=1,2,\cdots,\left|\mathcal{O}\right|$, then agent $a$ with $p_{a}$ is said to weakly prefer (in the sense of first-order stochastic dominance) $x$ to $x^{\prime}$.
-Moreover, $a$ is said to strictly prefers $x$ to $x^{\prime}$ if $a$ weakly prefers $x$ to $x^{\prime}$ and (1) is satisfied with inequality for some $k=1,2,\cdots,\left|\mathcal{O}\right|-1$.
-Assignment $x$ weakly ordinally dominates assignment $x^{\prime}$ for $p$ if each agent $a$ with $p_{a}$ weakly prefers $x$ to $x^{\prime}$.
-If $x$ weakly ordinally dominates assignment $x^{\prime}$ for $p$ and (1) is satisfied with inequality for at least one agent–rank pair, then $x$ ordinally dominates assignment $x^{\prime}$ for $p$.
-An assignment that is not ordinally dominated by any other assignments is said to be ordinally efficient for $p$.
+Moreover, $a$ is said to strictly prefer $x$ to $x^{\prime}$ if $a$ weakly prefers $x$ to $x^{\prime}$ and (1) is satisfied with inequality for some $k=1,2,\cdots,\left|\mathcal{O}\right|-1$.
 Assignment $x$ is wasteful for $p$ if there are $o,o^{\prime}\in\mathcal{O}$ and $a\in\mathcal{A}$ such that $\sum\nolimits_{a^{\prime}\in\mathcal{A}}x_{a^{\prime}o}<q_{o}$, $x_{ao^{\prime}}>0$ and $R\left(o,p_{a}\right)<R\left(o^{\prime},p_{a}\right)$.
-Any wasteful assignment for $p$ must not be ordinally efficient.
-[^1] Moreover, a wasteful assignment for $p$ must not be efficient in a weaker efficiency concept such as ex-post efficiency, which is discussed several previous studies such as Bogomolnaia and Moulin (2001).
-Thus, not being wasteful is a very weak requirement for efficiency.
 For an assignment $x\in\mathcal{X}$ and a preference profile $p=\left(p_{a}\right)_{a\in\mathcal{A}}\in\mathcal{P}^{\left|\mathcal{A}\right|}$, let $RV\left(x,p\right)$ be such that
@@
 which is the sum of expected rank.
-An assignment $x^{\ast}\in\mathcal{X}$ is rank efficient for $p$ if $RV\left(x^{\ast},p\right)\leq RV\left(x,p\right)$ for all $x\in\mathcal{X}$.
-Let $\mathcal{X}^{\ast}\left(p\right)$ be the set of rank efficient assignments for $p$.
-The following result is shown by Featherstone (2020).
-Remark 2
-Any rank efficient assignment for $p$ is ordinally efficient for $p$.
-In the subsequent section, we focus on the mechanisms whose input is $p\in\mathcal{P}^{\left|\mathcal{A}\right|}$ and output is rank efficient for any $p\in\mathcal{P}^{\left|\mathcal{A}\right|}$.
+An assignment $x^{\ast}\in\mathcal{X}$ is rank-minimizing for $p$ if $RV\left(x^{\ast},p\right)\leq RV\left(x,p\right)$ for all $x\in\mathcal{X}$.
+Let $\mathcal{X}^{\ast}\left(p\right)$ be the set of rank-minimizing assignments for $p$.
+If $x^{\ast}\in\mathcal{X}$ is rank-minimizing for $p$, then $x$ is not wasteful for $p$.
+The efficiency property of ”rank-minimizing” is stronger than several efficiency properties, such as rank efficiency and ordinal efficiency, analyzed in this literature (see Figure 1 in Feizi (2024)).
+Among them, not being wasteful is a very weak property.
+In the subsequent section, we focus on the mechanisms whose input is $p\in\mathcal{P}^{\left|\mathcal{A}\right|}$ and output is rank-minimizing for any $p\in\mathcal{P}^{\left|\mathcal{A}\right|}$.
 Before that we introduce some technical results.
@@
 $x^{\ast}=\sum\limits_{i=1}^{I}\sigma^{i}y^{i}\in\mathcal{X}\text{ where }\sum\limits_{i=1}^{I}\sigma^{i}=1$
-and $\sigma^{i}\in\left[0,1\right]$ for all $i=1,\cdots,I$. If $x^{\ast}$ is rank efficient, then $y^{i}$ is also rank efficient for all $i=1,\cdots,I$.
+and $\sigma^{i}\in\left[0,1\right]$ for all $i=1,\cdots,I$. If $x^{\ast}$ is rank-minimizing, then $y^{i}$ is also rank-minimizing for all $i=1,\cdots,I$.
 We provide the proofs of our results in the Appendix.
@@
 Corollary 1
-There must exist at least one deterministic rank efficient assignment.
-Nikzad (2022) and Troyan (2024) consider a deterministic rank efficient assignment; that is, they focus on a deterministic assignment $y^{\ast}\in\mathcal{Y}$ such that $RV\left(y^{\ast},p\right)\leq RV\left(y,p\right)$ for all $y\in\mathcal{Y}$.
-However, by Corollary 1, such $y^{\ast}$ must also be rank efficient in our definition; that is, $RV\left(y^{\ast},p\right)\leq RV\left(x,p\right)$ for all $x\in\mathcal{X}$.
+There must exist at least one deterministic rank-minimizing assignment.
+Nikzad (2022) and Troyan (2024) consider a deterministic rank-minimizing assignment; that is, they focus on a deterministic assignment $y^{\ast}\in\mathcal{Y}$ such that $RV\left(y^{\ast},p\right)\leq RV\left(y,p\right)$ for all $y\in\mathcal{Y}$.
+However, by Corollary 1, such $y^{\ast}$ must also be rank-minimizing in our definition; that is, $RV\left(y^{\ast},p\right)\leq RV\left(x,p\right)$ for all $x\in\mathcal{X}$.
 Moreover, this result allows us to show that the mechanism introduced later is well-defined.
@@
 First, we consider an efficiency property on mechanisms.
-We call $f$ a rank-minimizing mechanism if $f(p)$ is rank efficient for all $p$.
+We call $f$ a rank-minimizing mechanism if $f(p)$ is rank-minimizing for all $p$.
 Second, we introduce a fairness property.
@@
 Unlike in our model, they take into account the priorities of objects and define agents as equal only if they are equal with respect to the priorities of all objects.
+Other fairness properties discussed in the literature include weak envy-freeness and the equal division lower bound.
+However, as shown by Feizi (2024), both of these properties are incompatible with rank-minimizing.
+Since ETE is weaker than these fairness properties, it is compatible with rank-minimizing and is therefore adopted in this study.
 By Lemma 2, as long as the assignment is not wasteful, any type that is ranked lower than $\bar{k}\left(p_{a}\right)$th in the preference ranking revealed by $a$ will not be assigned to $a$, because there are sufficient copies of types ranked higher than or equal to $\bar{k}\left(p_{a}\right)$th.
 Thus, if two individuals reveal rankings that are identical for positions at or above $\bar{k}\left(p_{a}\right)$th rank, those rankings are considered essentially the same, and they should receive the same assignment in an ex-ante sense.
-We focus on the efficient and fair mechanisms; that is, we call $f$ a fair rank-minimizing mechanism if $f(p)$ is rank efficient for all $p$ and satisfies ETE. The following mechanism is one of them.
-Let $\mathcal{Y}^{\ast}\left(p\right)\subseteq\mathcal{Y}$ be the set of deterministic rank efficient assignments; that is, $y^{\ast}\in\mathcal{Y}^{\ast}\left(p\right),$ $RV\left(y^{\ast},p\right)\leq RV\left(x,p\right)$ for any $x\in\mathcal{X}$.
+We focus on the efficient and fair mechanisms; that is, we call $f$ a fair rank-minimizing mechanism if $f(p)$ is rank-minimizing for all $p$ and satisfies ETE. The following mechanism is one of them.
+Let $\mathcal{Y}^{\ast}\left(p\right)\subseteq\mathcal{Y}$ be the set of deterministic rank-minimizing assignments; that is, $y^{\ast}\in\mathcal{Y}^{\ast}\left(p\right),$ $RV\left(y^{\ast},p\right)\leq RV\left(x,p\right)$ for any $x\in\mathcal{X}$.
 By Corollary 1, $\mathcal{Y}^{\ast}\left(p\right)\neq\emptyset$.
@@
 $RV\left(f^{U}\left(p\right),p\right)=\sum\limits_{y^{\ast}\in\mathcal{Y}^{\ast}\left(p\right)}\frac{1}{\left|\mathcal{Y}^{\ast}\left(p\right)\right|}RV\left(y^{\ast},p\right).$
-Thus, $RV\left(f^{U}\left(p\right),p\right)\leq RV\left(x,p\right)$ for any $x\in\mathcal{X}$; that is, $f^{U}\left(p\right)$ must be rank efficient for $p$.
+Thus, $RV\left(f^{U}\left(p\right),p\right)\leq RV\left(x,p\right)$ for any $x\in\mathcal{X}$; that is, $f^{U}\left(p\right)$ must be rank-minimizing for $p$.
 Since Nikzad (2022), Ortega and Klein (2023) and Troyan (2024) also discuss the uniform rank-minimizing mechanism, this is one of the most common rank-minimizing mechanisms in the literature.
@@
 In other word, we consider the case where the number of object types that are more preferable than the $k-1$th best one in $p_{a}$ is insufficient to distribute to all agents.
-In this case, for some $p_{-a}\in\mathcal{P}^{\left|\mathcal{A}\right|-1}$ and some fair rank-minimizing mechanism, $a$ is assigned to an object type that is less preferable or equal to the $k$th best one.
+In this case, for some $p_{-a}\in\mathcal{P}^{\left|\mathcal{A}\right|-1}$ and some fair rank-minimizing mechanism, $a$ is assigned to an object type that is less preferable or equal to the $k$th-best one.
 Lemma 3
@@
 if $\left(p_{a_{1}},p_{a_{2}},p_{a_{3}}\right)$ is neither $\left(\psi^{\prime},\psi,\psi\right)$ nor $\left(\psi,\psi^{\prime},\psi\right)$ nor $\left(\psi,\psi,\psi^{\prime}\right)$.
-That is, if the best type and the second best type of the three agents are respectively $o_{1}$ and $o_{2}$ and the third best type of the two of them is $\varnothing$, then the agent revealing $o_{3}$ as the third best one obtains $o_{1}$ with probability 00.
+That is, if the first-best type and the second-best type of the three agents are respectively $o_{1}$ and $o_{2}$ and the third-best type of the two of them is $\varnothing$, then the agent revealing $o_{3}$ as the third-best one obtains $o_{1}$ with probability 00.
 Trivially, $f$ satisfies weak ETE.
@@
 Thus, in this case, $r_{a_{1}}=\psi^{\prime}$ is strategically dominated by $\psi$.
-Therefore, even if a fair rank-minimizing mechanism is used, truth-telling may be strategically dominated by some other strategy.
+Therefore, even if a rank-minimizing mechanism is used, truth-telling may be strategically dominated by some other strategy.
 Moreover, in this case, $f$ does not satisfy ETE.
@@
 In this section, we consider the following three-stage game, which is a modification of that in Section 3; that is, the first and second stages in this game are the same as in the previous game.
-In the third stage, all agents decide whether accept or refuse the object type that is assigned to them.
+In the third stage, all agents decide whether to accept or refuse the assigned object type.
 If an agent refuses their assignment, then the agent instead obtains the null object type $\varnothing$ (the outside option).
-In this game, the null object type has two characteristics: first, there are sufficiently many copies of this type to allocate to all individuals, and second, agents can always obtain it at no cost as long as they abandon their assignment.
+That is, in this game, the null object type has two characteristics: first, there are sufficiently many copies of this type to allocate to all individuals, and second, agents can always obtain it at no cost as long as they abandon their assignment.
 This third stage is also considered by Featherstone (2020) but is not explicitly addressed by almost previous studies related to ours.
-[^3] Doğan and Yenmez (2019) also explicitly consider this stage.
+[^3] As exceptions, Feigenbaum et al.
+(2020) and Afacan (2022) consider an assignment problem where, after the initial allocation is decided, if someone refuses the assigned object, it is reallocated.
+Moreover, Doğan and Yenmez (2019) also explicitly consider this stage.
 However, this stage is implicitly considered by many of them; that is, they consider an ex-post individually rational (or more strictly an ex-post efficient) matching as a requirement and if an assignment is ex-post individually rational, no agent refuses their assignment.
@@
 Featherstone (2020) defines the full extension of $r_{a}$denoted by $e_{a}$ in this study such that (i) $R\left(o,r_{a}\right)=R\left(o,e_{a}\right)$ for all $o\in\mathcal{O}$ such that $R\left(o,r_{a}\right)<R\left(\varnothing,r_{a}\right)$, and (ii) $R\left(\varnothing,e_{a}\right)=\left|\mathcal{O}\right|$, and $R\left(o,r_{a}\right)-1=R\left(o,e_{a}\right)$ for all $o\in\mathcal{O}$ such that $R\left(o,r_{a}\right)>R\left(\varnothing,r_{a}\right)$; that is, $e_{a}\in\mathcal{D}_{a}$ is such that all non-null object types are placed in their true preference orders.
+Revealing the full extension of $r_{a}$ is one of the ODSs of $a$.
 First, we consider the following example to briefly consider the main result.
@@
 $f^{U}\left(d_{a_{1}},p_{a_{2}},p_{a_{3}}\right)=\left(\begin{array}[]{ccc}1/2&0&1/2\\ 1/2&0&1/2\\ 0&1&0\end{array}\right)$
-Thus, revealing $d_{a_{1}}$ instead of $r_{a_{1}}$ may increase the probability that $a_{1}$ is assigned to their best object $o_{1}$. On the other hand, suppose
+Thus, revealing $d_{a_{1}}$ instead of $r_{a_{1}}$ may increase the probability that $a_{1}$ is assigned to their first-best object $o_{1}$. On the other hand, suppose
 $p_{a_{2}}^{\prime}=(o_{1},\varnothing,o_{2})\text{ and }p_{a_{3}}^{\prime}=(o_{1},\varnothing,o_{2}).$
@@
 and $\text{and }g\left(f^{U}\left(d_{a_{1}},\hat{p}_{a_{2}},\hat{p}_{a_{3}}\right),r\right)_{a^{\prime}\varnothing}$ $=$ $\frac{1}{3}>0,$
-for $a^{\prime}=a_{2},a_{3}$. Thus, it is ordinally dominated by
-$\tilde{x}=\left(\begin{array}[]{ccc}1/3&0&2/3\\ 1/3&1/2&1/6\\ 1/3&1/2&1/6\end{array}\right).$
+for $a^{\prime}=a_{2},a_{3}$. Thus, if
+$\tilde{x}=\left(\begin{array}[]{ccc}1/3&0&2/3\\ 1/3&1/2&1/6\\ 1/3&1/2&1/6\end{array}\right),$
+then $a_{2}$ and $a_{3}$ strictly prefer and $a_{1}$ weakly prefers $\tilde{x}$ to $g\left(f^{U}\left(d_{a_{1}},\hat{p}_{a_{2}},\hat{p}_{a_{3}}\right),r\right)$.
 In the case where $a_{1}$ is assigned to $o_{2}$, either $a_{2}$ or $a_{3}$ is assigned to $\varnothing$ but $a_{1}$ will refuse $o_{2}$.
@@
 By this result, an agent is never hurt by using an ODS instead of truth-telling.
-Next, we show that with some conditions, there exists some ODS such that an agent may be better off by using it instead of truth-telling.
+Next, we show that with some conditions, there is some ODS such that an agent may be better off by using it instead of truth-telling.
 Theorem 2
@@
 Thus, for all $a^{\prime}$, $p_{a^{\prime}}$ may also be an ODS of $a^{\prime}$.
-Theorems 1 and 2 do not say that an ODS is a dominant strategy.
-However, when if there is $a\in\mathcal{A}$ whose true preference satisfies (3) for some $o,o^{\prime}\in\mathcal{O}$, since truth-telling is strategically dominated by another strategy, $a$ has a strong incentive to manipulate its preference ranking, while one of the ODSs may not necessarily be chosen.
+Theorems 1 and 2 do not imply that an ODS is a dominant strategy.
+However, if there is $a\in\mathcal{A}$ whose true preference satisfies (3) for some $o,o^{\prime}\in\mathcal{O}$, since truth-telling is strategically dominated by another strategy, $a$ has a strong incentive to manipulate its preference ranking, while one of the ODSs may not necessarily be chosen.
 As stated earlier, there may exist multiple ODSs. In that case, an ODS may be strategically dominated by another one. We have the following result.
@@
 That is, the copies of $o_{1}$ and $o_{2}$ are sufficient to allocate all agents, but those of $o_{1}$ and $o_{3}$ are insufficient.
-Thus, revealing $o_{2}$ as a second best type does not increase the probability that $a_{1}$ is assigned to $o_{1}$.
-On the other hand, revealing $o_{3}$ as a second best type increases it, when $o_{3}$ is the first best type of another agent.
+Thus, revealing $o_{2}$ as a second-best type does not increase the probability that $a_{1}$ is assigned to $o_{1}$.
+On the other hand, revealing $o_{3}$ as a second-best type increases it, when $o_{3}$ is the first-best type of another agent.
 Thus, revealing the full extension may be strategically dominated by another ODS when an unacceptable type (in this case $o_{3}$) is less preferable than the other (in this case $o_{2}$) but the number of copies of the former is less than that of the latter.
@@
 $\sum\limits_{\bar{o}:R\left(\bar{o},p_{a^{\prime}}\right)<R\left(\varnothing,p_{a^{\prime}}\right)}q_{\bar{o}}<\left|\mathcal{A}\right|,$
-if $\left|\mathcal{A}\right|>q_{o}+1$, then the first best type of the other agents ($\left|\mathcal{A}\right|-\left(q_{o}+1\right)$ agents) represented by $a^{\prime\prime}$ is $\varnothing$, then
+if $\left|\mathcal{A}\right|>q_{o}+1$, then the first-best type of the other agents ($\left|\mathcal{A}\right|-\left(q_{o}+1\right)$ agents) represented by $a^{\prime\prime}$ is $\varnothing$, then
 $\left(f^{M}\left(p\right)\right)_{ao}=0,\text{ }\left(f^{M}\left(p\right)\right)_{a^{\prime}o}=1,\text{ }\left(f^{M}\left(p\right)\right)_{ap_{a}^{2}}=1\text{ and }\left(f^{M}\left(p\right)\right)_{a^{\prime\prime}\varnothing}=1,$
@@
 The strategic issues, mentioned in Theorem 2, do not arise under this mechanism $f^{M}$.
-First, we explain why the strategic issues mentioned in Example 2 does not arise in the modified uniform rank-minimizing mechanism.
+First, we explain why the strategic issue mentioned in Example 2 does not arise in the modified uniform rank-minimizing mechanism.
 If $r_{a}=p_{a^{\prime}}$ where $p_{a^{\prime}}$ is defined in (ii), then $\left(d_{a},p_{-a}\right)$ satisfies (i), (ii) and (iii) where $d_{a}\in\mathcal{D}_{a}$.
@@
 Let $R\left(\varnothing,r_{a}\right)=k$.
-Suppose $k\geq 3$; that is, $\varnothing$ is neither the best type nor the second best type for $r_{a}$.
+Suppose $k\geq 3$; that is, $\varnothing$ is neither the first-best type nor the second-best type for $r_{a}$.
 Thus, we let $r_{a}^{1}=o$ and $r_{a}^{2}=o^{\prime}$.
@@
 Second, we consider efficient and fair mechanisms other than the fair rank-minimizing mechanisms, such as the (general) probabilistic serial mechanism introduced by Bogomolnaia and Moulin (2001) and Budish et al.
-(2013), because it is fair and ordinally efficient, although as stated earlier, this efficiency property is weaker than the rank efficiency.
+(2013), because it is fair and ordinally efficient, although this efficiency property is weaker than the rank-minimizing.
 Appendix
@@
 $=$ $\sum\limits_{i=1}^{I}\sigma^{i}RV\left(y^{i},p\right).$
-Thus, $RV\left(y^{i},p\right)>RV\left(x^{\ast},p\right)$ implies that there is $y^{j}$ such that $RV\left(y^{j},p\right)<RV\left(x^{\ast},p\right)$, which contradicts that $x^{\ast}$ is a rank efficient assignment.
+Thus, $RV\left(y^{i},p\right)>RV\left(x^{\ast},p\right)$ implies that there is $y^{j}$ such that $RV\left(y^{j},p\right)<RV\left(x^{\ast},p\right)$, which contradicts that $x^{\ast}$ is a rank-minimizing assignment.
 Proof of Lemma 2
@@
 Note that $a$ itself belongs to this group.
-For the second group, the best object type is $\varnothing$ under $\left(r_{a},p_{-a}\right)$.
+For the second group, the first-best object type is $\varnothing$ under $\left(r_{a},p_{-a}\right)$.
 Then, since all agents in the second group always assign $\varnothing$ with probability $1$ and $f$ satisfies ETE,
@@
 Next, we consider $j$th group for each $j=2,3,\cdots,k-1$.
-In $j$th group, there are just $q_{r_{a}^{j}}$ agents whose best object type is $r_{a}^{j}$ and the second best object is $r_{a}^{j+1},$ the third best $r_{a}^{j+2},$cdots and the $k$th best $r_{a}^{k}$ under $\left(r_{a},p_{-a}\right)$.
+In $j$th group, there are just $q_{r_{a}^{j}}$ agents whose first-best object type is $r_{a}^{j}$ and the second-best object is $r_{a}^{j+1},$ the third-best $r_{a}^{j+2},$cdots and the $k$th-best $r_{a}^{k}$ under $\left(r_{a},p_{-a}\right)$.
 If $a_{j}$ belongs $j$th group for $j=1,2,\cdots,k-1$, the preference orders of them are as follows:
@@
 and thus $\left(r_{a},p_{-a}\right)$ is well-defined so far.
-Finally, for all the remaining agents (belonging to the $k$th group), the best object type is $\varnothing$ under $\left(r_{a},p_{-a}\right)$.
-In this case, at least one agent cannot be assigned to their best object.
-Thus, in a rank-minimizing assignment, all agents in $j$th group for all $j=2,3,\cdots,k-1$ is assigned to their best object type $r_{a}^{j}$, and $q_{r_{a}^{1}}$ agents in the first group is also assigned to their best object type $r_{a}^{1},$ and moreover, just one agent in the first group is also assigned to $r_{a}^{k}$.
-Moreover, an agent who belongs to the $k$th group is assigned to their best object type $\varnothing$.
-Since all agents excepting those belonging to the first group are assigned to their best alternative with probability $1$ and $f^{U}$ satisfies ETE,
+Finally, for all the remaining agents (belonging to the $k$th group), the first-best object type is $\varnothing$ under $\left(r_{a},p_{-a}\right)$.
+In this case, at least one agent cannot be assigned to their first-best object.
+Thus, in a rank-minimizing assignment, all agents in $j$th group for all $j=2,3,\cdots,k-1$ is assigned to their first-best object type $r_{a}^{j}$, and $q_{r_{a}^{1}}$ agents in the first group is also assigned to their first-best object type $r_{a}^{1},$ and moreover, just one agent in the first group is also assigned to $r_{a}^{k}$.
+Moreover, an agent who belongs to the $k$th group is assigned to their first-best object type $\varnothing$.
+Since all agents excepting those belonging to the first group are assigned to their first-best alternative with probability $1$ and $f^{U}$ satisfies ETE,
 $\left(f\left(r_{a},p_{-a}\right)\right)_{ar_{a}^{1}}$ $=$ $\frac{q_{r_{a}^{1}}}{q_{r_{a}^{1}}+1},$
@@
 Therefore, $R(\varnothing,p_{a})>R(\varnothing,r_{a})=k$.
-Let $p_{-a}$ be such that there are $q_{o}$ students represented by $a^{\prime}$ whose revealed preference is equivalent to $r_{a}(=p_{a^{\prime}})$ and the first best type of the other agents represented by $a^{\prime\prime}$ is $\varnothing$.
+Let $p_{-a}$ be such that there are $q_{o}$ students represented by $a^{\prime}$ whose revealed preference is equivalent to $r_{a}(=p_{a^{\prime}})$ and the first-best type of the other agents represented by $a^{\prime\prime}$ is $\varnothing$.
 Then, $\left(p_{a},p_{-a}\right)$ satisfies (i), (ii) and (iii) with $l=k$, because $k\leq\bar{k}\left(p_{a}\right)$.

### v5

@@
-Strategic Analysis of Fair Rank-Minimizing Mechanisms with Agent Refusal Option
+
 Abstract
@@
 We focus on the case where agents can refuse their assignment and obtain the outside option instead.
-Without the refusal option, truth-telling is not strategically dominated by any strategies if a fair rank-minimizing mechanism is used.
-However, if agents have the option and the uniform rank-minimizing mechanism is used, then a strategy called an outside option demotion strategy strategically dominates truth-telling.
+Without the refusal option, truth-telling is not strategically dominated by any strategies if a fair rank-minimizing mechanism is employed.
+However, if agents have the option and the uniform rank-minimizing mechanism is employed, then a strategy called an outside option demotion strategy strategically dominates truth-telling.
 Moreover, we show that adopting this strategy may lead to inefficient assignments.
-To counter this, we propose a modification of the uniform rank-minimizing mechanism, though it may lead agents to strategically reduce the number of acceptable types.
+To counter this, we consider a modification of the uniform rank-minimizing mechanism, though it may lead agents to strategically reduce the number of acceptable types.
 JEL classification: C78; D47
@@
 There are several agents and types of object, including a null object type, which corresponds to the outside option.
-In the mechanism, each agent reveals one preference ranking, which may or may not be their true preference ranking.
+In a mechanism, each agent reveals one preference ranking, which may or may not be their true preference ranking.
 In rank-minimizing mechanisms, assignments are determined as to minimize the total (or average) revealed rank of the objects to which the agents are assigned.
-The rank distributions of assignments are a critical consideration for many real-world matching authorities.
+The rank distribution of assignments is a critical consideration for many real-world matching authorities.
 For instance, Featherstone (2020) notes that Teach For America takes rank distribution into account when making assignment decisions.
@@
 Moreover, Ortega and Klein (2023) compare the outcomes of rank-minimizing mechanisms with those of the DA mechanism and the TTC mechanism, showing that rank-minimizing mechanisms improve upon the others in several aspects of efficiency.
-In this study, we especially consider the rank-minimizing mechanisms that satisfy a fairness property called equal treatment of equals, which requires that if two agents reveal the same preference ranking, then they are assigned to each object type with the same probability.
+In this study, we focus on rank-minimizing mechanisms that satisfy the fairness property of equal treatment of equals, which ensures that agents revealing identical preference rankings are assigned to each object type with the same probability.
+[^1] Yokote et al.
+(2019) mention that this fairness property is a classical concept discussed by Aristotle in Book V of Nicomachean Ethics.
 This fairness property is discussed by Bogomolnaia and Moulin (2001) and many other previous studies on various allocation problems (Thomson 2011).
@@
 Related to but logically independent to their results, we show that, in this case, truth-telling is not strategically dominated by any untrue preference under any fair rank-minimizing mechanism.
-These results imply that in this case, strategic manipulation may not be a critical problem of rank-minimizing mechanisms.
+These results imply that in this case, strategic manipulation is a critical problem of rank-minimizing mechanisms.
 Next, as the main analysis, we consider the case where agents have the refusal options.
@@
 We consider an outside option demotion strategy, which involves revealing a ranking where the rank of the outside option is placed at the bottom, without changing the order of the truly acceptable object types.
-We show that, if the uniform rank-minimizing mechanism is used and agents have the refusal option, then any outside option demotion strategy weakly strategically dominates truth-telling.
+We show that, if the uniform rank-minimizing mechanism is employed and agents have the refusal option, then any outside option demotion strategy weakly strategically dominates truth-telling.
 Moreover, with some additional conditions, some outside option demotion strategy (strictly) strategically dominates truth-telling.
-The reason is as follows: first, for some preference profiles of other agents, an agent succeeds in avoiding the outside option and getting a truly acceptable object type because assigning the outside option to them significantly increases the sum of ranks.
-Of course, due to this manipulation, the agent faces the risk of being assigned an unacceptable object type.
-However, in the case where the agents have the refusal option, the agent can refuse any of such types and obtain the outside option instead.
-Moreover, we show that if an agent adopts an outside option demotion strategy and the uniform rank-minimizing mechanism is used, an inefficient assignment may be realized.
+The reason is as follows: First, for some preference profiles of other agents, an agent can successfully avoid the outside option and obtain a truly acceptable object type because assigning them the outside option would significantly increase the sum of ranks.
+However, this manipulation carries the risk of being assigned an unacceptable object type.
+Yet, when agents have the refusal option, they can simply reject such assignments and choose the outside option instead.
+Thus, in this case, the risk associated with adopting this strategy is entirely eliminated.
+Moreover, we show that if an agent adopts an outside option demotion strategy and the uniform rank-minimizing mechanism is employed, an inefficient assignment may be realized.
 This is because, in that case, an unacceptable type may be assigned to the agent who adopts the strategy, and then the agent refuses it, even though the type is desired by other agents.
@@
 Each agent must be allocated one copy of an object whose type is an element of a set denoted by $\mathcal{O}$.
-Let $q_{o}\geq 1$ be the capacity of an object type $o\in\mathcal{O}$.
+Let $q_{o}\geq 1$ be the number of copies (the capacity) of an object type $o\in\mathcal{O}$.
 We assume that there is a null object type denoted by $\varnothing\in\mathcal{O}$ satisfying $q_{\varnothing}\geq\left|\mathcal{A}\right|$.
@@
 To avoid trivial cases, we assume $q_{o}\in\left[1,\left|\mathcal{A}\right|\right)$ for all $o\in\mathcal{O\setminus}\left\{\varnothing\right\}$.
-Let $\mathcal{P}$ be the set of all possible strict rankings (linear orders) of the elements in $\mathcal{O}$.
-Let $p_{a}=\left(p_{a}^{1},\cdots,p_{a}^{\left|\mathcal{O}\right|}\right)\in\mathcal{P}^{\left|\mathcal{A}\right|}$ be a preference order revealed by $a$, where $p_{a}^{k}=o$ means that type $o$ object is ranked $k$th-best one under $p_{a}$.
+Let $\mathcal{P}$ be the set of all possible strict rankings of the elements in $\mathcal{O}$.
+Let $p_{a}=\left(p_{a}^{1},\cdots,p_{a}^{\left|\mathcal{O}\right|}\right)\in\mathcal{P}$ be a preference order revealed by $a$, where $p_{a}^{k}=o$ means that type $o$ object is ranked $k$th-best one under $p_{a}$.
 For notational convenience, we let $R:\mathcal{O\times P}\rightarrow\left\{1,\cdots,\left|\mathcal{O}\right|\right\}$ be a rank function; that is, $p_{a}^{k}=o$ if and only if $R\left(o,p_{a}\right)=k$.
@@
 An assignment is represented as lotteries over types.
-Let $x_{ao}$ be the probability that agent $a$ assigns (a copy of) type $o$ object.
+Let $x_{ao}$ represent the probability that agent $a$ assigns (a copy of) type $o$ object.
 Moreover, let $x$ be a matrix whose $\left(a,o\right)$ entry is equal to $x_{ao}$.
@@
 Let $\mathcal{X}$ be the set of all possible assignments.
-Specifically, let $y\in\mathcal{X}$ be a deterministic assignment if $y_{ao}$ is either 00 or $1$ for all $\left(a,o\right)\in\mathcal{A}\times\mathcal{O}$.
+Specifically, $y\in\mathcal{X}$ is said to be a deterministic assignment if $y_{ao}$ is either 00 or $1$ for all $\left(a,o\right)\in\mathcal{A}\times\mathcal{O}$.
 Let $\mathcal{Y}\subseteq\mathcal{X}$ be a set of all possible deterministic assignments.
@@
 Every assignment can be written as a convex combination of deterministic assignments.
-We consider the preference of agent $a$ with $p_{a}$ with regard to two assignments $x$ and $x^{\prime}$. If
+We consider the preference of agent $a$ denoted by $p_{a}$ with regard to two assignments $x$ and $x^{\prime}$. If
 $\sum\limits_{o^{\prime}:\text{ }R\left(o^{\prime},p_{a}\right)\leq k}x_{ao^{\prime}}\geq\sum\limits_{o^{\prime}:\text{ }R\left(o^{\prime},p_{a}\right)\leq k}x_{ao^{\prime}}^{\prime}$ (1)
@@
 Assignment $x$ is wasteful for $p$ if there are $o,o^{\prime}\in\mathcal{O}$ and $a\in\mathcal{A}$ such that $\sum\nolimits_{a^{\prime}\in\mathcal{A}}x_{a^{\prime}o}<q_{o}$, $x_{ao^{\prime}}>0$ and $R\left(o,p_{a}\right)<R\left(o^{\prime},p_{a}\right)$.
+In words, an assignment is considered wasteful for a given preference profile if there is at least one agent who, despite the fact that a more desirable object is still available, does not receive as much of it as they could, and instead receives some amount of a less preferred object.
 For an assignment $x\in\mathcal{X}$ and a preference profile $p=\left(p_{a}\right)_{a\in\mathcal{A}}\in\mathcal{P}^{\left|\mathcal{A}\right|}$, let $RV\left(x,p\right)$ be such that
@@
 If $x^{\ast}\in\mathcal{X}$ is rank-minimizing for $p$, then $x$ is not wasteful for $p$.
-The efficiency property of ”rank-minimizing” is stronger than several efficiency properties, such as rank efficiency and ordinal efficiency, analyzed in this literature (see Figure 1 in Feizi (2024)).
+The efficiency property of rank-minimizing is stronger than several other efficiency properties, such as rank efficiency and ordinal efficiency, analyzed in this literature (see Figure 1 in Feizi (2024)).
 Among them, not being wasteful is a very weak property.
@@
 There must exist at least one deterministic rank-minimizing assignment.
-Nikzad (2022) and Troyan (2024) consider a deterministic rank-minimizing assignment; that is, they focus on a deterministic assignment $y^{\ast}\in\mathcal{Y}$ such that $RV\left(y^{\ast},p\right)\leq RV\left(y,p\right)$ for all $y\in\mathcal{Y}$.
+Nikzad (2022) and Troyan (2024) focus only on deterministic rank-minimizing assignments; that is, they consider a deterministic assignment $y^{\ast}\in\mathcal{Y}$ such that $RV\left(y^{\ast},p\right)\leq RV\left(y,p\right)$ for all $y\in\mathcal{Y}$.
 However, by Corollary 1, such $y^{\ast}$ must also be rank-minimizing in our definition; that is, $RV\left(y^{\ast},p\right)\leq RV\left(x,p\right)$ for all $x\in\mathcal{X}$.
@@
 which is well-defined because $\sum\nolimits_{\bar{o}\in\mathcal{O}}q_{\bar{o}}>\left|\mathcal{A}\right|$.
-This means that for a given $p_{a}$, there are sufficient number of copies of object types that ranked higher than or equal to $\bar{k}\left(p_{a}\right)$.
+This means that for a given $p_{a}\in\mathcal{P}$, there are sufficient number of copies of object types that ranked higher than or equal to $\bar{k}\left(p_{a}\right)$.
 Trivially, if $p_{a}^{k}=\varnothing$, then $\bar{k}\left(p_{a}\right)\leq k$.
@@
 Lemma 2
-Let $x\in\mathcal{X}$ be an assignment that is not wasteful for $p$. Then, $x_{ao}=0$ for all $o$ such that $R(o,p_{a})>\bar{k}\left(p_{a}\right)$.
-Then, if an assignment is not wasteful, then $a$ must be assigned to an object type that ranked higher than or equals to $\bar{k}\left(p_{a}\right)$.
-Typically, in that case, the object types that ranked lower than $\varnothing$ in the preference ranking of agent $a$ are never assigned to $a$, because $\varnothing$ is the object with a sufficient number of copies.
+If $x\in\mathcal{X}$ be an assignment that is not wasteful for $p$, then $x_{ao}=0$ for all $o$ such that $R(o,p_{a})>\bar{k}\left(p_{a}\right)$.
+This means that if an assignment is not wasteful for $p$, then $a$ must be assigned to an object type that ranked higher than or equals to $\bar{k}\left(p_{a}\right)$.
+Typically, under a nonwasteful assignment, the object types that ranked lower than $\varnothing$ in the preference ranking of agent $a$ are never assigned to $a$, because $\varnothing$ is the object with a sufficient number of copies.
 3 Rank-minimizing Mechanism
@@
 Second, based on $p$, an assignment is determined via a mechanism.
-We assume that all agents know the capacities of all object types $\left(q_{o}\right)_{o\in\mathcal{O}}$ and the mechanism used to determine the assignment, when they reveal a preference order.
+We assume that all agents know the capacities of all object types $\left(q_{o}\right)_{o\in\mathcal{O}}$ and the mechanism is employed to determine the assignment, when they reveal a preference order.
 Note that in this section, the null object type is just that with sufficient number of copies.
-A mechanism used in the second stage is represented by $f:\mathcal{P}^{\left|\mathcal{A}\right|}\rightarrow\mathcal{X}$.
+A mechanism employed in the second stage is represented by $f:\mathcal{P}^{\left|\mathcal{A}\right|}\rightarrow\mathcal{X}$.
 First, we consider an efficiency property on mechanisms.
@@
 Unlike in our model, they take into account the priorities of objects and define agents as equal only if they are equal with respect to the priorities of all objects.
+By Lemma 2, as long as the assignment is not wasteful, any type that is ranked lower than $\bar{k}\left(p_{a}\right)$th in the preference ranking revealed by $a$ will not be assigned to $a$, because there are sufficient copies of types ranked higher than or equal to $\bar{k}\left(p_{a}\right)$th.
+Thus, if two individuals reveal rankings that are identical for positions at or above $\bar{k}\left(p_{a}\right)$th rank, those rankings are considered essentially the same, and they should receive the same assignment in an ex-ante sense.
 Other fairness properties discussed in the literature include weak envy-freeness and the equal division lower bound.
@@
 Since ETE is weaker than these fairness properties, it is compatible with rank-minimizing and is therefore adopted in this study.
-By Lemma 2, as long as the assignment is not wasteful, any type that is ranked lower than $\bar{k}\left(p_{a}\right)$th in the preference ranking revealed by $a$ will not be assigned to $a$, because there are sufficient copies of types ranked higher than or equal to $\bar{k}\left(p_{a}\right)$th.
-Thus, if two individuals reveal rankings that are identical for positions at or above $\bar{k}\left(p_{a}\right)$th rank, those rankings are considered essentially the same, and they should receive the same assignment in an ex-ante sense.
 We focus on the efficient and fair mechanisms; that is, we call $f$ a fair rank-minimizing mechanism if $f(p)$ is rank-minimizing for all $p$ and satisfies ETE. The following mechanism is one of them.
@@
 We call the strategy of revealing their true preference truth-telling.
-Featherstone (2020) shows that any rank-minimizing mechanism does not satisfy strategy-proofness.
+Featherstone (2020, Proposition 10) shows that any rank-minimizing mechanism does not satisfy strategy-proofness.
 Therefore, in this paper, we consider a weaker strategic property.
-We discuss whether truth-telling ($r_{a}$) can be strategically dominated by an untrue preference $p_{a}\neq r_{a}$.
+For a mechanism $f$ and an agent $a$, (revealing) $p_{a}^{\prime}$ is said to be weakly strategically dominated by (revealing) $p_{a}$ if $a$ weakly prefers $f\left(p_{a},p_{-a}\right)$ to $f\left(p_{a}^{\prime},p_{-a}\right)$ for all $p_{-a}\in\mathcal{P}^{\left|\mathcal{A}\right|-1}$.
+Moreover, for $f$ and $a$, (revealing) $p_{a}^{\prime}$ is said to be (strictly) strategically dominated by (revealing) $p_{a}$ if $p_{a}^{\prime}$ is weakly strategically dominated by $p_{a}$ and there exists some $p_{-a}^{\prime}\in\mathcal{P}^{\left|\mathcal{A}\right|-1}$ such that $a$ strictly prefers $f\left(p_{a},p_{-a}^{\prime}\right)$ to $f\left(p_{a}^{\prime},p_{-a}^{\prime}\right)$.
+We discuss whether truth-telling ($r_{a}$) can be strategically dominated by revealing an untrue preference $p_{a}\neq r_{a}$.
+Fix any $r_{a}$ and any $p_{a}\neq r_{a}$.
 Let $k$ be the smallest integer such that $p_{ak}\neq r_{ak}$.
 By Lemma 2, if $k>\bar{k}\left(p_{a}\right)$, then $p_{a}$ and $r_{a}$ are essentially equivalent under a fair rank-minimizing mechanism, because any result of a rank-minimizing mechanism is not wasteful.
-Thus, we focus only on the case where $k\leq\bar{k}\left(p_{a}\right)$ defined above; that is,
+Thus, we focus only on $r_{a}$ and $p_{a}$ satisfying $k\leq\bar{k}\left(p_{a}\right)$ defined above; that is,
 $\sum\limits_{\bar{o}:R(\bar{o},p_{a})<k}q_{\bar{o}}\left(=\sum\limits_{\bar{o}:R(\bar{o},r_{a})<k}q_{\bar{o}}\right)<\left|\mathcal{A}\right|$
@@
 If $k\leq\bar{k}\left(p_{a}\right)$, then there is $p_{-a}\in\mathcal{P}^{\left|\mathcal{A}\right|-1}$ such that agent $a$ (with $r_{a}$) strictly prefers $f\left(r_{a},p_{-a}\right)$ to $f\left(p_{a},p_{-a}\right)$.
-By this lemma, we have the following result.
+By Lemma 3, we have the following result.
 Proposition 2
-Suppose that a fair rank-minimizing mechanism is used. For any $r_{a}\in\mathcal{P}$, there is no $p_{a}\in\mathcal{P}$ such that revealing $p_{a}$ strategically dominates truth-telling.
+Suppose that a fair rank-minimizing mechanism is employed. For any $r_{a}\in\mathcal{P}$, there is no $p_{a}\in\mathcal{P}$ such that revealing $p_{a}$ strategically dominates truth-telling.
 This result implies that any fair rank-minimizing mechanism satisfies a weak strategic characteristic; that is, for each untrue preference $p_{a}$, either $\left(f\left(r_{a},p_{-a}\right)\right)_{ao}=\left(f\left(p_{a},p_{-a}\right)\right)_{ao}$ for all $o\in\mathcal{O}$ and all $p_{-a}\in\mathcal{P}^{\left|\mathcal{A}\right|-1}$ or there must exist $p_{-a}$ such that $a$ is worse off by revealing it instead of $r_{a}$.
-Thus, if a fair rank-minimizing mechanism is used, any (essential) strategic manipulation by an agent carries some level of risk.
+Thus, if a fair rank-minimizing mechanism is employed, any (essential) strategic manipulation by an agent carries some level of risk.
 This characteristic is quite weak, but in the subsequent section, we show that this is not satisfied when the agents can refuse their assignment and instead obtain the null object type.
-Note that Proposition 2 holds for any fair rank-minimizing mechanisms but cannot be generalized to the rank-minimizing mechanisms satisfying weak ETE.
-We show this by introducing the following example.
+Note that Proposition 2 holds for any fair rank-minimizing mechanisms but cannot be generalized to rank-minimizing mechanisms satisfying weak ETE. We show this by introducing the following example.
 Example 1
 Let $\mathcal{A=}\left\{a_{1},a_{2},a_{3}\right\}$, $\mathcal{O=}\left\{o_{1},o_{2},o_{3},\varnothing\right\}$ and $q_{o_{1}}=1$.
-Let $\psi=\left(o_{1},o_{2},\varnothing,o_{3}\right)$ and $\psi^{\prime}=\left(o_{1},o_{2},o_{3},\varnothing\right)$.
-A rank-minimizing mechanism $f$ is such that
-$\left(f\left(\psi^{\prime},\psi,\psi\right)\right)_{a_{1}o_{1}}$ $=$ $\left(f\left(\psi,\psi^{\prime},\psi\right)\right)_{a_{2}o_{1}}=\left(f\left(\psi,\psi,\psi^{\prime}\right)\right)_{a_{3}o_{1}}=0,$ (2)
-$f\left(p_{a_{1}},p_{a_{2}},p_{a_{3}}\right)$ $=$ $\sum\limits_{y^{\ast}\in\mathcal{Y}^{\ast}\left(p\right)}\frac{y^{\ast}}{\left|\mathcal{Y}^{\ast}\left(p\right)\right|}$
-if $\left(p_{a_{1}},p_{a_{2}},p_{a_{3}}\right)$ is neither $\left(\psi^{\prime},\psi,\psi\right)$ nor $\left(\psi,\psi^{\prime},\psi\right)$ nor $\left(\psi,\psi,\psi^{\prime}\right)$.
-That is, if the first-best type and the second-best type of the three agents are respectively $o_{1}$ and $o_{2}$ and the third-best type of the two of them is $\varnothing$, then the agent revealing $o_{3}$ as the third-best one obtains $o_{1}$ with probability 00.
-Trivially, $f$ satisfies weak ETE.
-However, whether $f$ satisfies ETE or not is dependent on $q_{o_{2}}$.
-We assume that $r_{a_{1}}=\psi^{\prime}$ and this rank-minimizing mechanism is used.
+Let $\alpha=\left(o_{1},o_{2},\varnothing,o_{3}\right)$ and $\alpha^{\prime}=\left(o_{1},o_{2},o_{3},\varnothing\right)$.
+Let a rank-minimizing mechanism $f^{\ast}$ be such that, if $\left(p_{a_{i}},p_{a_{j}},p_{a_{k}}\right)=\left(\alpha^{\prime},\alpha,\alpha\right)$ where $i,j,k=1,2,3$ and they are distinct, then
+$\left(f^{\ast}\left(p_{a_{1}},p_{a_{2}},p_{a_{3}}\right)\right)_{a_{i}o_{1}}$ $=$ $0,$ (2)
+$f^{\ast}\left(p_{a_{1}},p_{a_{2}},p_{a_{3}}\right)$ $=$ $\sum\limits_{y^{\ast}\in\mathcal{Y}^{\ast}\left(p\right)}\frac{y^{\ast}}{\left|\mathcal{Y}^{\ast}\left(p\right)\right|}$
+otherwise; that is, if $\left(p_{a_{1}},p_{a_{2}},p_{a_{3}}\right)$ is neither $\left(\alpha^{\prime},\alpha,\alpha\right)$ nor $\left(\alpha,\alpha^{\prime},\alpha\right)$ nor $\left(\alpha,\alpha,\alpha^{\prime}\right)$.
+In words, if the first-best type and the second-best type of the three agents are respectively $o_{1}$ and $o_{2}$ and the third-best type of the two of them is $\varnothing$, then the agent revealing $o_{3}$ as the third-best one obtains $o_{1}$ with probability 00.
+We additionally assume that if $\left(p_{a_{i}},p_{a_{j}},p_{a_{k}}\right)=\left(\alpha^{\prime},\alpha,\alpha\right)$, then
+$\left(f^{\ast}\left(p_{a_{1}},p_{a_{2}},p_{a_{3}}\right)\right)_{a_{j}\bar{o}}=\left(f^{\ast}\left(p_{a_{1}},p_{a_{2}},p_{a_{3}}\right)\right)_{a_{k}\bar{o}}$
+for all $\bar{o}\in\mathcal{O}$. Then, $f^{\ast}$ satisfies weak ETE. However, whether $f^{\ast}$ satisfies ETE or not is dependent on $q_{o_{2}}$.
+We assume that $r_{a_{1}}=\alpha^{\prime}$ and this rank-minimizing mechanism $f^{\ast}$ is employed.
 First, suppose $q_{o_{2}}=1$.
-In this case, $\bar{k}\left(\psi\right)=\bar{k}\left(\psi^{\prime}\right)=3$.
-Then, $r_{a_{1}}=\psi^{\prime}$ is not strategically dominated by $\psi$ and any other preference orders, because
-$\left(f\left(\psi^{\prime},\psi^{\prime},\psi^{\prime}\right)\right)_{a_{1}o_{1}}$ $=$ $\left(f\left(\psi^{\prime},\psi^{\prime},\psi^{\prime}\right)\right)_{a_{1}o_{2}}=\left(f\left(\psi^{\prime},\psi^{\prime},\psi^{\prime}\right)\right)_{a_{1}o_{3}}=\frac{1}{3},$
-$\left(f\left(\psi,\psi^{\prime},\psi^{\prime}\right)\right)_{a_{1}o_{1}}$ $=$ $\left(f\left(\psi,\psi^{\prime},\psi^{\prime}\right)\right)_{a_{1}o_{2}}=\left(f\left(\psi,\psi^{\prime},\psi^{\prime}\right)\right)_{a_{1}\varnothing}=\frac{1}{3}.$
-That is, if $p_{a_{2}}=p_{a_{3}}=\psi^{\prime}$, then $a_{1}$ is worse off by revealing $\psi$ instead of $\psi^{\prime}$.
-Moreover, in this case, $f$ satisfies ETE and thus this is a fair rank-minimizing mechanism.
+In this case, $f^{\ast}$ satisfies ETE and thus this is a fair rank-minimizing mechanism.
+In this case, $\bar{k}\left(\alpha\right)=\bar{k}\left(\alpha^{\prime}\right)=3$.
+Then, for example, $r_{a_{1}}=\alpha^{\prime}$ is not strategically dominated by $\alpha$ and any other preference orders, because
+$\left(f^{\ast}\left(\alpha^{\prime},\alpha^{\prime},\alpha^{\prime}\right)\right)_{a_{1}o_{1}}$ $=$ $\left(f^{\ast}\left(\alpha^{\prime},\alpha^{\prime},\alpha^{\prime}\right)\right)_{a_{1}o_{2}}=\left(f^{\ast}\left(\alpha^{\prime},\alpha^{\prime},\alpha^{\prime}\right)\right)_{a_{1}o_{3}}=\frac{1}{3},$
+$\left(f^{\ast}\left(\alpha,\alpha^{\prime},\alpha^{\prime}\right)\right)_{a_{1}o_{1}}$ $=$ $\left(f^{\ast}\left(\alpha,\alpha^{\prime},\alpha^{\prime}\right)\right)_{a_{1}o_{2}}=\left(f^{\ast}\left(\alpha,\alpha^{\prime},\alpha^{\prime}\right)\right)_{a_{1}\varnothing}=\frac{1}{3}.$
+That is, if $p_{a_{2}}=p_{a_{3}}=\alpha^{\prime}$, then $a_{1}$ is worse off by revealing $\alpha$ instead of $\alpha^{\prime}=r_{a_{1}}\,$(truth-telling).
 Second, suppose $q_{o_{2}}\geq 2$.
-In this case, $\bar{k}\left(\psi\right)=\bar{k}\left(\psi^{\prime}\right)=2$.
-Then, $a_{1}$ must have either $o_{1}$ or $o_{2}$ if it chooses either $\psi$ or $\psi^{\prime}$ and thus, the unique interest of $a_{1}$ is increasing the probability to obtain $o_{1}$.
-Thus, in this case, $r_{a_{1}}=\psi^{\prime}$ is strategically dominated by $\psi$.
-Therefore, even if a rank-minimizing mechanism is used, truth-telling may be strategically dominated by some other strategy.
-Moreover, in this case, $f$ does not satisfy ETE.
+In this case, $\bar{k}\left(\alpha\right)=\bar{k}\left(\alpha^{\prime}\right)=2$.
+Moreover, if $p_{a_{i}}$ is either $\alpha$ or $\alpha^{\prime}$, then a rank-minimizing mechanism satisfies
+$\left(f^{\ast}\left(p_{a_{1}},p_{a_{2}},p_{a_{3}}\right)\right)_{a_{i}o_{1}}+\left(f^{\ast}\left(p_{a_{1}},p_{a_{2}},p_{a_{3}}\right)\right)_{a_{i}o_{2}}=1.$
+That is, if a rank-minimizing mechanism is employed and $a_{1}$ chooses either $\alpha$ or $\alpha^{\prime}$, then $a_{1}$ must have either $o_{1}$ or $o_{2}$.
+Thus, the only strategic interest of $a_{1}$ is increasing the probability to obtain $o_{1}$.
+If $f^{\ast}$ is employed, then $\alpha^{\prime}=r_{a_{1}}$ is strategically dominated by $\alpha$ for $a_{1}$ due to (2).
+Therefore, we conclude that even if a rank-minimizing mechanism that satisfies weak ETE is employed, truth-telling may be strategically dominated by some other strategy.
+Note that in the same case, $f^{\ast}$ does not satisfy ETE.
 First, since $q_{o_{1}}+q_{o_{2}}\geq\left|\mathcal{A}\right|$, any agent is assigned to neither $o_{3}$ nor $\varnothing$,∅ , as long as the assignment is not wasteful.
-Thus, $\psi$ and $\psi^{\prime}$ are essentially the same in this case.
-However, $\left(f\left(\psi^{\prime},\psi,\psi\right)\right)_{a_{1}o_{1}}=0,$ but $\left(f\left(\psi^{\prime},\psi,\psi\right)\right)_{a_{1}o_{2}}>0$ or $\left(f\left(\psi^{\prime},\psi,\psi\right)\right)_{a_{1}o_{3}}>0$ must be satisfied.
-Therefore, in this case, $f$ does not satisfy ETE.
+Thus, $\alpha$ and $\alpha^{\prime}$ are essentially the same in this case.
+However, $\left(f^{\ast}\left(\alpha^{\prime},\alpha,\alpha\right)\right)_{a_{1}o_{1}}=0,$ but $\left(f^{\ast}\left(\alpha^{\prime},\alpha,\alpha\right)\right)_{a_{1}o_{2}}>0$ or $\left(f^{\ast}\left(\alpha^{\prime},\alpha,\alpha\right)\right)_{a_{1}o_{3}}>0$ must be satisfied.
+Hence $f^{\ast}$ does not satisfy ETE.
 4 Refusal option
@@
 That is, in this game, the null object type has two characteristics: first, there are sufficiently many copies of this type to allocate to all individuals, and second, agents can always obtain it at no cost as long as they abandon their assignment.
-This third stage is also considered by Featherstone (2020) but is not explicitly addressed by almost previous studies related to ours.
+This third stage is also considered by Featherstone (2020) but is not explicitly addressed by almost other previous studies related to ours.
 [^4] As exceptions, Feigenbaum et al.
@@
 That is, since an agent refuses any of the unacceptable object types when it is assigned, the probability that the null object type is assigned will increase by the probability that any of the unacceptable object types is assigned.
-We reconsider strategic problems of a fair rank-minimizing mechanism.
-For a mechanism $f$, (revealing) $p_{a}^{\prime}$ (for $a$) is said to be strategically dominated by (revealing) $p_{a}$ with the refusal option if $a$ weakly prefers $g\left(f\left(p_{a},p_{-a}\right),r\right)$ to $g\left(f\left(p_{a}^{\prime},p_{-a}\right),r\right)$ for all $p_{-a}\in\mathcal{P}^{\left|\mathcal{A}\right|-1}$ and strictly prefers $g\left(f\left(p_{a},p_{-a}^{\prime}\right),r\right)$ to $g\left(f\left(p_{a}^{\prime},p_{-a}^{\prime}\right),r\right)$ for some $p_{-a}^{\prime}\in\mathcal{P}^{\left|\mathcal{A}\right|-1}$.
+We reconsider strategic problems of a fair rank-minimizing mechanism in this three-stage game.
+For a mechanism $f$ and an agent $a$, (revealing) $p_{a}^{\prime}$ is said to be weakly strategically dominated by (revealing) $p_{a}$ with the refusal option if $a$ weakly prefers $g\left(f\left(p_{a},p_{-a}\right),r\right)$ to $g\left(f\left(p_{a}^{\prime},p_{-a}\right),r\right)$ for all $p_{-a}\in\mathcal{P}^{\left|\mathcal{A}\right|-1}$.
+Moreover, for $f$ and $a$, (revealing) $p_{a}^{\prime}$ is said to be (strictly) strategically dominated by (revealing) $p_{a}$ with the refusal option if $p_{a}^{\prime}$ is weakly strategically dominated by $p_{a}$ with the refusal option and there exists some $p_{-a}^{\prime}\in\mathcal{P}^{\left|\mathcal{A}\right|-1}$ such that $a$ strictly prefers $g\left(f\left(p_{a},p_{-a}^{\prime}\right),r\right)$ to $g\left(f\left(p_{a}^{\prime},p_{-a}^{\prime}\right),r\right)$.
 We show that, under the uniform rank-minimizing mechanism, truth-telling is strategically dominated by another strategy when the agents have the refusal option.
-We let $\mathcal{D}_{a}\subseteq\mathcal{P}$ be a set of preferences such that $d_{a}\in\mathcal{D}_{a}$ if (i) $R\left(o,r_{a}\right)=R\left(o,d_{a}\right)$ for all $o\in\mathcal{O}$ such that $R\left(o,r_{a}\right)<R\left(\varnothing,r_{a}\right)$ and (ii) $R\left(\varnothing,d_{a}\right)=\left|\mathcal{O}\right|$.
+We define the following specific strategies.
+Let $\mathcal{D}_{a}\subseteq\mathcal{P}$ be a set of preferences such that $d_{a}\in\mathcal{D}_{a}$ if (i) $R\left(o,r_{a}\right)=R\left(o,d_{a}\right)$ for all $o\in\mathcal{O}$ such that $R\left(o,r_{a}\right)<R\left(\varnothing,r_{a}\right)$ and (ii) $R\left(\varnothing,d_{a}\right)=\left|\mathcal{O}\right|$.
 First, the ranks of the truly acceptable object types are unchanged.
 Second, the null object type $\varnothing$ is the least preferable one under $d_{a}\in\mathcal{D}_{a}$.
-Thus, any object type that is less preferable than $\varnothing$ in the true preference has changed to be more preferable than $\varnothing$ in $d_{a}$.
 We call the strategy of agent $a$ of revealing a preference in $\mathcal{D}_{a}$ an outside option demotion strategy (hereafter ODS) of agent $a$.
+Under an ODS, every object type that is ranked below $\varnothing$ in the agent’s true preference is instead ranked above $\varnothing$ in $d_{a}$.
 Note that an ODS (revealing $d_{a}\in\mathcal{D}_{a}$) increases the number of acceptable object types revealed by $a$; that is, $R\left(\varnothing,r_{a}\right)\leq R\left(\varnothing,d_{a}\right)=\left|\mathcal{O}\right|$ is satisfied.
@@
 Let $\mathcal{A=}\left\{a_{1},a_{2},a_{3}\right\}$ and $\mathcal{O=}\left\{o_{1},o_{2},\varnothing\right\}$.
-Suppose $q_{o_{1}}=q_{o_{2}}=1$ and $r_{a_{1}}=(o_{1},\varnothing,o_{2})$.
-Then, the unique ODS is revealing $d_{a_{1}}=(o_{1},o_{2},\varnothing)$.
-First, suppose
-$p_{a_{2}}=(o_{1},o_{2},\varnothing)\text{ and }p_{a_{3}}=(o_{2},o_{1},\varnothing).$
-Then,
-$f^{U}\left(r_{a_{1}},p_{a_{2}},p_{a_{3}}\right)=y=\left(\begin{array}[]{ccc}y_{a_{1}o_{1}}&y_{a_{1}o_{2}}&y_{a_{1}\varnothing}\\ y_{a_{2}o_{1}}&y_{a_{2}o_{2}}&y_{a_{2}\varnothing}\\ y_{a_{3}o_{1}}&y_{a_{3}o_{2}}&y_{a_{3}\varnothing}\end{array}\right)=\left(\begin{array}[]{ccc}0&0&1\\ 1&0&0\\ 0&1&0\end{array}\right).$
+Let $\beta=\left(o_{1},o_{2},\varnothing\right)$, $\beta^{\prime}=\left(o_{2},o_{1},\varnothing\right)$ and $\beta^{\prime\prime}=(o_{1},\varnothing,o_{2})$.
+Suppose $q_{o_{1}}=q_{o_{2}}=1$ and $r_{a_{1}}=\beta^{\prime\prime}$.
+Then, the unique ODS is revealing $d_{a_{1}}=\beta$.
+In this example, we assume $\left(p_{a_{2}},p_{a_{3}}\right)=\left(r_{a_{2}},r_{a_{3}}\right);$ that is, $a_{2}$ and $a_{3}$ always adopt truth-telling.
+First, we consider the case where $p_{a_{2}}=\beta$ and $p_{a_{3}}=\beta^{\prime}$. Then,
+$f^{U}\left(r_{a_{1}},\beta,\beta^{\prime}\right)=y=\left(\begin{array}[]{ccc}y_{a_{1}o_{1}}&y_{a_{1}o_{2}}&y_{a_{1}\varnothing}\\ y_{a_{2}o_{1}}&y_{a_{2}o_{2}}&y_{a_{2}\varnothing}\\ y_{a_{3}o_{1}}&y_{a_{3}o_{2}}&y_{a_{3}\varnothing}\end{array}\right)=\left(\begin{array}[]{ccc}0&0&1\\ 1&0&0\\ 0&1&0\end{array}\right).$
 Whenever $\left|\mathcal{A}\right|=\left|\mathcal{O}\right|=3$, we write the matrix representing an assignment in this manner. On the other hand,
-$f^{U}\left(d_{a_{1}},p_{a_{2}},p_{a_{3}}\right)=\left(\begin{array}[]{ccc}1/2&0&1/2\\ 1/2&0&1/2\\ 0&1&0\end{array}\right)$
-Thus, revealing $d_{a_{1}}$ instead of $r_{a_{1}}$ may increase the probability that $a_{1}$ is assigned to their first-best object $o_{1}$. On the other hand, suppose
-$p_{a_{2}}^{\prime}=(o_{1},\varnothing,o_{2})\text{ and }p_{a_{3}}^{\prime}=(o_{1},\varnothing,o_{2}).$
-Then,
-$f^{U}\left(r_{a_{1}},p_{a_{2}}^{\prime},p_{a_{3}}^{\prime}\right)=\left(\begin{array}[]{ccc}1/3&0&2/3\\ 1/3&0&2/3\\ 1/3&0&2/3\end{array}\right),$
-$f^{U}\left(d_{a_{1}},p_{a_{2}}^{\prime},p_{a_{3}}^{\prime}\right)=\left(\begin{array}[]{ccc}1/3&2/3&0\\ 1/3&0&2/3\\ 1/3&0&2/3\end{array}\right).$
+$f^{U}\left(d_{a_{1}},\beta,\beta^{\prime}\right)=\left(\begin{array}[]{ccc}1/2&0&1/2\\ 1/2&0&1/2\\ 0&1&0\end{array}\right)$
+Thus, revealing $d_{a_{1}}$ instead of $r_{a_{1}}$ may increase the probability that $a_{1}$ is assigned to their first-best object $o_{1}$.
+First, under $\left(r_{a_{1}},\beta,\beta^{\prime}\right),$varnothing is ranked lowest by $a_{2}$ and $a_{3}$ but $a_{1}$ ranks $\varnothing$ as their second-best.
+Thus, in any rank-minimizing mechanism, $a_{1}$ is assigned $\varnothing$ with probability 1.
+On the other hand, since $a_{1}$ places $\varnothing$ at the bottom of their ranking.
+As a result.
+under $\left(d_{a_{1}},p_{a_{2}},p_{a_{3}}\right)$, $a_{1}$ is assigned $o_{1}$—the object that is truly their best—with probability $1/2$, and $\varnothing$ with probability $1/2$.
+Thus, in this case, $a_{1}$ is better off by choosing $d_{a_{1}}$ instead of $r_{a_{1}}$.
+Second, suppose $p_{a_{2}}=p_{a_{3}}=\beta^{\prime\prime}.$ Then,
+$f^{U}\left(r_{a_{1}},\beta^{\prime\prime},\beta^{\prime\prime}\right)=\left(\begin{array}[]{ccc}1/3&0&2/3\\ 1/3&0&2/3\\ 1/3&0&2/3\end{array}\right),$
+$f^{U}\left(d_{a_{1}},\beta^{\prime\prime},\beta^{\prime\prime}\right)=\left(\begin{array}[]{ccc}1/3&2/3&0\\ 1/3&0&2/3\\ 1/3&0&2/3\end{array}\right).$
 Thus, in this case, $a_{1}$ seems to be worse off by revealing $d_{a_{1}}$ instead of $r_{a_{1}}$, because $a_{1}$ prefers $\varnothing$ to $o_{2}$.
-However, in the model introduced in this section, $a_{1}$ can refuse $o_{2}$ and in that case, $\varnothing$ is assigned.
-Therefore,
-$g\left(f^{U}\left(d_{a_{1}},p_{a_{2}}^{\prime},p_{a_{3}}^{\prime}\right),r\right)=g\left(f^{U}\left(r_{a_{1}},p_{a_{2}}^{\prime},p_{a_{3}}^{\prime}\right),r\right)\text{;}$
+However, $a_{1}$ can also refuse $o_{2}$ and obtain $\varnothing$.
+Therefore, if $r_{a_{2}}=r_{a_{3}}=\beta^{\prime\prime}$,
+$g\left(f^{U}\left(d_{a_{1}},\beta^{\prime\prime},\beta^{\prime\prime}\right),r\right)=g\left(f^{U}\left(r_{a_{1}},\beta^{\prime\prime},\beta^{\prime\prime}\right),r\right)=f^{U}\left(r_{a_{1}},\beta^{\prime\prime},\beta^{\prime\prime}\right)\text{;}$
 that is, $a_{1}$ is not worse off by choosing $d_{a_{1}}$ instead of $r_{a_{1}}$.
-Next, let $\hat{p}_{a_{2}}=\hat{p}_{a_{3}}=(o_{1},o_{2},\varnothing)$. Then,
-$f^{U}\left(r_{a_{1}},\hat{p}_{a_{2}},\hat{p}_{a_{3}}\right)=\left(\begin{array}[]{ccc}0&0&1\\ 1/2&1/2&0\\ 1/2&1/2&0\end{array}\right),$
-$f^{U}\left(d_{a_{1}},\hat{p}_{a_{2}},\hat{p}_{a_{3}}\right)=\left(\begin{array}[]{ccc}1/3&1/3&1/3\\ 1/3&1/3&1/3\\ 1/3&1/3&1/3\end{array}\right).$
-In the latter case, $a_{1}$ may refuse the assignment; that is,
-$g\left(f^{U}\left(d_{a_{1}},\hat{p}_{a_{2}},\hat{p}_{a_{3}}\right),r\right)=\left(\begin{array}[]{ccc}1/3&0&2/3\\ 1/3&1/3&1/3\\ 1/3&1/3&1/3\end{array}\right),$
-which is wasteful, because
-$\sum\limits_{a\in\left\{a_{1},a_{2},a_{3}\right\}}g\left(f^{U}\left(d_{a_{1}},\hat{p}_{a_{2}},\hat{p}_{a_{3}}\right),r\right)_{ao_{2}}$ $=$ $\frac{2}{3}<1,$
-and $\text{and }g\left(f^{U}\left(d_{a_{1}},\hat{p}_{a_{2}},\hat{p}_{a_{3}}\right),r\right)_{a^{\prime}\varnothing}$ $=$ $\frac{1}{3}>0,$
-for $a^{\prime}=a_{2},a_{3}$. Thus, if
+Thus, in this example, $a_{1}$ is never worse off and may even be better off by selecting $d_{a_{1}}$ (the ODS) instead of $r_{a_{1}}$.
+Next, we show that the ODS by $a_{1}$ may make the assignment wasteful.
+Third, suppose $p_{a_{2}}=p_{a_{3}}=\beta$. Then,
+$f^{U}\left(r_{a_{1}},\beta,\beta\right)=\left(\begin{array}[]{ccc}0&0&1\\ 1/2&1/2&0\\ 1/2&1/2&0\end{array}\right),$
+$f^{U}\left(d_{a_{1}},\beta,\beta\right)=\left(\begin{array}[]{ccc}1/3&1/3&1/3\\ 1/3&1/3&1/3\\ 1/3&1/3&1/3\end{array}\right).$
+In the latter case, $a_{1}$ refuses the assignment; that is,
+$g\left(f^{U}\left(d_{a_{1}},\beta,\beta\right),r\right)=\left(\begin{array}[]{ccc}1/3&0&2/3\\ 1/3&1/3&1/3\\ 1/3&1/3&1/3\end{array}\right).$
+Then, $a_{1}$ is better off by adopting the ODS instead of truth-telling. However, $g\left(f^{U}\left(d_{a_{1}},\beta,\beta\right),r\right)$ is wasteful, because
+$\sum\limits_{a\in\left\{a_{1},a_{2},a_{3}\right\}}g\left(f^{U}\left(d_{a_{1}},\beta,\beta\right),r\right)_{ao_{2}}$ $=$ $\frac{2}{3}<1,$
+and $\text{and }g\left(f^{U}\left(d_{a_{1}},\beta,\beta\right),r\right)_{a^{\prime}\varnothing}$ $=$ $\frac{1}{3}>0,$
+for $a^{\prime}=a_{2},a_{3}$. That is, for
 $\tilde{x}=\left(\begin{array}[]{ccc}1/3&0&2/3\\ 1/3&1/2&1/6\\ 1/3&1/2&1/6\end{array}\right),$
-then $a_{2}$ and $a_{3}$ strictly prefer and $a_{1}$ weakly prefers $\tilde{x}$ to $g\left(f^{U}\left(d_{a_{1}},\hat{p}_{a_{2}},\hat{p}_{a_{3}}\right),r\right)$.
-In the case where $a_{1}$ is assigned to $o_{2}$, either $a_{2}$ or $a_{3}$ is assigned to $\varnothing$ but $a_{1}$ will refuse $o_{2}$.
-Thus, in that case, it is more efficient that the agent who is assigned to $\varnothing$ is assigned to $o_{2}$ and $a_{1}$ is assigned to $\varnothing$.
-That is, revealing $d_{a_{1}}$ and refusing some unacceptable assignment may make $a_{1}$ better off but result in a wasteful assignment, because some agents may desire the object type refused by $a_{1}$.
-Note that $\left(d_{a_{1}},\hat{p}_{a_{2}},\hat{p}_{a_{3}}\right)$ can be a Nash equilibrium when $r_{a_{2}}=\hat{p}_{a_{2}}$ and $r_{a_{3}}=\hat{p}_{a_{3}}$.
+$a_{2}$ and $a_{3}$ strictly prefer and $a_{1}$ weakly prefers $\tilde{x}$ to $g\left(f^{U}\left(d_{a_{1}},\beta,\beta\right),r\right)$.
+If $a_{1}$ is assigned to $o_{2}$, then either $a_{2}$ or $a_{3}$ will be assigned to $\varnothing$.
+However, in this case, $a_{1}$ will refuse $o_{2}$ and thus $o_{2}$ will be remained unallocated.
+In this case, rather than leaving $o_{2}$ unallocated, it is more efficient to assign $o_{2}$ to the agent ($a_{2}$ or $a_{3}$) who would otherwise receive $\varnothing$ and instead assign $a_{1}$ to $\varnothing$ from the beginning.
+This adjustment ensures that $o_{2}$ is allocated efficiently.
+Hence revealing $d_{a_{1}}$ and refusing an unacceptable assignment may benefit $a_{1}$ but lead to a wasteful allocation, as some agents may prefer the object type that $a_{1}$ refuses.
+Note that $\left(d_{a_{1}}\left(=\beta\right),\beta,\beta\right)$ can be a Nash equilibrium.
 Trivially, $a_{1}$ has no incentive to change $d_{a_{1}}$.
 On the other hand, suppose that both $a_{2}$ and $a_{3}$ prefer the assignment such that each object $o\in\left\{o_{1},o_{2},\varnothing\right\}$ assigns with the probability $1/3$ to the assignment such that $o_{2}$ assigns with probability $1$.
-This ensures that neither $a_{2}$ nor $a_{3}$ has an incentive to change $\hat{p}_{a_{2}}$ and $\hat{p}_{a_{3}}$ to $(o_{2},o_{1},\varnothing)$ or $(o_{2},\varnothing,o_{1})$.
-Then, in this case, $\left(d_{a_{1}},\hat{p}_{a_{2}},\hat{p}_{a_{3}}\right)$ is a Nash equilibrium.
+This ensures that neither $a_{2}$ nor $a_{3}$ has an incentive to change $\beta$ and $\beta$ to $(o_{2},o_{1},\varnothing)$ or $(o_{2},\varnothing,o_{1})$.
+Then, in this case, $\left(d_{a_{1}},\beta,\beta\right)$ is a Nash equilibrium.
 Thus, a wasteful assignment can occur as a Nash equilibrium of agents’ strategies when the assignment is determined by the uniform rank-minimizing mechanism.
@@
 Theorem 1
-Suppose that the uniform rank-minimizing mechanism used. Then, for all agents, truth-telling is weakly strategically dominated by any ODS with the refusal option.
+Suppose that the uniform rank-minimizing mechanism is employed. Then, for all agents, truth-telling is weakly strategically dominated by any ODS with the refusal option.
 By this result, an agent is never hurt by using an ODS instead of truth-telling.
@@
 The condition given in (3) is that there is a truly unacceptable type $o^{\prime}$ such that the number of copies of $o^{\prime}$, plus the sum of the copies of the truly acceptable types, is sufficiently small.
-We discuss the cases where (3) is satisfied for some $a,o,o^{\prime}$ in the subsequent section.
-Note that there is $p_{-a}\in\mathcal{P}^{\left|\mathcal{A}\right|-1}$ such that $R\left(\varnothing,p_{a^{\prime}}\right)=\left|\mathcal{O}\right|$ for all $a^{\prime}\in\mathcal{A\setminus}\left\{a\right\}$ and $a$ strictly prefers $g\left(f^{U}\left(d_{a},p_{-a}\right),r\right)$ to $g\left(f^{U}\left(r_{a},p_{-a}\right),r\right)$.
-Thus, for all $a^{\prime}$, $p_{a^{\prime}}$ may also be an ODS of $a^{\prime}$.
+We discuss real-world scenarios where (3) is satisfied for some $a,o,o^{\prime}$ in the subsequent section.
 Theorems 1 and 2 do not imply that an ODS is a dominant strategy.
 However, if there is $a\in\mathcal{A}$ whose true preference satisfies (3) for some $o,o^{\prime}\in\mathcal{O}$, since truth-telling is strategically dominated by another strategy, $a$ has a strong incentive to manipulate its preference ranking, while one of the ODSs may not necessarily be chosen.
-As stated earlier, there may exist multiple ODSs. In that case, an ODS may be strategically dominated by another one. We have the following result.
-Proposition 3
-Suppose that there is $a\in\mathcal{A}$ whose true preference $r_{a}$ satisfies
-$R\left(o,r_{a}\right)<R\left(\varnothing,r_{a}\right)<R\left(o^{\prime},r_{a}\right)\text{ and }\sum\limits_{\bar{o}:R\left(\bar{o},r_{a}\right)<R\left(\varnothing,r_{a}\right)}q_{\bar{o}}+q_{o^{\prime}}\geq\left|\mathcal{A}\right|,$
-for some $o,o^{\prime}\in\mathcal{O}$.
-Let $d_{a}\in\mathcal{D}_{a}$ be such that $R\left(\varnothing,r_{a}\right)=R\left(o^{\prime},d_{a}\right)$.
-Then, for all $p_{-a}\in\mathcal{P}^{\left|\mathcal{A}\right|-1}$ and all $\bar{o}\in\mathcal{O}$,
-$g\left(f^{U}\left(d_{a},p_{-a}\right),r\right)_{a\bar{o}}=g\left(f^{U}\left(r_{a},p_{-a}\right),r\right)_{a\bar{o}}.$
-In this case, truth-telling is not strategically dominated by an ODS of revealing such $d_{a}$.
-Note that revealing the full extension may be such an ODS, when $o^{\prime}$ defined in Proposition 3 satisfies $R\left(\varnothing,r_{a}\right)+1=R\left(o^{\prime},r_{a}\right)$.
-Moreover, by Theorem 1 and Proposition 3, we show that for an agent, revealing the full extension may be strategically dominated by another ODS. First, we provide an example to briefly see this fact.
+As stated earlier, multiple ODSs may exist.
+Even if an ODS strategically dominates truth-telling, another ODS that does not strategically dominate truth-telling may also exist.
+In such a case, the latter ODS is strategically dominated by the former.
+The strategy of revealing the full extension is not an exception to this.
+We show this fact by introducing the following example.
 Example 3
 Let $\mathcal{A=}\left\{a_{1},a_{2},a_{3}\right\}$ and $\mathcal{O=}\left\{o_{1},o_{2},o_{3},\varnothing\right\}$.
-Suppose $q_{o_{1}}=q_{o_{3}}=1,$ $q_{o_{2}}=2$ and $r_{a_{1}}=(o_{1},\varnothing,o_{2},o_{3})$.
+Let $\gamma=(o_{1},o_{2},o_{3},\varnothing),$ $\gamma^{\prime}=(o_{1},o_{3},o_{2},\varnothing)$ and $\gamma^{\prime\prime}=(o_{3},o_{1},o_{2},\varnothing).$ Suppose $q_{o_{1}}=q_{o_{3}}=1,$ $q_{o_{2}}=2$ and $r_{a_{1}}=(o_{1},\varnothing,o_{2},o_{3})$.
 Then, there are two ODSs.
-One is $e_{a_{1}}=(o_{1},o_{2},o_{3},\varnothing)$, which is the full extension of $a_{1}$, and the other is $d_{a_{1}}=(o_{1},o_{3},o_{2},\varnothing)$.
-First, suppose
-$\bar{p}_{a_{2}}=(o_{1},o_{3},o_{2},\varnothing)\text{ and }\bar{p}_{a_{3}}=(o_{3},o_{1},o_{2},\varnothing).$
-Then,
-$f^{U}\left(r_{a_{1}},\bar{p}_{a_{2}},\bar{p}_{a_{3}}\right)=y=\left(\begin{array}[]{cccc}y_{a_{1}o_{1}}&y_{a_{1}o_{2}}&y_{a_{1}o_{3}}&y_{a_{1}\varnothing}\\ y_{a_{2}o_{1}}&y_{a_{2}o_{2}}&y_{a_{2}o_{3}}&y_{a_{2}\varnothing}\\ y_{a_{3}o_{1}}&y_{a_{3}o_{2}}&y_{a_{3}o_{3}}&y_{a_{3}\varnothing}\end{array}\right)=\left(\begin{array}[]{cccc}0&0&0&1\\ 1&0&0&0\\ 0&0&1&0\end{array}\right),$
+One is $e_{a_{1}}=\gamma$, which is the full extension of $a_{1}$, and the other is $d_{a_{1}}=\gamma^{\prime}$.
+In this example, we assume $\left(p_{a_{2}},p_{a_{3}}\right)=\left(r_{a_{2}},r_{a_{3}}\right);$ that is, $a_{2}$ and $a_{3}$ always adopt truth-telling.
+First, suppose $p_{a_{2}}=\gamma^{\prime}$ and $p_{a_{3}}=\gamma^{\prime\prime}.$ Then,
+$f^{U}\left(r_{a_{1}},\gamma^{\prime},\gamma^{\prime\prime}\right)=y=\left(\begin{array}[]{cccc}y_{a_{1}o_{1}}&y_{a_{1}o_{2}}&y_{a_{1}o_{3}}&y_{a_{1}\varnothing}\\ y_{a_{2}o_{1}}&y_{a_{2}o_{2}}&y_{a_{2}o_{3}}&y_{a_{2}\varnothing}\\ y_{a_{3}o_{1}}&y_{a_{3}o_{2}}&y_{a_{3}o_{3}}&y_{a_{3}\varnothing}\end{array}\right)=\left(\begin{array}[]{cccc}0&0&0&1\\ 1&0&0&0\\ 0&0&1&0\end{array}\right),$
 In this example, we write the assignments in this manner. Moreover,
-$g\left(f^{U}\left(d_{a_{1}},\bar{p}_{a_{2}},\bar{p}_{a_{3}}\right),r\right)=\left(\begin{array}[]{cccc}1/2&0&0&1/2\\ 1/2&1/2&0&0\\ 0&0&1&0\end{array}\right)$
-Therefore, $a_{1}$ (strictly) prefers $g\left(f^{U}\left(d_{a_{1}},\bar{p}_{a_{2}},\bar{p}_{a_{3}}\right),r\right)$ to $g\left(f^{U}\left(r_{a_{1}},\bar{p}_{a_{2}},\bar{p}_{a_{3}}\right),r\right)$.
+$g\left(f^{U}\left(d_{a_{1}},\gamma^{\prime},\gamma^{\prime\prime}\right),r\right)=\left(\begin{array}[]{cccc}1/2&0&0&1/2\\ 1/2&1/2&0&0\\ 0&0&1&0\end{array}\right)\text{.}$
+Therefore, $a_{1}$ (strictly) prefers $g\left(f^{U}\left(d_{a_{1}},\gamma^{\prime},\gamma^{\prime\prime}\right),r\right)$ to $g\left(f^{U}\left(r_{a_{1}},\gamma^{\prime},\gamma^{\prime\prime}\right),r\right)$.
 On the other hand, there is no pair $\left(p_{a_{2}},p_{a_{3}}\right)$ of preference orders of $a_{2}$ and $a_{3}$ such that $a_{1}$ (strictly) prefers $g\left(f^{U}\left(e_{a_{1}},p_{a_{2}},p_{a_{3}}\right),r\right)$ to $g\left(f^{U}\left(r_{a_{1}},p_{a_{2}},p_{a_{3}}\right),r\right)$.
-This is because, by Proposition 3 and $q_{o_{1}}+q_{o_{2}}=3=\left|\mathcal{A}\right|$, for any $\left(p_{a_{2}},p_{a_{3}}\right)\in\mathcal{P}^{2}$ and any $\bar{o}\in\mathcal{O}$,
+This is because, by $q_{o_{1}}+q_{o_{2}}=3=\left|\mathcal{A}\right|$, for any $\left(p_{a_{2}},p_{a_{3}}\right)\in\mathcal{P}^{2}$ and any $\bar{o}\in\mathcal{O}$,
 $\left(g\left(f^{U}\left(e_{a_{1}},p_{a_{2}},p_{a_{3}}\right),r\right)\right)_{a_{1}\bar{o}}=\left(g\left(f^{U}\left(r_{a_{1}},p_{a_{2}},p_{a_{3}}\right),r\right)\right)_{a_{1}\bar{o}}\text{.}$
@@
 On the other hand, revealing $o_{3}$ as a second-best type increases it, when $o_{3}$ is the first-best type of another agent.
+This is because, to minimize the total rank, $o_{3}$ must be assigned to an agent who ranks it as their first-best type.
+However, since $a_{1}$ ranks $o_{2}$ and $\varnothing$ lower, minimizing the total rank requires increasing the probability that $o_{1}$ is allocated to $a_{1}$.
 Thus, revealing the full extension may be strategically dominated by another ODS when an unacceptable type (in this case $o_{3}$) is less preferable than the other (in this case $o_{2}$) but the number of copies of the former is less than that of the latter.
-Generally, we have the following result.
-Lemma 4
-Suppose that there is $a\in\mathcal{A}$ whose true preference satisfies
-$R\left(o,r_{a}\right)<R\left(\varnothing,r_{a}\right)=R\left(o^{\prime},r_{a}\right)-1<R\left(o^{\prime\prime},r_{a}\right),\text{ }$
-$\sum\limits_{\bar{o}:R\left(\bar{o},r_{a}\right)<R\left(\varnothing,r_{a}\right)}q_{\bar{o}}+q_{o^{\prime}}\geq\left|\mathcal{A}\right|,\text{ and}$
-$\sum\limits_{\bar{o}:R\left(\bar{o},r_{a}\right)<R\left(\varnothing,r_{a}\right)}q_{\bar{o}}+q_{o^{\prime\prime}}<\left|\mathcal{A}\right|,$
-[…]
-$R\left(\varnothing,r_{a}\right)=R\left(o^{\prime},e_{a}\right)=R\left(o^{\prime\prime},d_{a}\right),$
-with the refusal option.
-Note that Lemma 4 does not contradict a result of Featherstone (2020, Proposition 13).
-In his model, he assumes that the agents have no information on the capacities of object types.
+Note that this fact does not contradict a result of Featherstone (2020, Proposition 13).
+In their model, they assume that the agents have no information on the capacities of object types.
 Revealing the full extension can be strategically dominated by another ODS only when the capacities of the unacceptable types are different and agents know them.
-Thus, the strategic domination in Lemma 4 does not occur in the low-information situation discussed by Featherstone (2020).
-Next, we show that if agents adopt an ODS, then the result may be wasteful and thus inefficient.
+Thus, the strategic domination in Example 3 does not occur in the low-information situation discussed by Featherstone (2020).
+Next, we show that if some agent adopts an ODS, then the result may be wasteful and thus inefficient.
 Proposition 3
@@
 By Theorem 2, for an agent $a$ with $r_{a}$ that satisfy (3) for some $o,o^{\prime}\in\mathcal{O}$, an ODS dominates truth-telling.
-Moreover, by Proposition 1, in the same case, the strategy may make the assignment wasteful and thus inefficient.
+Moreover, by Proposition 3, in the same case, the strategy may make the assignment wasteful.
 In an ODS, an agent reveals actually unacceptable types as acceptable types.
@@
 Since the type may be desirable for another agent, in that case, the assignment is wasteful.
-By Theorems 1 and 2 and Proposition 3, if the uniform rank-minimizing mechanism is used, the agents may have a strong incentive to manipulate their preferences, and these manipulations may result in an inefficient assignment.
-Thus, next, we consider another fair rank-minimizing mechanism such that, contrary to $f^{U}$, if it is used as the mechanism in the second stage, then for any agent, truth-telling is not strategically dominated by any other strategy.
+By Theorems 1 and 2 and Proposition 3, if the uniform rank-minimizing mechanism is employed, the agents may have a strong incentive to manipulate their preferences, and these manipulations may result in an inefficient assignment.
+Thus, next, we consider another fair rank-minimizing mechanism such that, contrary to $f^{U}$, if it is employed as the mechanism in the second stage, then for any agent, truth-telling is not strategically dominated by any other strategy.
 We let $f^{M}$ be the modified uniform rank-minimizing mechanism such that if $p$ satisfies
-for agent $a,$ $R\left(o,p_{a}\right)=1$ and $R\left(\varnothing,p_{a}\right)\geq 3,$
+there is one agent represented by $a$ such that $R\left(o,p_{a}\right)=1$ and $R\left(\varnothing,p_{a}\right)\geq 3,$
 there are $q_{o}$ agents represented by $a^{\prime}$ revealing $p_{a^{\prime}}$, such that $R\left(\varnothing,p_{a^{\prime}}\right)=l\geq 2,$ $p_{a}^{l^{\prime}}=p_{a^{\prime}}^{l}$ for all $l^{\prime}=1,\cdots,l-1,$ $R\left(\varnothing,p_{a}\right)>l,$ and $\bar{k}(p_{a^{\prime}})=l$; that is,
@@
 $\left(f^{M}\left(p\right)\right)_{ao}=0,\text{ }\left(f^{M}\left(p\right)\right)_{a^{\prime}o}=1,\text{ }\left(f^{M}\left(p\right)\right)_{ap_{a}^{2}}=1\text{ and }\left(f^{M}\left(p\right)\right)_{a^{\prime\prime}\varnothing}=1,$
-and if $p$ does not satisfy all of them, then
+and otherwise; that is, if $p$ does not satisfy all of them, then
 $f^{M}\left(p\right)=\sum\limits_{y^{\ast}\in\mathcal{Y}^{\ast}\left(p\right)}\frac{y^{\ast}}{\left|\mathcal{Y}^{\ast}\left(p\right)\right|}=f^{U}\left(p\right).$
@@
 First, suppose $q_{o_{1}}=q_{o_{2}}=1$.
-If $\left(p_{a_{1}},p_{a_{2}},p_{a_{3}}\right)=\left(\psi^{\prime},\psi,\psi\right)$, $\left(\psi,\psi^{\prime},\psi\right)$ or $\left(\psi,\psi,\psi^{\prime}\right)$, then this satisfies (i), (ii) and (iii).
-In these cases, a student who chooses $\psi^{\prime}$ is corresponding to $a,$ another student is corresponding to $a^{\prime}$, and $o_{1}$ is corresponding to $o$.
-The rank-minimizing mechanism $f$ introduced in Example 1 is the modified uniform rank-minimizing mechanism, because a student who chooses $\psi^{\prime}$ obtains $o_{1}$ with probability 00 and another student obtains $o_{1}$ with probability $1$
+If $\left(p_{a_{1}},p_{a_{2}},p_{a_{3}}\right)=\left(\alpha^{\prime},\alpha,\alpha\right)$, $\left(\alpha,\alpha^{\prime},\alpha\right)$ or $\left(\alpha,\alpha,\alpha^{\prime}\right)$, then $\left(p_{a_{1}},p_{a_{2}},p_{a_{3}}\right)$ satisfies (i), (ii) and (iii).
+In these cases, the student who chooses $\alpha^{\prime}$ is corresponding to $a,$ another student is corresponding to $a^{\prime}$, and $o_{1}$ is corresponding to $o$.
+The rank-minimizing mechanism $f^{\ast}$ introduced in Example 1 is the modified uniform rank-minimizing mechanism, because a student who chooses $\alpha^{\prime}$ obtains $o_{1}$ with probability 00 and another student obtains $o_{1}$ with probability $1$
 On the other hand, suppose $q_{o_{1}}=1$ and $q_{o_{2}}\geq 2$.
-Even if $\left(p_{a_{1}},p_{a_{2}},p_{a_{3}}\right)=\left(\psi^{\prime},\psi,\psi\right)$, $\left(\psi,\psi^{\prime},\psi\right)$ or $\left(\psi,\psi,\psi^{\prime}\right)$, this does not satisfy (ii), because in this case, $p_{a^{\prime}}=\psi$ and $l=3.$ Then,
+Even if $\left(p_{a_{1}},p_{a_{2}},p_{a_{3}}\right)=\left(\alpha^{\prime},\alpha,\alpha\right)$, $\left(\alpha,\alpha^{\prime},\alpha\right)$ or $\left(\alpha,\alpha,\alpha^{\prime}\right)$, this does not satisfy (ii).
+This is because in this case, $p_{a^{\prime}}=\alpha$ and $l=3,$ but
 $\sum\limits_{\bar{o}:R\left(\bar{o},p_{a^{\prime}}\right)<R\left(\varnothing,p_{a^{\prime}}\right)}q_{\bar{o}}=q_{o_{1}}+q_{o_{2}}=3=\left|\mathcal{A}\right|.$
-Therefore, (ii) is not satisfied with $p_{a^{\prime}}=\psi$.
+Therefore, (ii) is not satisfied with $p_{a^{\prime}}=\alpha$ in the case of $q_{o_{1}}=1$ and $q_{o_{2}}\geq 2$.
 We have the following result.
@@
 The modified uniform rank-minimizing mechanism $f^{M}$ satisfies ETE.
-The strategic issues, mentioned in Theorem 2, do not arise under this mechanism $f^{M}$.
+The strategic issues, mentioned in Theorems 1 and 2, do not arise under this mechanism $f^{M}$.
 First, we explain why the strategic issue mentioned in Example 2 does not arise in the modified uniform rank-minimizing mechanism.
 If $r_{a}=p_{a^{\prime}}$ where $p_{a^{\prime}}$ is defined in (ii), then $\left(d_{a},p_{-a}\right)$ satisfies (i), (ii) and (iii) where $d_{a}\in\mathcal{D}_{a}$.
-This implies that for some $p_{-a}$, if agent $a$ adopts an ODS instead of truth-telling and $f^{M}$ is used, then $a$ is worse off.
+This implies that for some $p_{-a}$, if agent $a$ adopts an ODS instead of truth-telling and $f^{M}$ is employed, then $a$ is worse off.
 To see this fact, we revisit Example 2.
-In this example, $\left(d_{a_{1}},p_{a_{2}}^{\prime},p_{a_{3}}^{\prime}\right)$ satisfies (i), (ii) and (iii), but $\left(r_{a_{1}},p_{a_{2}}^{\prime},p_{a_{3}}^{\prime}\right)$ does not.
+In this example, $\left(d_{a_{1}},p_{a_{2}},p_{a_{3}}\right)=\left(\beta,\beta^{\prime\prime},\beta^{\prime\prime}\right)$ satisfies (i), (ii) and (iii), but $\left(r_{a_{1}},p_{a_{2}},p_{a_{3}}\right)=\left(\beta^{\prime\prime},\beta^{\prime\prime},\beta^{\prime\prime}\right)$ does not.
 Thus,
-$f^{M}\left(r_{a_{1}},p_{a_{2}}^{\prime},p_{a_{3}}^{\prime}\right)$ $=$ $f^{U}\left(r_{a_{1}},p_{a_{2}}^{\prime},p_{a_{3}}^{\prime}\right)=\left(\begin{array}[]{ccc}1/3&0&2/3\\ 1/3&0&2/3\\ 1/3&0&2/3\end{array}\right),$
-$f^{M}\left(d_{a_{1}},p_{a_{2}}^{\prime},p_{a_{3}}^{\prime}\right)$ $=$ $\left(\begin{array}[]{ccc}0&1&0\\ 1/2&0&1/2\\ 1/2&0&1/2\end{array}\right),$
-$g\left(f^{M}\left(d_{a_{1}},p_{a_{2}}^{\prime},p_{a_{3}}^{\prime}\right)\right)$ $=$ $\left(\begin{array}[]{ccc}0&0&1\\ 1/2&0&1/2\\ 1/2&0&1/2\end{array}\right).$
+$f^{M}\left(r_{a_{1}},\beta^{\prime\prime},\beta^{\prime\prime}\right)$ $=$ $f^{U}\left(r_{a_{1}},\beta^{\prime\prime},\beta^{\prime\prime}\right)=\left(\begin{array}[]{ccc}1/3&0&2/3\\ 1/3&0&2/3\\ 1/3&0&2/3\end{array}\right),$
+$f^{M}\left(d_{a_{1}},\beta^{\prime\prime},\beta^{\prime\prime}\right)$ $=$ $\left(\begin{array}[]{ccc}0&1&0\\ 1/2&0&1/2\\ 1/2&0&1/2\end{array}\right),$
+$g\left(f^{M}\left(d_{a_{1}},\beta^{\prime\prime},\beta^{\prime\prime}\right)\right)$ $=$ $\left(\begin{array}[]{ccc}0&0&1\\ 1/2&0&1/2\\ 1/2&0&1/2\end{array}\right).$
 This implies that $a_{1}$ may be worse off by revealing $d_{a_{1}}$ instead of $r_{a_{1}}$, even if $a_{1}$ refuses any unacceptable objects.
-Therefore, $r_{a_{1}}$ is not strategically dominated by $d_{a_{1}}$, if $f^{M}$ is used.
+Therefore, $r_{a_{1}}$ is not strategically dominated by $d_{a_{1}}$, if $f^{M}$ is employed.
 Formally, we have the following fact.
 Proposition 5
-Suppose that the modified uniform rank-minimizing mechanism is used.
+Suppose that the modified uniform rank-minimizing mechanism is employed.
 For any $r_{a}\in\mathcal{P}$, there is no $p_{a}\in\mathcal{P}$ that strategically dominates $r_{a}$ with (or without) the refusal option.
-By adopting $f^{M}$, any ODS can fail, because increasing the number of acceptable objects may lead the agent to be unassigned acceptable objects in their true preference.
+By adopting $f^{M}$, any ODS can fail because increasing the number of acceptable objects may lead the agent to be unassigned to truly acceptable objects, depending on the strategy profile of the other agents.
 Thus, in some profiles of other agents’ (revealed) preferences, truth-telling is better than adopting any ODS.
@@
 Example 4
-Let $R\left(\varnothing,r_{a}\right)=k$.
-Suppose $k\geq 3$; that is, $\varnothing$ is neither the first-best type nor the second-best type for $r_{a}$.
-Thus, we let $r_{a}^{1}=o$ and $r_{a}^{2}=o^{\prime}$.
-For $p_{-a}$, there are $q_{o}$ agents represented by $a^{\prime}$ whose revealed preference order is such that $p_{a^{\prime}}^{k^{\prime}}=r_{a^{\prime}}^{k^{\prime}}$ for all $k^{\prime}=1,\cdots,k-2$ and $p_{a^{\prime}}^{k-1}=\varnothing$.
-We assume
-$q_{r_{a}^{1}}+\cdots+q_{r_{a}^{k-2}}<\left|\mathcal{A}\right|\text{.}$
-Moreover, for $p_{-a}$ and any other agent represented by $a^{\prime\prime}$, $R\left(\varnothing,r_{a^{\prime\prime}}\right)=1$.
-In this case, since $\left(r_{a},p_{-a}\right)$ satisfies (i), (ii) and (iii),
-$\left(f^{M}\left(r_{a},p_{-a}\right)\right)_{ao}$ $=$ 0, ($0,\text{ }\left(f^{M}\left(r_{a},p_{-a}\right)\right)_{ao^{\prime}}=1,$
-$\left(f^{U}\left(r_{a},p_{-a}\right)\right)_{ao}$ $=$ $\frac{q_{o}}{q_{o}+1},$
-$\left(f^{U}\left(r_{a},p_{-a}\right)\right)_{ao^{\prime}}$ $=$ $\frac{1}{q_{o}+1}.$
-On the other hand, we let $p_{a}=p_{a^{\prime}}$. Since $\left(p_{a},p_{-a}\right)$ does not satisfy all of (i), (ii) and (iii),
-$\left(f^{M}\left(p_{a},p_{-a}\right)\right)_{ao}$ $=$ $\left(f^{U}\left(p_{a},p_{-a}\right)\right)_{ao}=\frac{q_{o}}{q_{o}+1},$
-$\left(f^{M}\left(p_{a},p_{-a}\right)\right)_{ao^{\prime}}$ $=$ $\left(f^{U}\left(p_{a},p_{-a}\right)\right)_{ao^{\prime}}=\frac{1}{q_{o}+1}.$
-Thus, in this case, if $f^{M}$ is used, then $a$ has an incentive to choose $p_{a}$.
-In other words, in this case, $a$ has an incentive to reveal by reducing its truly acceptable types.
-On the other hand, if $f^{U}$ is used, then $a$ has no such an incentive.
-Of course, this does not mean that truth-telling is strategically dominated revealing some reduction of $r_{a}$. However, agents may have such incentives even if they do not have the refusal option.
+Let $\mathcal{A=}\left\{a_{1},a_{2}\right\}$ and $\mathcal{O=}\left\{o_{1},o_{2},\varnothing\right\}$.
+Let $\delta=\left(o_{1},o_{2},\varnothing\right)$ and $\delta^{\prime}=\left(o_{1},\varnothing,o_{2}\right)$.
+We assume $q_{o}=1$.
+Suppose $r_{a_{1}}=\delta$ and $p_{a_{2}}=\delta$.
+In this case, since $\left(\delta(=r_{a_{1}}),\delta\right)$ does not satisfy all of (i), (ii) and (iii),
+$\left(f^{M}\left(\delta,\delta\right)\right)_{a_{1}o_{1}}$ $=$ $\left(f^{M}\left(\delta,\delta\right)\right)_{a_{1}o_{2}}=\frac{1}{2},$
+$\left(f^{U}\left(\delta,\delta\right)\right)_{a_{1}o_{1}}$ $=$ $\left(f^{U}\left(\delta,\delta\right)\right)_{a_{1}o_{2}}=\frac{1}{2}.$
+On the other hand, since $\left(\delta^{\prime}(\neq r_{a_{1}}),\delta\right)$ satisfies all of (i), (ii) and (iii),
+$\left(f^{M}\left(\delta^{\prime},\delta\right)\right)_{a_{1}o_{1}}$ $=$ 1, ($1,\text{ }\left(f^{U}\left(\delta^{\prime},\delta\right)\right)_{a_{1}o_{2}}=0,$
+$\left(f^{U}\left(\delta^{\prime},\delta\right)\right)_{a_{1}o_{1}}$ $=$ $\left(f^{U}\left(\delta^{\prime},\delta\right)\right)_{a_{1}o_{2}}=\frac{1}{2}.$
+Thus, in this case, if $f^{M}$ is employed, then $a_{1}$ has an incentive to choose $\delta^{\prime}$ instead of $r_{a_{1}}=\delta$.
+In other words, in this case, $a_{1}$ has an incentive to reveal by reducing its truly acceptable types.
+On the other hand, if $f^{U}$ is employed, then $a_{1}$ has no such an incentive.
+Of course, this does not mean that truth-telling is strategically dominated revealing some reduction of $r_{a_{1}}$.
+However, agents may have such incentives even if they do not have the refusal option.
 5 Concluding Remarks
-We show that if an assignment is determined by the uniform rank-minimizing mechanism, which is a fair one, and agents can refuse the assignment and obtain the outside option instead, then the true preference is strategically dominated by an ODS.
+We show that if an assignment is determined by the uniform rank-minimizing mechanism, which is a fair one, and agents can refuse the assignment and obtain the outside option instead, then truth-telling is strategically dominated by an ODS.
 The adoption of such strategies by agents may result in inefficient assignments; that is, an agent who adopts an ODS refuses their assigned type, although some other agents desire that type.
@@
 Our result implies that a student may have an incentive to reveal truly unacceptable schools as acceptable schools, and they may be assigned to one of them and then refuse the admission even if some other students desire the school.
-In such a case, the school may re-enroll students, but even in that case, it would cause market disruption.
+In such a case, the school might re-enroll students, but even in that case, it would cause market disruption.
 Our result implies that if all agents prefer many other types over the outside option, then the strategic problem considered in this study is not so severe.
@@
 (2013), because it is fair and ordinally efficient, although this efficiency property is weaker than the rank-minimizing.
+Finally, our study does not address reallocation after assignment refusals.
+However, Feigenbaum et al.
+(2020) and Afacan (2022) consider dynamic markets, such as school choice markets with a second-round admission process, where reallocation after refusals is possible.
+Incorporating such a framework could enhance the practical applicability of our findings.
+Specifically, analyzing how refused objects are redistributed and how agents’ strategic behavior changes in response could provide new insights into the design of rank-minimizing mechanisms.
 Appendix
@@
 Proof of Proposition 3
-Suppose that there is $a\in\mathcal{A}$ whose true preference $r_{a}$ satisfies
-$R\left(o,r_{a}\right)$ $<$ $R\left(\varnothing,r_{a}\right)<R\left(o^{\prime},r_{a}\right)$
-$\sum\limits_{\bar{o}:R\left(\bar{o},r_{a}\right)<R\left(\varnothing,r_{a}\right)}q_{\bar{o}}+q_{o^{\prime}}$$\geq$ $\left|\mathcal{A}\right|,$ (9)
-for some $o,o^{\prime}\in\mathcal{O}$.
-Fix $p_{-a}\in\mathcal{P}^{\left|\mathcal{A}\right|-1}$ arbitrarily.
-Moreover, we let $d_{a}\in\mathcal{D}_{a}$ be such that $R\left(\varnothing,r_{a}\right)=R\left(o^{\prime},d_{a}\right)$.
-We show the following fact.
-Claim 3
-Let $y$ and $y^{\prime}$ be such that $y_{a\varnothing}=y_{ao^{\prime}}^{\prime}=1$ and $y_{a^{\prime}\bar{o}}=y_{a^{\prime}\bar{o}}^{\prime}$ for all $a^{\prime}\in\mathcal{A}$ and all $\bar{o}\in\mathcal{O}$.
-Then, $y\in\mathcal{Y}^{\ast}\left(r_{a},p_{-a}\right)$ if and only if $y^{\prime}\in\mathcal{Y}^{\ast}\left(d_{a},p_{-a}\right)$.
-Proof. By the construction,
-$RV\left(y,\left(r_{a},p_{-a}\right)\right)=RV\left(y^{\prime},\left(d_{a},p_{-a}\right)\right)\text{.}$
-First, suppose $y\in\mathcal{Y}^{\ast}\left(r_{a},p_{-a}\right)$.
-Then, by (9), $y^{\prime}$ is an allocation.
-Suppose not; that is, there is $y^{\prime\prime}\in\mathcal{Y}$ such that $RV\left(y^{\prime\prime},\left(d_{a},p_{-a}\right)\right)<RV\left(y^{\prime},\left(d_{a},p_{-a}\right)\right)$.
-Then, by (9), $o^{\prime\prime}$ such that $y_{ao^{\prime\prime}}^{\prime\prime}=1$ satisfies $R\left(o,d_{a}\right)\leq R\left(o^{\prime},d_{a}\right)$.
-Let
-$o^{\prime\prime\prime}=o^{\prime\prime}\text{ if }o\neq o^{\prime},\text{ }o^{\prime\prime\prime}=\varnothing\text{ if }o^{\prime\prime}=o^{\prime}\text{.}$
-Moreover, let $y^{\prime\prime\prime}$ be such that $y_{ao^{\prime\prime\prime}}^{\prime\prime\prime}=1$ and $y_{a^{\prime}\bar{o}}^{\prime\prime\prime}=y_{a^{\prime}\bar{o}}^{\prime\prime}$ for all $a^{\prime}\in\mathcal{A}$ and all $\bar{o}\in\mathcal{O}$.
-Then,
-$RV\left(y^{\prime\prime\prime},\left(r_{a},p_{-a}\right)\right)\leq RV\left(y^{\prime\prime},\left(d_{a},p_{-a}\right)\right)<RV\left(y^{\prime},\left(d_{a},p_{-a}\right)\right)=RV\left(y,\left(r_{a},p_{-a}\right)\right)$
-contradicting $y\in\mathcal{Y}^{\ast}\left(r_{a},p_{-a}\right)$.
-Second, suppose $y^{\prime}\in\mathcal{Y}^{\ast}\left(d_{a},p_{-a}\right)$.
-Suppose not; that is, there is $y^{\prime\prime}\in\mathcal{Y}$ such that $RV\left(y^{\prime\prime},\left(r_{a},p_{-a}\right)\right)<RV\left(y,\left(r_{a},p_{-a}\right)\right)$.
-Then, by (9), $o^{\prime\prime}\in\mathcal{O}$ such that $y_{ao^{\prime\prime}}^{\prime\prime}=1$ and $R\left(o^{\prime\prime},r_{a}\right)\leq R\left(\varnothing,r_{a}\right)$.
-Let
-$o^{\prime\prime\prime}=o^{\prime\prime}\text{ if }o^{\prime\prime}\neq\varnothing,\text{ }o^{\prime\prime\prime}=\bar{o}\text{ if }o^{\prime\prime}=\varnothing\text{,}$
-where $\bar{o}$ satisfies $\sum y_{a\bar{o}}^{\prime\prime}<q_{\bar{o}}$ and $R\left(\bar{o},d_{a}\right)\leq R\left(o^{\prime},d_{a}\right)$.
-There exist such an $\bar{o}$, because (9) is satisfied.
-Then, let $y^{\prime\prime\prime}$ be such that $y_{ao^{\prime\prime\prime}}^{\prime\prime\prime}=1$ and $y_{a^{\prime}\bar{o}}^{\prime\prime\prime}=y_{a^{\prime}\bar{o}}^{\prime\prime}$ for all $a^{\prime}\in\mathcal{A}$ and all $\bar{o}\in\mathcal{O}$.
-Then,
-$RV\left(y^{\prime\prime\prime},\left(d_{a},p_{-a}\right)\right)\leq RV\left(y^{\prime\prime},\left(r_{a},p_{-a}\right)\right)<RV\left(y^{\prime},\left(r_{a},p_{-a}\right)\right)=RV\left(y,\left(d_{a},p_{-a}\right)\right)$
-contradicting $y\in\mathcal{Y}^{\ast}\left(d_{a},p_{-a}\right)$.
-Therefore, $y\in\mathcal{Y}^{\ast}\left(r_{a},p_{-a}\right)$ if and only if $y^{\prime}\in\mathcal{Y}^{\ast}\left(d_{a},p_{-a}\right)$. Q.E.D.
-By Claims 1 and 3, we have Proposition 3.
-Proof of Lemma 4
-We assume that the assumptions of Lemma 4 is satisfied. First, by Proposition 3, for all $p_{-a}\in\mathcal{P}^{\left|\mathcal{A}\right|-1}$ and all $\bar{o}\in\mathcal{O}$,
-$g\left(f^{U}\left(e_{a},p_{-a}\right),r\right)_{a\bar{o}}=g\left(f^{U}\left(r_{a},p_{-a}\right),r\right)_{a\bar{o}}.$
-Second, by Theorem 2, $r_{a}$ is strategically dominated by $d_{a}$ with the refusal option. Therefore, $e_{a}$ is also strategically dominated by $d_{a}$ with the refusal option.
-Proof of Proposition 4
 We use the example introduced in the proof of Theorem 2.

### v6

@@
 Abstract
-This study examines strategic issues in fair rank-minimizing mechanisms, which choose an assignment that minimizes the average rank of object types to which agents are assigned and satisfy a fairness property called equal treatment of equals.
-As one of these fair mechanisms, the uniform rank-minimizing mechanism is considered.
-We focus on the case where agents can refuse their assignment and obtain the outside option instead.
-Without the refusal option, truth-telling is not strategically dominated by any strategies if a fair rank-minimizing mechanism is employed.
-However, if agents have the option and the uniform rank-minimizing mechanism is employed, then a strategy called an outside option demotion strategy strategically dominates truth-telling.
-Moreover, we show that adopting this strategy may lead to inefficient assignments.
-To counter this, we consider a modification of the uniform rank-minimizing mechanism, though it may lead agents to strategically reduce the number of acceptable types.
+This paper investigates the strategic implications of the uniform rank-minimizing mechanism (URM), an assignment rule that selects uniformly from the set of deterministic assignments minimizing the sum of agents’ reported ranks.
+We focus on settings in which agents may refuse their assignment and instead receive an outside option.
+Without the refusal option, we show that truth-telling is not strictly dominated under any fair rank-minimizing mechanism; that is, one satisfying equal treatment of equals.
+However, introducing the refusal option significantly changes strategic incentives: specific manipulations, called outside option demotion strategies, dominate truth-telling under the URM.
+Moreover, such manipulations can lead to inefficient outcomes, as desirable objects may be refused by misreporting agents and consequently remain unassigned.
+To address this issue, we propose a modification of the URM that restores undominated truth-telling, although it introduces incentives to underreport acceptable objects.
+Our results highlight a fundamental trade-off in the design of fair rank-minimizing mechanisms when agents can refuse their assignments.
 JEL classification: C78; D47
@@
 1 Introduction
-In this study, we discuss the strategic issues of fair rank-minimizing allocation mechanisms mainly when agents have the option to refuse their assignment determined by one of the mechanisms and obtain their outside option instead.
-There are several agents and types of object, including a null object type, which corresponds to the outside option.
-In a mechanism, each agent reveals one preference ranking, which may or may not be their true preference ranking.
-In rank-minimizing mechanisms, assignments are determined as to minimize the total (or average) revealed rank of the objects to which the agents are assigned.
-The rank distribution of assignments is a critical consideration for many real-world matching authorities.
-For instance, Featherstone (2020) notes that Teach For America takes rank distribution into account when making assignment decisions.
-Consequently, rank-minimizing mechanisms are discussed in recent studies, including Featherstone (2020), Nikzad (2022), Ortega and Klein (2023), Feizi (2024), and Troyan (2024).
-Several of these studies highlight the advantages of rank-minimizing mechanisms, particularly regarding the efficiency of assignments.
-By the results of Featherstone (2020) and Feizi (2024), any rank-minimizing mechanism satisfies other efficiency properties, such as ordinal efficiency, which is first introduced by Bogomolnaia and Moulin (2001).
-Moreover, Ortega and Klein (2023) compare the outcomes of rank-minimizing mechanisms with those of the DA mechanism and the TTC mechanism, showing that rank-minimizing mechanisms improve upon the others in several aspects of efficiency.
-In this study, we focus on rank-minimizing mechanisms that satisfy the fairness property of equal treatment of equals, which ensures that agents revealing identical preference rankings are assigned to each object type with the same probability.
-[^1] Yokote et al.
-(2019) mention that this fairness property is a classical concept discussed by Aristotle in Book V of Nicomachean Ethics.
-This fairness property is discussed by Bogomolnaia and Moulin (2001) and many other previous studies on various allocation problems (Thomson 2011).
-We consider a slightly stronger version of this property.
-We call a rank-minimizing mechanism satisfying equal treatment of equals a fair rank-minimizing mechanism.
-To consider these mechanisms, we adopt the probabilistic assignment framework.
-Among the mechanisms, we first consider a representative one called the uniform rank-minimizing mechanism.
-Since Nikzad (2022), Ortega and Klein (2023) and Troyan (2024) also mainly consider this mechanism, at least in the literature, this is the most common rank-minimizing mechanism.
-We discuss the strategic manipulation incentives of agents when the assignment is determined by a fair rank-minimizing mechanism.
-Specifically, we consider the case where agents can refuse their assignment, and if they refuse it, then the outside option is assigned to them.
-First, as a benchmark case, we discuss the situation where agents cannot refuse their assignment.
-This case is discussed by many previous studies, such as those mentioned above.
-Featherstone (2020) shows that any rank-minimizing mechanism is not strategy-proof.
-On the other hand, Troyan (2024) shows that the uniform rank-minimizing mechanism is not obviously manipulable, which is a strategic property introduced by Troyan and Morrill (2020).
-Moreover, Nikzad (2022) shows that the uniform rank-minimizing mechanism is Bayesian incentive compatible in uniform markets, which he primarily discusses.
-Related to but logically independent to their results, we show that, in this case, truth-telling is not strategically dominated by any untrue preference under any fair rank-minimizing mechanism.
-These results imply that in this case, strategic manipulation is a critical problem of rank-minimizing mechanisms.
-Next, as the main analysis, we consider the case where agents have the refusal options.
+Many assignment problems such as school choice or job assignment require matching agents with heterogeneous objects based on ordinal preferences.
+In such contexts, policymakers often seek to minimize the sum (or average) of the ranks of assigned objects.
+For instance, Featherstone (2020) observes that Teach For America explicitly considers rank distribution in its assignment decisions.
+However, the authority can optimize only the ranks reported by individuals, which may differ from their true preference ranks.
+These considerations motivate the study of rank-minimizing mechanisms, which determine assignments to minimize the total reported ranks.
+A well-known example is the uniform rank-minimizing mechanism (hereafter, URM), which selects uniformly at random from the set of deterministic assignments that minimize the total rank.
+The URM satisfies equal treatment of equals (Bogomolnaia and Moulin, 2001), a fundamental fairness notion requiring agents who report identical preferences to receive identical (ex-ante) treatment.
+Mechanisms that are both rank-minimizing and satisfy this fairness condition are referred to as fair rank-minimizing mechanisms, with the URM being their most representative example.
+While it is known that rank-minimizing mechanisms are not strategy-proof (Featherstone, 2020), we mainly study whether truth-telling is a dominated strategy.
+We first consider the benchmark setting without any refusal options, where each agent must accept the assigned object.
+In this case, we show that truth-telling is not a dominated strategy.
+Our main analysis concerns the case in which agents can refuse their assignment and instead receive an outside option.
 [^1] Feigenbaum et al.
@@
 For example, about 10% of the students withdrew from the New York City public high school placement.
-Featherstone (2020) theoretically considers the situation, but focuses only on the case where each agent has limited information on both the preferences of other agents and the capacities of all object types except the outside option.
-In contrast, we assume that each agent has complete information on the capacities of all object types.
-For example, we consider a school choice market where information about all schools, including their capacities, is disclosed to all students.
-We do not need any assumptions about the information structure of agents regarding the preferences and rationality of other agents, because we focus only on strategic dominance relations.
-We consider an outside option demotion strategy, which involves revealing a ranking where the rank of the outside option is placed at the bottom, without changing the order of the truly acceptable object types.
-We show that, if the uniform rank-minimizing mechanism is employed and agents have the refusal option, then any outside option demotion strategy weakly strategically dominates truth-telling.
-Moreover, with some additional conditions, some outside option demotion strategy (strictly) strategically dominates truth-telling.
-The reason is as follows: First, for some preference profiles of other agents, an agent can successfully avoid the outside option and obtain a truly acceptable object type because assigning them the outside option would significantly increase the sum of ranks.
-However, this manipulation carries the risk of being assigned an unacceptable object type.
-Yet, when agents have the refusal option, they can simply reject such assignments and choose the outside option instead.
-Thus, in this case, the risk associated with adopting this strategy is entirely eliminated.
-Moreover, we show that if an agent adopts an outside option demotion strategy and the uniform rank-minimizing mechanism is employed, an inefficient assignment may be realized.
-This is because, in that case, an unacceptable type may be assigned to the agent who adopts the strategy, and then the agent refuses it, even though the type is desired by other agents.
-Therefore, in this case, strategic manipulation is a critical problem for rank-minimizing mechanisms.
-Featherstone (2020) shows that revealing the full extension, which is one of the outside option demotion strategies, is a weakly better strategy than any other strategies in the case where the agents have little information and the refusal option.
-The implication of Featherstone (2020) is that rank-efficient mechanisms have strategic issues if the agents have little information and the outside option is somewhat valued by some agents.
-On the other hand, by our result, this implication is robust for the uniform rank-minimizing mechanism even without the low information assumption.
-Moreover, we show that revealing the full extension may be strategically dominated by another outside option demotion strategy because, contrary to our model, Featherstone (2020) assumes that the agents have no information on the capacities of object types other than their outside option.
-Due to the negative result on the uniform rank-minimizing mechanism introduced above, we modify it so that truth-telling is not strategically dominated by any other strategy.
-We call this mechanism the modified uniform rank-minimizing mechanism, which is shown to be fair.
-That is, to ensure that any outside option demotion strategy can fail in some case, in the mechanism, increasing acceptable object types can make the agent less likely to be assigned their acceptable types.
-However, using the modified uniform rank-minimizing mechanism may create another strategic issue: an agent may have an incentive to reduce the number of truly acceptable object types.
+In this setting, we show that a class of strategies called outside option demotion strategies, which place the outside option at the bottom of the reported ranking while preserving the order over acceptable objects, can strictly dominate truth-telling under the URM.
+The intuition is as follows: for some preference profiles of other agents, an agent who adopts such a strategy may succeed in avoiding the outside option and instead receive a truly acceptable object, because assigning the outside option to that agent would substantially increase the total rank.
+In the standard setting, this manipulation would be risky, as the agent might be assigned an unacceptable object.
+However, with the refusal option in place, the agent can simply reject any such assignment and receive the outside option instead.
+Thus, the strategies become effectively risk-free.
+We further show that such strategies may lead to inefficient outcomes: some objects that are desired by agents may be left unassigned if misreporting agents refuse them.
+In addition, we provide a concrete example in which an outside option demotion strategy forms part of a pure-strategy Nash equilibrium, reinforcing the strategic vulnerability of the URM with the refusal option.
+Finally, we propose a modified version of the URM that prevents truth-telling from being strictly dominated and is fair.
+However, this mechanism introduces new incentives: agents may benefit by promoting the outside option.
+Our results thus reveal a fundamental trade-off in the design of fair rank-minimizing assignment mechanisms in the presence of refusal options.
+Related Literature
+There is an extensive literature on fair assignment in probabilistic assignment problems.
+Prominent studies include Bogomolnaia and Moulin (2001), Budish et al.
+(2013), and Nesterov (2017), who explore fairness notions such as weak envy-freeness and the equal division lower bound.
+However, Feizi (2024) shows that these fairness notions are incompatible with rank-minimization.
+In contrast, ETE, which we adopt in this study, is compatible with rank-minimizing mechanisms and thus suitable for our analysis.
+The efficiency of rank-minimizing mechanisms is recently also well studied.
+Featherstone (2020) and Feizi (2024) show that rank-minimizing mechanisms satisfy stronger efficiency properties than commonly used ones, including ordinal efficiency, originally introduced by Bogomolnaia and Moulin (2001).
+[^2] Although Featherstone (2020) introduce the concept of rank-minimizing assignments in the economic literature, a broader body of work exists in mathematics and operations research.
+See Krokhmal and Pardalos (2009) for a comprehensive survey.
+Ortega and Klein (2023) compare outcomes under rank-minimizing mechanisms with those from the deferred acceptance and top trading cycles mechanisms, showing that the former improve on the latter two in several dimensions of efficiency.
+Regarding strategic behavior, much of the existing literature focuses on settings without the refusal option.
+In this context, Nikzad (2022) establishes Bayesian incentive compatibility for the URM in uniform markets, and Troyan (2024) shows that the URM is not obviously manipulable, in the sense of Troyan and Morrill (2020).
+Our result in the no-refusal setting is related but logically independent: we show that under any fair rank-minimizing mechanism, truth-telling is not a dominated strategy.
+On the other hand, several studies have pointed out strategic issues associated with rank-minimizing mechanisms even without the refusal option.
+Tasnim et al.
+(2024) propose strategies for improving individual profit under rank-minimizing mechanisms and show, using both simulated markets and the real matching market in Amsterdam, that applying these strategies can potentially reduce overall matching performance.
+Bando et al.
+(2025) also introduce a computationally efficient algorithm for finding profitable manipulations when agents are pessimistic.
+Our study complements them by theoretically identifying strategic dominance relations in settings where agents can refuse their assignments.
+Featherstone (2020) considers settings where agents can refuse their assignment.
+His analysis assumes that agents possess limited information—specifically, they are unaware of other agents’ preferences and of the capacities of all objects except their own outside option.
+Our model departs from this assumption: we suppose that agents know the capacities of all objects, as may be the case in a school choice context where this information is publicly available.
+Moreover, we abstract from assumptions about agents’ beliefs or rationality regarding others, focusing mainly on strategic dominance.
+Featherstone (2020) shows that when information is limited as above and agents have the refusal option, revealing the full extension, which is an instance of an ODS, is weakly better than any alternative.
+His results imply that rank-minimizing mechanisms are vulnerable to manipulation when outside options are valued to some extent.
+Our findings show that this vulnerability persists even in the absence of information limitation.
+Moreover, we show that the full extension may itself be strictly dominated by another ODS when agents are informed about object capacities.
+This highlights a new dimension of strategic risk that does not arise under Featherstone’s setting.
 2 Model
@@
 Let $\mathcal{X}$ be the set of all possible assignments.
-Specifically, $y\in\mathcal{X}$ is said to be a deterministic assignment if $y_{ao}$ is either 00 or $1$ for all $\left(a,o\right)\in\mathcal{A}\times\mathcal{O}$.
+Specifically, $y\in\mathcal{X}$ is said to be a deterministic assignment if $y_{ao}$ is either 0 or $1$ for all $\left(a,o\right)\in\mathcal{A}\times\mathcal{O}$.
 Let $\mathcal{Y}\subseteq\mathcal{X}$ be a set of all possible deterministic assignments.
@@
 for all $k=1,2,\cdots,\left|\mathcal{O}\right|$, then agent $a$ with $p_{a}$ is said to weakly prefer (in the sense of first-order stochastic dominance) $x$ to $x^{\prime}$.
-Moreover, $a$ is said to strictly prefer $x$ to $x^{\prime}$ if $a$ weakly prefers $x$ to $x^{\prime}$ and (1) is satisfied with inequality for some $k=1,2,\cdots,\left|\mathcal{O}\right|-1$.
+Moreover, $a$ is said to strictly prefer $x$ to $x^{\prime}$ if $a$ weakly prefers $x$ to $x^{\prime}$ and (1) is satisfied with strict inequality for some $k=1,2,\cdots,\left|\mathcal{O}\right|-1$.
 Assignment $x$ is wasteful for $p$ if there are $o,o^{\prime}\in\mathcal{O}$ and $a\in\mathcal{A}$ such that $\sum\nolimits_{a^{\prime}\in\mathcal{A}}x_{a^{\prime}o}<q_{o}$, $x_{ao^{\prime}}>0$ and $R\left(o,p_{a}\right)<R\left(o^{\prime},p_{a}\right)$.
@@
 If $x^{\ast}\in\mathcal{X}$ is rank-minimizing for $p$, then $x$ is not wasteful for $p$.
-The efficiency property of rank-minimizing is stronger than several other efficiency properties, such as rank efficiency and ordinal efficiency, analyzed in this literature (see Figure 1 in Feizi (2024)).
-Among them, not being wasteful is a very weak property.
-In the subsequent section, we focus on the mechanisms whose input is $p\in\mathcal{P}^{\left|\mathcal{A}\right|}$ and output is rank-minimizing for any $p\in\mathcal{P}^{\left|\mathcal{A}\right|}$.
+In the subsequent section, we focus on the mechanisms whose input is $p\in\mathcal{P}^{\left|\mathcal{A}\right|}$ and output is a rank-minimizing assignment for any $p\in\mathcal{P}^{\left|\mathcal{A}\right|}$.
 Before that we introduce some technical results.
@@
 Finally, we introduce another technical result that is used to show our main results introduced later. Let
-$\bar{k}\left(p_{a}\right)=\min\left\{k\in\left\{1,\cdots,\left|\mathcal{O}\right|\right\}\text{ }\left|\text{ }\sum\limits_{\bar{o}:R(\bar{o},p_{a})\leq k}q_{\bar{o}}\geq\left|\mathcal{A}\right|\right.\right\},$
+[…]
 which is well-defined because $\sum\nolimits_{\bar{o}\in\mathcal{O}}q_{\bar{o}}>\left|\mathcal{A}\right|$.
@@
 Lemma 2
-If $x\in\mathcal{X}$ be an assignment that is not wasteful for $p$, then $x_{ao}=0$ for all $o$ such that $R(o,p_{a})>\bar{k}\left(p_{a}\right)$.
+If an assignment $x\in\mathcal{X}$ is not wasteful for $p$, then $x_{ao}=0$ for all $o$ such that $R(o,p_{a})>\bar{k}\left(p_{a}\right)$.
 This means that if an assignment is not wasteful for $p$, then $a$ must be assigned to an object type that ranked higher than or equals to $\bar{k}\left(p_{a}\right)$.
-Typically, under a nonwasteful assignment, the object types that ranked lower than $\varnothing$ in the preference ranking of agent $a$ are never assigned to $a$, because $\varnothing$ is the object with a sufficient number of copies.
+Typically, under a nonwasteful assignment, the unacceptable object types in the preference ranking of agent $a$ are never assigned to $a$, because $\varnothing$ is the object with a sufficient number of copies.
 3 Rank-minimizing Mechanism
 In this section, we consider the following game.
-First, all agents simultaneously reveal a preference order in $\mathcal{P}$, which may or may not be their true preference order $r_{a}\in\mathcal{P}$.
+All agents simultaneously reveal a preference order in $\mathcal{P}$, which may or may not be their true preference order $r_{a}\in\mathcal{P}$.
 Let $p_{a}\in\mathcal{P}$ be the preference order revealed by $a\in\mathcal{A}$ and $p=\left(p_{a}\right)_{a\in\mathcal{A}}\in\mathcal{P}^{\left|\mathcal{A}\right|}$.
-Second, based on $p$, an assignment is determined via a mechanism.
+Then, based on $p$, an assignment is determined via a mechanism.
 We assume that all agents know the capacities of all object types $\left(q_{o}\right)_{o\in\mathcal{O}}$ and the mechanism is employed to determine the assignment, when they reveal a preference order.
 Note that in this section, the null object type is just that with sufficient number of copies.
-A mechanism employed in the second stage is represented by $f:\mathcal{P}^{\left|\mathcal{A}\right|}\rightarrow\mathcal{X}$.
+A mechanism employed in this game is represented by $f:\mathcal{P}^{\left|\mathcal{A}\right|}\rightarrow\mathcal{X}$.
 First, we consider an efficiency property on mechanisms.
 We call $f$ a rank-minimizing mechanism if $f(p)$ is rank-minimizing for all $p$.
-Second, we introduce a fairness property.
+In this study, we mainly focus on the following fair mechanism.
+Let $\mathcal{Y}^{\ast}\left(p\right)\subseteq\mathcal{Y}$ be the set of deterministic rank-minimizing assignments; that is, $y^{\ast}\in\mathcal{Y}^{\ast}\left(p\right),$ $RV\left(y^{\ast},p\right)\leq RV\left(x,p\right)$ for any $x\in\mathcal{X}$.
+By Corollary 1, $\mathcal{Y}^{\ast}\left(p\right)\neq\emptyset$.
+Let
+$f^{U}\left(p\right)=\sum\limits_{y^{\ast}\in\mathcal{Y}^{\ast}\left(p\right)}\frac{y^{\ast}}{\left|\mathcal{Y}^{\ast}\left(p\right)\right|}\text{ for all }p\in\mathcal{P}^{\left|\mathcal{A}\right|}$
+be the uniform rank-minimizing mechanism (URM). By (16),
+$RV\left(f^{U}\left(p\right),p\right)=\sum\limits_{y^{\ast}\in\mathcal{Y}^{\ast}\left(p\right)}\frac{1}{\left|\mathcal{Y}^{\ast}\left(p\right)\right|}RV\left(y^{\ast},p\right).$
+Thus, $RV\left(f^{U}\left(p\right),p\right)\leq RV\left(x,p\right)$ for any $x\in\mathcal{X}$; that is, $f^{U}\left(p\right)$ must be rank-minimizing for $p$.
+Since Nikzad (2022), Ortega and Klein (2023) and Troyan (2024) also discuss the URM, this is one of the most common rank-minimizing mechanisms in the literature.
+We show that the URM satisfies fundamental fairness properties.
 First, we introduce a well-known one.
@@
 This property is called just “equal treatment of equals” by many previous studies such as Bogomolnaia and Moulin (2001) and Troyan (2024).
-In this study, we mainly use a slightly stronger one.
-A mechanism $f$ is said to satisfy equal treatment of equals (hereafter ETE) if $p\in\mathcal{P}^{\left|\mathcal{A}\right|}$ satisfying $p_{a}^{k^{\prime}}=p_{a^{\prime}}^{k^{\prime}}$ for all $1,\cdots,\bar{k}\left(p_{a}\right)$ implies $\left(f\left(p\right)\right)_{ao}=\left(f\left(p\right)\right)_{a^{\prime}o}$ for all $o\in\mathcal{O}$ [^3] Han (2024) also considers a similar property, called symmetricity, which is slightly stronger than equal treatment of equals.
+We can show that the URM satisfies a slightly stronger version.
+A mechanism $f$ is said to satisfy equal treatment of equals (hereafter ETE) if $p\in\mathcal{P}^{\left|\mathcal{A}\right|}$ satisfying $p_{a}^{k^{\prime}}=p_{a^{\prime}}^{k^{\prime}}$ for all $1,\cdots,\bar{k}\left(p_{a}\right)$ implies $\left(f\left(p\right)\right)_{ao}=\left(f\left(p\right)\right)_{a^{\prime}o}$ for all $o\in\mathcal{O}$.
+[^3] Han (2024) also considers a similar property, called symmetricity, which is slightly stronger than equal treatment of equals.
 Unlike in our model, they take into account the priorities of objects and define agents as equal only if they are equal with respect to the priorities of all objects.
 By Lemma 2, as long as the assignment is not wasteful, any type that is ranked lower than $\bar{k}\left(p_{a}\right)$th in the preference ranking revealed by $a$ will not be assigned to $a$, because there are sufficient copies of types ranked higher than or equal to $\bar{k}\left(p_{a}\right)$th.
-Thus, if two individuals reveal rankings that are identical for positions at or above $\bar{k}\left(p_{a}\right)$th rank, those rankings are considered essentially the same, and they should receive the same assignment in an ex-ante sense.
-Other fairness properties discussed in the literature include weak envy-freeness and the equal division lower bound.
-However, as shown by Feizi (2024), both of these properties are incompatible with rank-minimizing.
-Since ETE is weaker than these fairness properties, it is compatible with rank-minimizing and is therefore adopted in this study.
-We focus on the efficient and fair mechanisms; that is, we call $f$ a fair rank-minimizing mechanism if $f(p)$ is rank-minimizing for all $p$ and satisfies ETE. The following mechanism is one of them.
-Let $\mathcal{Y}^{\ast}\left(p\right)\subseteq\mathcal{Y}$ be the set of deterministic rank-minimizing assignments; that is, $y^{\ast}\in\mathcal{Y}^{\ast}\left(p\right),$ $RV\left(y^{\ast},p\right)\leq RV\left(x,p\right)$ for any $x\in\mathcal{X}$.
-By Corollary 1, $\mathcal{Y}^{\ast}\left(p\right)\neq\emptyset$.
-Let
-$f^{U}\left(p\right)=\sum\limits_{y^{\ast}\in\mathcal{Y}^{\ast}\left(p\right)}\frac{y^{\ast}}{\left|\mathcal{Y}^{\ast}\left(p\right)\right|}\text{ for all }p\in\mathcal{P}^{\left|\mathcal{A}\right|}$
-be the uniform rank-minimizing mechanism. By (7),
-$RV\left(f^{U}\left(p\right),p\right)=\sum\limits_{y^{\ast}\in\mathcal{Y}^{\ast}\left(p\right)}\frac{1}{\left|\mathcal{Y}^{\ast}\left(p\right)\right|}RV\left(y^{\ast},p\right).$
-Thus, $RV\left(f^{U}\left(p\right),p\right)\leq RV\left(x,p\right)$ for any $x\in\mathcal{X}$; that is, $f^{U}\left(p\right)$ must be rank-minimizing for $p$.
-Since Nikzad (2022), Ortega and Klein (2023) and Troyan (2024) also discuss the uniform rank-minimizing mechanism, this is one of the most common rank-minimizing mechanisms in the literature.
-We have the following result.
+Thus, if two individuals reveal rankings that are identical for positions at or above $\bar{k}\left(p_{a}\right)$th rank, those rankings are considered essentially identical, and they should receive the same assignment in an ex-ante sense.
+We say that a mechanism satisfying ETE is a fair mechanism. We have the following result.
 Proposition 1
-The uniform rank-minimizing mechanism satisfies ETE.
-Therefore, the uniform rank-minimizing mechanism is a fair rank-minimizing one.
+The URM satisfies ETE.
+Therefore, the URM is a fair rank-minimizing one.
 We discuss strategic problems of fair rank-minimizing mechanisms.
@@
 Therefore, in this paper, we consider a weaker strategic property.
-For a mechanism $f$ and an agent $a$, (revealing) $p_{a}^{\prime}$ is said to be weakly strategically dominated by (revealing) $p_{a}$ if $a$ weakly prefers $f\left(p_{a},p_{-a}\right)$ to $f\left(p_{a}^{\prime},p_{-a}\right)$ for all $p_{-a}\in\mathcal{P}^{\left|\mathcal{A}\right|-1}$.
-Moreover, for $f$ and $a$, (revealing) $p_{a}^{\prime}$ is said to be (strictly) strategically dominated by (revealing) $p_{a}$ if $p_{a}^{\prime}$ is weakly strategically dominated by $p_{a}$ and there exists some $p_{-a}^{\prime}\in\mathcal{P}^{\left|\mathcal{A}\right|-1}$ such that $a$ strictly prefers $f\left(p_{a},p_{-a}^{\prime}\right)$ to $f\left(p_{a}^{\prime},p_{-a}^{\prime}\right)$.
-We discuss whether truth-telling ($r_{a}$) can be strategically dominated by revealing an untrue preference $p_{a}\neq r_{a}$.
+For a mechanism $f$ and an agent $a$, (revealing) $p_{a}^{\prime}\in\mathcal{P}$ is said to be weakly (strategically) dominated by (revealing) $p_{a}\in\mathcal{P}$ if $a$ weakly prefers $f\left(p_{a},p_{-a}\right)$ to $f\left(p_{a}^{\prime},p_{-a}\right)$ for all $p_{-a}\in\mathcal{P}^{\left|\mathcal{A}\right|-1}$.
+For $f$ and $a$, (revealing) $p_{a}^{\prime}$ is said to be (strictly) dominated by (revealing) $p_{a}$ if $p_{a}^{\prime}$ is weakly dominated by $p_{a}$ and there exists some $p_{-a}^{\prime}\in\mathcal{P}^{\left|\mathcal{A}\right|-1}$ such that $a$ strictly prefers $f\left(p_{a},p_{-a}^{\prime}\right)$ to $f\left(p_{a}^{\prime},p_{-a}^{\prime}\right)$.
+Moreover, for $f$ and $a$, (revealing) $p_{a}$ is a (weakly) dominated strategy if there is some $p_{a}^{\prime}\in\mathcal{P}$ that (weakly) dominates $p_{a}$.
+Moreover, we say that for $f$, $p\in\mathcal{P}^{\left|\mathcal{A}\right|}$ is a Nash equilibrium if there is no $a$ who strictly prefers $f\left(p_{a}^{\prime},p_{-a}\right)$ to $f\left(p_{a},p_{-a}\right)$ for some $p_{a}^{\prime}\in\mathcal{P}$.
+Therefore, in this context, we define a Nash equilibrium as a profile of strategies such that, for each agent, the resulting assignment is not first-order stochastically dominated by the assignment that results if the agent unilaterally deviated to any other strategy.
+We mainly discuss whether truth-telling ($r_{a}$) is a dominated strategy.
 Fix any $r_{a}$ and any $p_{a}\neq r_{a}$.
@@
 is satisfied for that $k$.
-In other word, we consider the case where the number of object types that are more preferable than the $k-1$th best one in $p_{a}$ is insufficient to distribute to all agents.
+In other word, we consider the case where the number of object types that are more preferable than the k1k-1th best one in $p_{a}$ is insufficient to distribute to all agents.
 In this case, for some $p_{-a}\in\mathcal{P}^{\left|\mathcal{A}\right|-1}$ and some fair rank-minimizing mechanism, $a$ is assigned to an object type that is less preferable or equal to the $k$th-best one.
@@
 Proposition 2
-Suppose that a fair rank-minimizing mechanism is employed. For any $r_{a}\in\mathcal{P}$, there is no $p_{a}\in\mathcal{P}$ such that revealing $p_{a}$ strategically dominates truth-telling.
+Suppose that $f$ is a fair rank-minimizing mechanism. For $f$ and $a$ with any $r_{a}\in\mathcal{P}$, truth-telling is not a dominated strategy.
 This result implies that any fair rank-minimizing mechanism satisfies a weak strategic characteristic; that is, for each untrue preference $p_{a}$, either $\left(f\left(r_{a},p_{-a}\right)\right)_{ao}=\left(f\left(p_{a},p_{-a}\right)\right)_{ao}$ for all $o\in\mathcal{O}$ and all $p_{-a}\in\mathcal{P}^{\left|\mathcal{A}\right|-1}$ or there must exist $p_{-a}$ such that $a$ is worse off by revealing it instead of $r_{a}$.
@@
 $\left(f^{\ast}\left(p_{a_{1}},p_{a_{2}},p_{a_{3}}\right)\right)_{a_{i}o_{1}}$ $=$ $0,$ (2)
-$f^{\ast}\left(p_{a_{1}},p_{a_{2}},p_{a_{3}}\right)$ $=$ $\sum\limits_{y^{\ast}\in\mathcal{Y}^{\ast}\left(p\right)}\frac{y^{\ast}}{\left|\mathcal{Y}^{\ast}\left(p\right)\right|}$
-otherwise; that is, if $\left(p_{a_{1}},p_{a_{2}},p_{a_{3}}\right)$ is neither $\left(\alpha^{\prime},\alpha,\alpha\right)$ nor $\left(\alpha,\alpha^{\prime},\alpha\right)$ nor $\left(\alpha,\alpha,\alpha^{\prime}\right)$.
-In words, if the first-best type and the second-best type of the three agents are respectively $o_{1}$ and $o_{2}$ and the third-best type of the two of them is $\varnothing$, then the agent revealing $o_{3}$ as the third-best one obtains $o_{1}$ with probability 00.
+$\left(f^{\ast}\left(p_{a_{1}},p_{a_{2}},p_{a_{3}}\right)\right)_{ao}$ $=$ $\sum\limits_{y^{\ast}\in\mathcal{Y}^{\ast}\left(p\right)}\frac{y^{\ast}}{\left|\mathcal{Y}^{\ast}\left(p\right)\right|}$
+for all $a\in\mathcal{A}$ and $o\in\mathcal{O}$ otherwise; that is, if $\left(p_{a_{1}},p_{a_{2}},p_{a_{3}}\right)$ is neither $\left(\alpha^{\prime},\alpha,\alpha\right)$ nor $\left(\alpha,\alpha^{\prime},\alpha\right)$ nor $\left(\alpha,\alpha,\alpha^{\prime}\right)$.
+In words, if the first-best type and the second-best type of the three agents are respectively $o_{1}$ and $o_{2}$ and the third-best type of the two of them is $\varnothing$, then the agent revealing $o_{3}$ as the third-best one obtains $o_{1}$ with probability 0.
 We additionally assume that if $\left(p_{a_{i}},p_{a_{j}},p_{a_{k}}\right)=\left(\alpha^{\prime},\alpha,\alpha\right)$, then
@@
 In this case, $\bar{k}\left(\alpha\right)=\bar{k}\left(\alpha^{\prime}\right)=3$.
-Then, for example, $r_{a_{1}}=\alpha^{\prime}$ is not strategically dominated by $\alpha$ and any other preference orders, because
+Then, for example, $r_{a_{1}}=\alpha^{\prime}$ is not dominated by $\alpha$ and any other preference orders, because
 $\left(f^{\ast}\left(\alpha^{\prime},\alpha^{\prime},\alpha^{\prime}\right)\right)_{a_{1}o_{1}}$ $=$ $\left(f^{\ast}\left(\alpha^{\prime},\alpha^{\prime},\alpha^{\prime}\right)\right)_{a_{1}o_{2}}=\left(f^{\ast}\left(\alpha^{\prime},\alpha^{\prime},\alpha^{\prime}\right)\right)_{a_{1}o_{3}}=\frac{1}{3},$
@@
 Thus, the only strategic interest of $a_{1}$ is increasing the probability to obtain $o_{1}$.
-If $f^{\ast}$ is employed, then $\alpha^{\prime}=r_{a_{1}}$ is strategically dominated by $\alpha$ for $a_{1}$ due to (2).
-Therefore, we conclude that even if a rank-minimizing mechanism that satisfies weak ETE is employed, truth-telling may be strategically dominated by some other strategy.
+If $f^{\ast}$ is employed, then $\alpha^{\prime}=r_{a_{1}}$ is dominated by $\alpha$ for $a_{1}$ due to (2).
+Therefore, we conclude that even if a rank-minimizing mechanism that satisfies weak ETE is employed, truth-telling may be dominated by some other strategy.
 Note that in the same case, $f^{\ast}$ does not satisfy ETE.
-First, since $q_{o_{1}}+q_{o_{2}}\geq\left|\mathcal{A}\right|$, any agent is assigned to neither $o_{3}$ nor $\varnothing$,∅ , as long as the assignment is not wasteful.
+First, since $q_{o_{1}}+q_{o_{2}}\geq\left|\mathcal{A}\right|$, any agent is assigned to neither $o_{3}$ nor $\varnothing,$ as long as the assignment is not wasteful.
 Thus, $\alpha$ and $\alpha^{\prime}$ are essentially the same in this case.
@@
 4 Refusal option
-In this section, we consider the following three-stage game, which is a modification of that in Section 3; that is, the first and second stages in this game are the same as in the previous game.
-In the third stage, all agents decide whether to accept or refuse the assigned object type.
+In this section, we consider the following two-stage game, which is a modification of that in Section 3.
+In the first stage, all agents simultaneously reveal a preference order in $\mathcal{P}$ and an assignment is determined via a mechanism based on $p$.
+In the second stage, all agents decide whether to accept the assignment determined in the first stage.
 If an agent refuses their assignment, then the agent instead obtains the null object type $\varnothing$ (the outside option).
 That is, in this game, the null object type has two characteristics: first, there are sufficiently many copies of this type to allocate to all individuals, and second, agents can always obtain it at no cost as long as they abandon their assignment.
-This third stage is also considered by Featherstone (2020) but is not explicitly addressed by almost other previous studies related to ours.
+This two-stage game is also considered by Featherstone (2020) but is not explicitly addressed by almost other previous studies related to ours.
 [^4] As exceptions, Feigenbaum et al.
@@
 Moreover, Doğan and Yenmez (2019) also explicitly consider this stage.
-However, this stage is implicitly considered by many of them; that is, they consider an ex-post individually rational (or more strictly an ex-post efficient) matching as a requirement and if an assignment is ex-post individually rational, no agent refuses their assignment.
+However, this stage is implicitly considered by many of them; that is, they consider an ex-post individually rationality (or more strictly an ex-post efficiency) as a requirement and if an assignment is ex-post individually rational, no agent refuses their assignment.
 See, for example, Bogomolnaia and Moulin (2001) on the property.
@@
 $g\left(x,r\right)_{ao}$ $=$ $x_{ao}\text{ for all }R\left(o,r_{a}\right)<R\left(\varnothing,r_{a}\right)$
-$g\left(x,r\right)_{ao}$ $=$ 0 for all $0\text{ for all }R\left(o,r_{a}\right)>R\left(\varnothing,r_{a}\right)$
+$g\left(x,r\right)_{ao}$ $=$ $0\text{ for all }R\left(o,r_{a}\right)>R\left(\varnothing,r_{a}\right)$
 $g\left(x,r\right)_{a\varnothing}$ $=$ $\sum\limits_{o:R\left(o,r_{a}\right)\geq R\left(\varnothing,r_{a}\right)}x_{ao}.$
@@
 We reconsider strategic problems of a fair rank-minimizing mechanism in this three-stage game.
-For a mechanism $f$ and an agent $a$, (revealing) $p_{a}^{\prime}$ is said to be weakly strategically dominated by (revealing) $p_{a}$ with the refusal option if $a$ weakly prefers $g\left(f\left(p_{a},p_{-a}\right),r\right)$ to $g\left(f\left(p_{a}^{\prime},p_{-a}\right),r\right)$ for all $p_{-a}\in\mathcal{P}^{\left|\mathcal{A}\right|-1}$.
-Moreover, for $f$ and $a$, (revealing) $p_{a}^{\prime}$ is said to be (strictly) strategically dominated by (revealing) $p_{a}$ with the refusal option if $p_{a}^{\prime}$ is weakly strategically dominated by $p_{a}$ with the refusal option and there exists some $p_{-a}^{\prime}\in\mathcal{P}^{\left|\mathcal{A}\right|-1}$ such that $a$ strictly prefers $g\left(f\left(p_{a},p_{-a}^{\prime}\right),r\right)$ to $g\left(f\left(p_{a}^{\prime},p_{-a}^{\prime}\right),r\right)$.
-We show that, under the uniform rank-minimizing mechanism, truth-telling is strategically dominated by another strategy when the agents have the refusal option.
+For a mechanism $f$ and an agent $a$, (revealing) $p_{a}^{\prime}$ is said to be weakly dominated by (revealing) $p_{a}$ with the refusal option if $a$ weakly prefers $g\left(f\left(p_{a},p_{-a}\right),r\right)$ to $g\left(f\left(p_{a}^{\prime},p_{-a}\right),r\right)$ for all $p_{-a}\in\mathcal{P}^{\left|\mathcal{A}\right|-1}$.
+Moreover, for $f$ and $a$, (revealing) $p_{a}^{\prime}$ is said to be (strictly) dominated by (revealing) $p_{a}$ with the refusal option if $p_{a}^{\prime}$ is weakly dominated by $p_{a}$ with the refusal option and there exists some $p_{-a}^{\prime}\in\mathcal{P}^{\left|\mathcal{A}\right|-1}$ such that $a$ strictly prefers $g\left(f\left(p_{a},p_{-a}^{\prime}\right),r\right)$ to $g\left(f\left(p_{a}^{\prime},p_{-a}^{\prime}\right),r\right)$.
+For a mechanism $f$ and an agent $a$, (revealing) $p_{a}\in\mathcal{P}$ is said to be a (weakly) dominated strategy with the refusal option if there is some $p_{a}^{\prime}\in\mathcal{P}$ that (weakly) dominates $p_{a}$ with the refusal option.
+Moreover, we say that $p\in\mathcal{P}^{\left|\mathcal{A}\right|}$ is a Nash equilibrium with the refusal option if there is no $a$ who strictly prefers $g\left(f\left(p_{a}^{\prime},p_{-a}\right),r\right)$ to $g\left(f\left(p_{a},p_{-a}\right),r\right)$ for some $p_{a}^{\prime}\in\mathcal{P}$.
+We show that, under the URM, truth-telling is dominated by another strategy when the agents have the refusal option.
 We define the following specific strategies.
@@
 Note that an ODS (revealing $d_{a}\in\mathcal{D}_{a}$) increases the number of acceptable object types revealed by $a$; that is, $R\left(\varnothing,r_{a}\right)\leq R\left(\varnothing,d_{a}\right)=\left|\mathcal{O}\right|$ is satisfied.
-Featherstone (2020) defines the full extension of $r_{a}$denoted by $e_{a}$ in this study such that (i) $R\left(o,r_{a}\right)=R\left(o,e_{a}\right)$ for all $o\in\mathcal{O}$ such that $R\left(o,r_{a}\right)<R\left(\varnothing,r_{a}\right)$, and (ii) $R\left(\varnothing,e_{a}\right)=\left|\mathcal{O}\right|$, and $R\left(o,r_{a}\right)-1=R\left(o,e_{a}\right)$ for all $o\in\mathcal{O}$ such that $R\left(o,r_{a}\right)>R\left(\varnothing,r_{a}\right)$; that is, $e_{a}\in\mathcal{D}_{a}$ is such that all non-null object types are placed in their true preference orders.
+If $R\left(\varnothing,r_{a}\right)=\left|\mathcal{O}\right|,$ then truth-telling is the unique ODS for agent $a$.
+Therefore, an ODS constitutes a strategic manipulation if and only if $R\left(\varnothing,r_{a}\right)<\left|\mathcal{O}\right|;$ that is, the null object is not the worst object type for $a$.
+Featherstone (2020) defines the full extension of $r_{a}\$denoted by $e_{a}$ in this study such that (i) $R\left(o,r_{a}\right)=R\left(o,e_{a}\right)$ for all $o\in\mathcal{O}$ such that $R\left(o,r_{a}\right)<R\left(\varnothing,r_{a}\right)$, and (ii) $R\left(\varnothing,e_{a}\right)=\left|\mathcal{O}\right|$, and $R\left(o,r_{a}\right)-1=R\left(o,e_{a}\right)$ for all $o\in\mathcal{O}$ such that $R\left(o,r_{a}\right)>R\left(\varnothing,r_{a}\right)$; that is, $e_{a}\in\mathcal{D}_{a}$ is such that all non-null object types are placed in their true preference orders.
 Revealing the full extension of $r_{a}$ is one of the ODSs of $a$.
@@
 Thus, revealing $d_{a_{1}}$ instead of $r_{a_{1}}$ may increase the probability that $a_{1}$ is assigned to their first-best object $o_{1}$.
-First, under $\left(r_{a_{1}},\beta,\beta^{\prime}\right),$varnothing is ranked lowest by $a_{2}$ and $a_{3}$ but $a_{1}$ ranks $\varnothing$ as their second-best.
+First, under $\left(r_{a_{1}},\beta,\beta^{\prime}\right),$ $\varnothing$ is ranked lowest by $a_{2}$ and $a_{3}$ but $a_{1}$ ranks $\varnothing$ as their second-best.
 Thus, in any rank-minimizing mechanism, $a_{1}$ is assigned $\varnothing$ with probability 1.
@@
 Third, suppose $p_{a_{2}}=p_{a_{3}}=\beta$. Then,
-$f^{U}\left(r_{a_{1}},\beta,\beta\right)=\left(\begin{array}[]{ccc}0&0&1\\ 1/2&1/2&0\\ 1/2&1/2&0\end{array}\right),$
-$f^{U}\left(d_{a_{1}},\beta,\beta\right)=\left(\begin{array}[]{ccc}1/3&1/3&1/3\\ 1/3&1/3&1/3\\ 1/3&1/3&1/3\end{array}\right).$
+$f^{U}\left(r_{a_{1}},\beta,\beta\right)=\left(\begin{array}[]{ccc}0&0&1\\ 1/2&1/2&0\\ 1/2&1/2&0\end{array}\right),$ (6)
+$f^{U}\left(d_{a_{1}},\beta,\beta\right)=\left(\begin{array}[]{ccc}1/3&1/3&1/3\\ 1/3&1/3&1/3\\ 1/3&1/3&1/3\end{array}\right).$ (10)
 In the latter case, $a_{1}$ refuses the assignment; that is,
-$g\left(f^{U}\left(d_{a_{1}},\beta,\beta\right),r\right)=\left(\begin{array}[]{ccc}1/3&0&2/3\\ 1/3&1/3&1/3\\ 1/3&1/3&1/3\end{array}\right).$
+$g\left(f^{U}\left(d_{a_{1}},\beta,\beta\right),r\right)=\left(\begin{array}[]{ccc}1/3&0&2/3\\ 1/3&1/3&1/3\\ 1/3&1/3&1/3\end{array}\right).$ (11)
 Then, $a_{1}$ is better off by adopting the ODS instead of truth-telling. However, $g\left(f^{U}\left(d_{a_{1}},\beta,\beta\right),r\right)$ is wasteful, because
 $\sum\limits_{a\in\left\{a_{1},a_{2},a_{3}\right\}}g\left(f^{U}\left(d_{a_{1}},\beta,\beta\right),r\right)_{ao_{2}}$ $=$ $\frac{2}{3}<1,$
-and $\text{and }g\left(f^{U}\left(d_{a_{1}},\beta,\beta\right),r\right)_{a^{\prime}\varnothing}$ $=$ $\frac{1}{3}>0,$
+$\text{and }g\left(f^{U}\left(d_{a_{1}},\beta,\beta\right),r\right)_{a^{\prime}\varnothing}$ $=$ $\frac{1}{3}>0,$
 for $a^{\prime}=a_{2},a_{3}$. That is, for
@@
 If $a_{1}$ is assigned to $o_{2}$, then either $a_{2}$ or $a_{3}$ will be assigned to $\varnothing$.
-However, in this case, $a_{1}$ will refuse $o_{2}$ and thus $o_{2}$ will be remained unallocated.
+However, in this case, $a_{1}$ will refuse $o_{2}$ and thus $o_{2}$ will remain unallocated.
 In this case, rather than leaving $o_{2}$ unallocated, it is more efficient to assign $o_{2}$ to the agent ($a_{2}$ or $a_{3}$) who would otherwise receive $\varnothing$ and instead assign $a_{1}$ to $\varnothing$ from the beginning.
@@
 Hence revealing $d_{a_{1}}$ and refusing an unacceptable assignment may benefit $a_{1}$ but lead to a wasteful allocation, as some agents may prefer the object type that $a_{1}$ refuses.
-Note that $\left(d_{a_{1}}\left(=\beta\right),\beta,\beta\right)$ can be a Nash equilibrium.
-Trivially, $a_{1}$ has no incentive to change $d_{a_{1}}$.
-On the other hand, suppose that both $a_{2}$ and $a_{3}$ prefer the assignment such that each object $o\in\left\{o_{1},o_{2},\varnothing\right\}$ assigns with the probability $1/3$ to the assignment such that $o_{2}$ assigns with probability $1$.
-This ensures that neither $a_{2}$ nor $a_{3}$ has an incentive to change $\beta$ and $\beta$ to $(o_{2},o_{1},\varnothing)$ or $(o_{2},\varnothing,o_{1})$.
-Then, in this case, $\left(d_{a_{1}},\beta,\beta\right)$ is a Nash equilibrium.
-Thus, a wasteful assignment can occur as a Nash equilibrium of agents’ strategies when the assignment is determined by the uniform rank-minimizing mechanism.
 Now, we formally introduce our main results.
 Theorem 1
-Suppose that the uniform rank-minimizing mechanism is employed. Then, for all agents, truth-telling is weakly strategically dominated by any ODS with the refusal option.
+For $f^{U}$ and any agent $a$, truth-telling is weakly dominated strategy; that is, it is weakly dominated by any ODS with the refusal option.
 By this result, an agent is never hurt by using an ODS instead of truth-telling.
@@
 Theorems 1 and 2 do not imply that an ODS is a dominant strategy.
-However, if there is $a\in\mathcal{A}$ whose true preference satisfies (3) for some $o,o^{\prime}\in\mathcal{O}$, since truth-telling is strategically dominated by another strategy, $a$ has a strong incentive to manipulate its preference ranking, while one of the ODSs may not necessarily be chosen.
+However, if there is $a\in\mathcal{A}$ whose true preference satisfies (12) for some $o,o^{\prime}\in\mathcal{O}$, since truth-telling is a dominated strategy, $a$ has a strong incentive to manipulate its preference ranking.
 As stated earlier, multiple ODSs may exist.
-Even if an ODS strategically dominates truth-telling, another ODS that does not strategically dominate truth-telling may also exist.
-In such a case, the latter ODS is strategically dominated by the former.
+Even if an ODS dominates truth-telling, another ODS that does not dominate truth-telling may also exist.
+In such a case, the latter ODS is dominated by the former.
 The strategy of revealing the full extension is not an exception to this.
@@
 $g\left(f^{U}\left(d_{a_{1}},\gamma^{\prime},\gamma^{\prime\prime}\right),r\right)=\left(\begin{array}[]{cccc}1/2&0&0&1/2\\ 1/2&1/2&0&0\\ 0&0&1&0\end{array}\right)\text{.}$
-Therefore, $a_{1}$ (strictly) prefers $g\left(f^{U}\left(d_{a_{1}},\gamma^{\prime},\gamma^{\prime\prime}\right),r\right)$ to $g\left(f^{U}\left(r_{a_{1}},\gamma^{\prime},\gamma^{\prime\prime}\right),r\right)$.
+Therefore, $a_{1}$ (strictly) prefers $g\left(f^{U}\left(d_{a_{1}},\gamma^{\prime},\gamma^{\prime\prime}\right),r\right)$ to $g\left(f^{U}\left(r_{a_{1}},\gamma^{\prime},\gamma^{\prime\prime}\right),r\right)$ and $r_{a_{1}}$ is dominated by $d_{a_{1}}$.
 On the other hand, there is no pair $\left(p_{a_{2}},p_{a_{3}}\right)$ of preference orders of $a_{2}$ and $a_{3}$ such that $a_{1}$ (strictly) prefers $g\left(f^{U}\left(e_{a_{1}},p_{a_{2}},p_{a_{3}}\right),r\right)$ to $g\left(f^{U}\left(r_{a_{1}},p_{a_{2}},p_{a_{3}}\right),r\right)$.
@@
 $\left(g\left(f^{U}\left(e_{a_{1}},p_{a_{2}},p_{a_{3}}\right),r\right)\right)_{a_{1}\bar{o}}=\left(g\left(f^{U}\left(r_{a_{1}},p_{a_{2}},p_{a_{3}}\right),r\right)\right)_{a_{1}\bar{o}}\text{.}$
-We explain why $d_{a_{1}}$ strategically dominates $e_{a_{1}}$.
+Thus, $e_{a_{1}}$ is also dominated by $d_{a_{1}}$.
+We intuitively explain why $d_{a_{1}}$ dominates $e_{a_{1}}$.
 In this example, $q_{o_{1}}+q_{o_{3}}=2<3=q_{o_{1}}+q_{o_{2}}$.
@@
 However, since $a_{1}$ ranks $o_{2}$ and $\varnothing$ lower, minimizing the total rank requires increasing the probability that $o_{1}$ is allocated to $a_{1}$.
-Thus, revealing the full extension may be strategically dominated by another ODS when an unacceptable type (in this case $o_{3}$) is less preferable than the other (in this case $o_{2}$) but the number of copies of the former is less than that of the latter.
+Thus, revealing the full extension may be dominated by another ODS when an unacceptable type (in this case $o_{3}$) is less preferable than the other (in this case $o_{2}$) but the number of copies of the former is less than that of the latter.
 Note that this fact does not contradict a result of Featherstone (2020, Proposition 13).
 In their model, they assume that the agents have no information on the capacities of object types.
-Revealing the full extension can be strategically dominated by another ODS only when the capacities of the unacceptable types are different and agents know them.
-Thus, the strategic domination in Example 3 does not occur in the low-information situation discussed by Featherstone (2020).
-Next, we show that if some agent adopts an ODS, then the result may be wasteful and thus inefficient.
+Revealing the full extension can be dominated by another ODS only when the capacities of the unacceptable types are different and agents know them.
+Hence the strategic domination in Example 3 does not occur in the low-information situation discussed by Featherstone (2020).
+Next, we show that if an agent adopts an ODS, the resulting assignment may be wasteful and thus inefficient. Moreover, such an outcome can arise in a Nash equilibrium.
 Proposition 3
-Suppose that there is $a\in\mathcal{A}$ whose true preference satisfies (3) for some $o,o^{\prime}\in\mathcal{O}$.
-Then, there are some ODS of revealing $d_{a}$ and some $p_{-a}\in\mathcal{P}^{\left|\mathcal{A}\right|-1}$ such that $R\left(\varnothing,p_{a^{\prime}}\right)=\left|\mathcal{O}\right|$ for all $a^{\prime}\in\mathcal{A\setminus}\left\{a\right\}$ and $g\left(f^{U}\left(d_{a},p_{-a}\right),r\right)$ is wasteful.
-By Theorem 2, for an agent $a$ with $r_{a}$ that satisfy (3) for some $o,o^{\prime}\in\mathcal{O}$, an ODS dominates truth-telling.
-Moreover, by Proposition 3, in the same case, the strategy may make the assignment wasteful.
-In an ODS, an agent reveals actually unacceptable types as acceptable types.
-Then, the agent may be assigned to an actually unacceptable type and refuse it.
-Since the type may be desirable for another agent, in that case, the assignment is wasteful.
-By Theorems 1 and 2 and Proposition 3, if the uniform rank-minimizing mechanism is employed, the agents may have a strong incentive to manipulate their preferences, and these manipulations may result in an inefficient assignment.
-Thus, next, we consider another fair rank-minimizing mechanism such that, contrary to $f^{U}$, if it is employed as the mechanism in the second stage, then for any agent, truth-telling is not strategically dominated by any other strategy.
-We let $f^{M}$ be the modified uniform rank-minimizing mechanism such that if $p$ satisfies
+There exist $r\in\mathcal{P}^{\left|\mathcal{A}\right|}$ and $d\in\mathcal{P}^{\left|\mathcal{A}\right|}$ such that $d,$ the profile in which all agents adopt ODSs, is a Nash equilibrium with the refusal option under $f^{U}$.
+Moreover, if there are $o,o^{\prime}\in\mathcal{O}$ such that $q_{o}+q_{o^{\prime}}<\left|\mathcal{A}\right|$, then such profiles $r\in\mathcal{P}^{\left|\mathcal{A}\right|}$ and $d\in\mathcal{P}^{\left|\mathcal{A}\right|}$ can also yield a wasteful assignment; that is, $g\left(f^{U}\left(d\right),r\right)$ is wasteful.
+We briefly show this by revisiting Example 2. Let
+$r=\left(r_{a_{1}},r_{a_{2}},r_{a_{3}}\right)=\left(\beta^{\prime\prime},\beta,\beta\right)\text{ and }d=\left(d_{a_{1}},d_{a_{2}},d_{a_{3}}\right)=\left(\beta,\beta,\beta\right).$
+Then, $g\left(f^{U}\left(d\right),r\right)$ is given by (11), which is a wasteful assignment.
+Moreover, $d=\left(\beta,\beta,\beta\right)$ is a Nash equilibrium; that is, for each agent $a$, $g\left(f^{U}\left(d\right),r\right)$ is not first-order stochastically dominated by $g\left(f^{U}\left(p_{a},d_{-a}\right),r\right)$ for any $p_{a}\in\mathcal{P}$.
+[^5] Since we consider the first-order stochastic dominance relation, our notion of Nash equilibrium is weaker than the standard one.
+Nevertheless, suppose that both $a_{2}$ and $a_{3}$ prefer the assignment such that each object $o\in\left\{o_{1},o_{2},\varnothing\right\}$ assigns with the probability $1/3$ to the assignment such that $o_{2}$ assigns with probability $1$.
+This assumption ensures that neither $a_{2}$ nor $a_{3}$ has an incentive to change $\beta$ and $\beta$ to $(o_{2},o_{1},\varnothing)$ or $(o_{2},\varnothing,o_{1})$.
+Under this condition, $d=\left(\beta,\beta,\beta\right)$ is a Nash equilibrium even in the standard sense.
+By Proposition 3, if agents adopt an ODS, then the result may be wasteful.
+Moreover, such a strategy profile can be a Nash equilibrium.
+Therefore, under the URM, a profile of ODS are stably adopted in place of truth-telling and that such strategies may result in inefficient outcomes.
+This suggests a significant drawback of the URM.
+Therefore, we next consider an alternative fair rank-minimizing mechanism which, unlike $f^{U}$, guarantees that for any agent, truth-telling is not dominated by any other strategy when it is used as the mechanism in the second stage.
+We let $f^{M}$ be the modified uniform rank-minimizing mechanism (hereafter modified URM) such that if $p$ satisfies
 there is one agent represented by $a$ such that $R\left(o,p_{a}\right)=1$ and $R\left(\varnothing,p_{a}\right)\geq 3,$
@@
 $f^{M}\left(p\right)=\sum\limits_{y^{\ast}\in\mathcal{Y}^{\ast}\left(p\right)}\frac{y^{\ast}}{\left|\mathcal{Y}^{\ast}\left(p\right)\right|}=f^{U}\left(p\right).$
-To understand this mechanism, we reconsider Example 1.
-First, suppose $q_{o_{1}}=q_{o_{2}}=1$.
-If $\left(p_{a_{1}},p_{a_{2}},p_{a_{3}}\right)=\left(\alpha^{\prime},\alpha,\alpha\right)$, $\left(\alpha,\alpha^{\prime},\alpha\right)$ or $\left(\alpha,\alpha,\alpha^{\prime}\right)$, then $\left(p_{a_{1}},p_{a_{2}},p_{a_{3}}\right)$ satisfies (i), (ii) and (iii).
-In these cases, the student who chooses $\alpha^{\prime}$ is corresponding to $a,$ another student is corresponding to $a^{\prime}$, and $o_{1}$ is corresponding to $o$.
-The rank-minimizing mechanism $f^{\ast}$ introduced in Example 1 is the modified uniform rank-minimizing mechanism, because a student who chooses $\alpha^{\prime}$ obtains $o_{1}$ with probability 00 and another student obtains $o_{1}$ with probability $1$
-On the other hand, suppose $q_{o_{1}}=1$ and $q_{o_{2}}\geq 2$.
-Even if $\left(p_{a_{1}},p_{a_{2}},p_{a_{3}}\right)=\left(\alpha^{\prime},\alpha,\alpha\right)$, $\left(\alpha,\alpha^{\prime},\alpha\right)$ or $\left(\alpha,\alpha,\alpha^{\prime}\right)$, this does not satisfy (ii).
-This is because in this case, $p_{a^{\prime}}=\alpha$ and $l=3,$ but
-$\sum\limits_{\bar{o}:R\left(\bar{o},p_{a^{\prime}}\right)<R\left(\varnothing,p_{a^{\prime}}\right)}q_{\bar{o}}=q_{o_{1}}+q_{o_{2}}=3=\left|\mathcal{A}\right|.$
-Therefore, (ii) is not satisfied with $p_{a^{\prime}}=\alpha$ in the case of $q_{o_{1}}=1$ and $q_{o_{2}}\geq 2$.
+Under this mechanism, the outcome is equivalent to that of the URM, except when a specific preference profile satisfying conditions (i)–(iii) is realized.
+The specific profile characterized by conditions (i)–(iii) is one in which: (i) a particular agent denoted by $a$ ranks object $o$ first and the outside option at position three or lower; (ii) there are $q_{o}$ other agents whose preferences match $a$’s for the top l1l-1 positions, rank the outside option exactly $l$th, and for whom the total capacity of objects ranked above $l$ is insufficient for all agents; and (iii) if other agents remain, all of them rank the outside option first.
 We have the following result.
 Proposition 4
-The modified uniform rank-minimizing mechanism $f^{M}$ satisfies ETE.
-The strategic issues, mentioned in Theorems 1 and 2, do not arise under this mechanism $f^{M}$.
-First, we explain why the strategic issue mentioned in Example 2 does not arise in the modified uniform rank-minimizing mechanism.
-If $r_{a}=p_{a^{\prime}}$ where $p_{a^{\prime}}$ is defined in (ii), then $\left(d_{a},p_{-a}\right)$ satisfies (i), (ii) and (iii) where $d_{a}\in\mathcal{D}_{a}$.
-This implies that for some $p_{-a}$, if agent $a$ adopts an ODS instead of truth-telling and $f^{M}$ is employed, then $a$ is worse off.
-To see this fact, we revisit Example 2.
+The modified URM satisfies ETE.
+Then, we explain why the strategic issues, mentioned in Theorems 1 and 2, do not arise under this mechanism $f^{M}$.
+To ensure that truth-telling is not dominated by any ODS, the mechanism must be such that, for every agent and ODS, there exists a strategy profile of the other agents under which the agent is not assigned to any of the more preferred object types.
+Suppose that the true preference of $a$ satisfies $r_{a}^{1}=o$ and $r_{a}^{2}=\varnothing$.
+Given $p_{-a}$ satisfying conditions (ii) and (iii), condition (i) is not satisfied if agent $a$ chooses truth-telling, but it is satisfied if $a$ adopts an ODS.
+Therefore, when agent $a$ adopts an ODS, the probability of being assigned their most preferred object $o$ becomes zero, whereas under truth-telling, this probability is positive.
+To understand this point more, we revisit Example  2.
 In this example, $\left(d_{a_{1}},p_{a_{2}},p_{a_{3}}\right)=\left(\beta,\beta^{\prime\prime},\beta^{\prime\prime}\right)$ satisfies (i), (ii) and (iii), but $\left(r_{a_{1}},p_{a_{2}},p_{a_{3}}\right)=\left(\beta^{\prime\prime},\beta^{\prime\prime},\beta^{\prime\prime}\right)$ does not.
@@
 This implies that $a_{1}$ may be worse off by revealing $d_{a_{1}}$ instead of $r_{a_{1}}$, even if $a_{1}$ refuses any unacceptable objects.
-Therefore, $r_{a_{1}}$ is not strategically dominated by $d_{a_{1}}$, if $f^{M}$ is employed.
+Therefore, $r_{a_{1}}$ is not dominated by $d_{a_{1}}$, if $f^{M}$ is employed.
 Formally, we have the following fact.
 Proposition 5
-Suppose that the modified uniform rank-minimizing mechanism is employed.
-For any $r_{a}\in\mathcal{P}$, there is no $p_{a}\in\mathcal{P}$ that strategically dominates $r_{a}$ with (or without) the refusal option.
+For $f^{M}$ and $a$ with any $r_{a}\in\mathcal{P}$, truth-telling is not a dominated strategy with the refusal option.
 By adopting $f^{M}$, any ODS can fail because increasing the number of acceptable objects may lead the agent to be unassigned to truly acceptable objects, depending on the strategy profile of the other agents.
 Thus, in some profiles of other agents’ (revealed) preferences, truth-telling is better than adopting any ODS.
-However, $f^{M}$ may lead to other strategic problems.
-That is, if $f^{M}$ is adopted instead of $f^{U}$, an agent may have an incentive to reduce the number of acceptable object types.
+However, the modified URM may introduce alternative strategic concerns. We now turn to a different class of strategic manipulations.
+Let $\Pi_{a}\subseteq\mathcal{P}$ be a set of preferences such that $\pi_{a}\in\Pi_{a}$ if (i) $R\left(o,r_{a}\right)=R\left(o,\pi_{a}\right)$ for all $o\in\mathcal{O}$ such that $R\left(o,\pi_{a}\right)>R\left(\varnothing,\pi_{a}\right)$, (ii) $R\left(\varnothing,\pi_{a}\right)<R\left(\varnothing,r_{a}\right)$, and (iii) $R\left(o,r_{a}\right)=R\left(o,\pi_{a}\right)+1$ for all $o$ such that $R\left(\varnothing,\pi_{a}\right)\leq R\left(o,r_{a}\right)<R\left(\varnothing,r_{a}\right)$, and (iv) $R\left(o,r_{a}\right)=R\left(o,\pi_{a}\right)$ for all $o$ such that $R\left(\varnothing,r_{a}\right)<R\left(o,r_{a}\right)$.
+We call the strategy of agent $a$ of revealing a preference in $\Pi_{a}$ an outside option promotion strategy (hereafter OPS) of agent $a$.
+In words, contrary to ODSs, an agent raises the rank of the outside option in their reported preference, making fewer object types appear acceptable.
+First, we have the following result on this class of strategies on the URM (not modified one).
+Proposition 6
+For $f^{M}$ and $a$ with any $r_{a}\in\mathcal{P}$, any OPS of $a$ is weakly dominated by truth-telling of $a$. Moreover, this is satisfied even without the refusal option.
+On the other hand, if the modified URM is adopted instead of the URM, an agent may have an incentive to adopt an OPS. We briefly show this fact by using the following example.
 Example 4
-Let $\mathcal{A=}\left\{a_{1},a_{2}\right\}$ and $\mathcal{O=}\left\{o_{1},o_{2},\varnothing\right\}$.
-Let $\delta=\left(o_{1},o_{2},\varnothing\right)$ and $\delta^{\prime}=\left(o_{1},\varnothing,o_{2}\right)$.
-We assume $q_{o}=1$.
-Suppose $r_{a_{1}}=\delta$ and $p_{a_{2}}=\delta$.
+Let $\mathcal{A=}\left\{a_{1},a_{2},a_{3}\right\}$ and $\mathcal{O=}\left\{o_{1},o_{2},\varnothing\right\}$.
+Let $\delta=\left(o_{1},o_{2},\varnothing\right)$, $\delta^{\prime}=\left(o_{1},\varnothing,o_{2}\right)$ and $\delta^{\prime\prime}=\left(\varnothing,o_{1},o_{2}\right)$.
+We assume $q_{o_{1}}=q_{o_{2}}=1$.
+Suppose $r_{a_{1}}=\delta$, $p_{a_{2}}=\delta$ and $p_{a_{3}}=\delta^{\prime\prime}$.
+Then, $\delta^{\prime}$ is an OPS of $a_{1}$.
 In this case, since $\left(\delta(=r_{a_{1}}),\delta\right)$ does not satisfy all of (i), (ii) and (iii),
-$\left(f^{M}\left(\delta,\delta\right)\right)_{a_{1}o_{1}}$ $=$ $\left(f^{M}\left(\delta,\delta\right)\right)_{a_{1}o_{2}}=\frac{1}{2},$
-$\left(f^{U}\left(\delta,\delta\right)\right)_{a_{1}o_{1}}$ $=$ $\left(f^{U}\left(\delta,\delta\right)\right)_{a_{1}o_{2}}=\frac{1}{2}.$
+$\left(f^{M}\left(\delta,\delta,\delta^{\prime\prime}\right)\right)_{a_{1}o_{1}}$ $=$ $\left(f^{M}\left(\delta,\delta,\delta^{\prime\prime}\right)\right)_{a_{1}o_{2}}=\frac{1}{2},$
+$\left(f^{U}\left(\delta,\delta,\delta^{\prime\prime}\right)\right)_{a_{1}o_{1}}$ $=$ $\left(f^{U}\left(\delta,\delta,\delta^{\prime\prime}\right)\right)_{a_{1}o_{2}}=\frac{1}{2}.$
 On the other hand, since $\left(\delta^{\prime}(\neq r_{a_{1}}),\delta\right)$ satisfies all of (i), (ii) and (iii),
-$\left(f^{M}\left(\delta^{\prime},\delta\right)\right)_{a_{1}o_{1}}$ $=$ 1, ($1,\text{ }\left(f^{U}\left(\delta^{\prime},\delta\right)\right)_{a_{1}o_{2}}=0,$
-$\left(f^{U}\left(\delta^{\prime},\delta\right)\right)_{a_{1}o_{1}}$ $=$ $\left(f^{U}\left(\delta^{\prime},\delta\right)\right)_{a_{1}o_{2}}=\frac{1}{2}.$
+$\left(f^{M}\left(\delta^{\prime},\delta,\delta^{\prime\prime}\right)\right)_{a_{1}o_{1}}$ $=$ $1,\text{ }\left(f^{M}\left(\delta^{\prime},\delta,\delta^{\prime\prime}\right)\right)_{a_{1}o_{2}}=0,$
+$\left(f^{U}\left(\delta^{\prime},\delta,\delta^{\prime\prime}\right)\right)_{a_{1}o_{1}}$ $=$ $\left(f^{U}\left(\delta^{\prime},\delta,\delta^{\prime\prime}\right)\right)_{a_{1}o_{2}}=\frac{1}{2}.$
 Thus, in this case, if $f^{M}$ is employed, then $a_{1}$ has an incentive to choose $\delta^{\prime}$ instead of $r_{a_{1}}=\delta$.
-In other words, in this case, $a_{1}$ has an incentive to reveal by reducing its truly acceptable types.
+In other words, in this case, $a_{1}$ has an incentive to use an OPS.
 On the other hand, if $f^{U}$ is employed, then $a_{1}$ has no such an incentive.
-Of course, this does not mean that truth-telling is strategically dominated revealing some reduction of $r_{a_{1}}$.
-However, agents may have such incentives even if they do not have the refusal option.
+Formally, we have the following result.
+Proposition 7
+Let $f$ be a fair rank-minimizing rule such that, for any $r_{a}\in\mathcal{P}$, there is no $p_{a}\in\mathcal{P}$ that dominates $r_{a}$ with a refusal option.
+Then, there is some $(r_{a},p_{-a})$ such that $a$ strictly prefers $g\left(f\left(\pi_{a},p_{-a}\right),r\right)$ to $g\left(f\left(r_{a},p_{-a}\right),r\right)$ where $\pi_{a}$ is an OPS of $a$.
+Moreover, this is satisfied even without the refusal option; that is, for that $(r_{a},p_{-a})$, $a$ strictly prefers $f\left(\pi_{a},p_{-a}\right)$ to $f\left(r_{a},p_{-a}\right)$.
+Proposition 5 shows that it is possible to construct a fair rank-minimizing mechanism under which truth-telling is not a dominated strategy.
+However, doing so gives rise to a new type of strategic concern that does not exist under the URM.
+Specifically, as shown in Proposition 6, under the URM, any OPS is dominated by truth-telling, thereby eliminating the risk of such manipulation.
+In contrast, Proposition 7 implies that under a fair rank-minimizing mechanism in which truth-telling is not dominated by any strategies, including the modified URM, an OPS is not dominated by truth-telling; that is, for an agent, there exsits some profile of other agents This suggests that manipulation through such strategies may persist even in mechanisms designed to eliminate strategic dominance of truth-telling.
+Of course, this does not mean that truth-telling is dominated by an OPS.
+However, this reveals a trade-off inherent in fair rank-minimizing mechanisms: while modifying them can eliminate the strategic dominance of truth-telling, such modifications may introduce incentives for other manipulations.
 5 Concluding Remarks
-We show that if an assignment is determined by the uniform rank-minimizing mechanism, which is a fair one, and agents can refuse the assignment and obtain the outside option instead, then truth-telling is strategically dominated by an ODS.
+We show that if an assignment is determined by the URM, which is a fair one, and agents can refuse the assignment and obtain the outside option instead, then truth-telling is dominated by an ODS.
 The adoption of such strategies by agents may result in inefficient assignments; that is, an agent who adopts an ODS refuses their assigned type, although some other agents desire that type.
@@
 Therefore, if resolving the problem in that way is difficult, it is necessary to consider alternative mechanisms.
-First, as shown in this study, if we modify the uniform rank-minimizing mechanism, then truth-telling is not strategically dominated by any other strategies, including the outside option demotion ones.
+First, as shown in this study, if we modify the URM, then truth-telling is not dominated by any other strategies, including the outside option demotion ones.
 However, in this modification, it is necessary to ensure that increasing the number of acceptable types can result in a disadvantage.
-Thus, this modification may lead to the opposite problem, where agents strategically decrease the number of acceptable types.
+Thus, this modification leads to the opposite strategic problem, where agents have an incentive to decrease the number of acceptable types.
 Second, we consider efficient and fair mechanisms other than the fair rank-minimizing mechanisms, such as the (general) probabilistic serial mechanism introduced by Bogomolnaia and Moulin (2001) and Budish et al.
@@
 $\sum\limits_{\bar{o}:R(\bar{o},p_{a})\leq\bar{k}\left(p_{a}\right)}\sum\limits_{\bar{a}\in\mathcal{A}}x_{\bar{a}o^{\prime}}$ $<$ $\left|\mathcal{A}\right|\leq\sum\limits_{\bar{o}:R(\bar{o},p_{a})\leq\bar{k}\left(p_{a}\right)}q_{o^{\prime}}\text{, and}$
-$\sum\limits_{\bar{o}:R(\bar{o},p_{a})\leq\bar{k}\left(p_{a}\right)}x_{a\bar{o}}$ $<$ $1$
+$\sum\limits_{\bar{o}:R(\bar{o},p_{a})\leq\bar{k}\left(p_{a}\right)}x_{a\bar{o}}$ $<$ $1.$
 Since $\sum_{\bar{a}\in\mathcal{A}}x_{\bar{a}o^{\prime\prime}}\leq q_{o^{\prime\prime}}$ for all $o^{\prime\prime}\in\mathcal{O}$, there is some $o^{\prime}\in\mathcal{O}$ such that $R(o^{\prime},p_{a})\leq\bar{k}\left(p_{a}\right)$ and $\sum_{\bar{a}\in\mathcal{A}}x_{\bar{a}o^{\prime}}<q_{o^{\prime}}$.
@@
 Now, let
-$Y=\left\{y\in\mathcal{Y}^{\ast}\left(p\right)\left|\text{ }y_{ao}=1\right.\right\},Y^{\prime}=\left\{y^{\ast}\in\mathcal{Y}^{\ast}\left(p\right)\left|\text{ }y_{a^{\prime}o}^{\ast}=1\right.\right\}.$
+[…]
 Since $\left(f^{U}\left(p\right)\right)_{ao}>0$, $Y\neq\emptyset$ and thus let $Y=\left\{y^{1},\cdots,y^{\left|Y\right|}\right\}$.
@@
 Next, we consider $j$th group for each $j=2,3,\cdots,k-1$.
-In $j$th group, there are just $q_{r_{a}^{j}}$ agents whose first-best object type is $r_{a}^{j}$ and the second-best object is $r_{a}^{j+1},$ the third-best $r_{a}^{j+2},$cdots and the $k$th-best $r_{a}^{k}$ under $\left(r_{a},p_{-a}\right)$.
+In $j$th group, there are just $q_{r_{a}^{j}}$ agents whose first-best object type is $r_{a}^{j}$ and the second-best object is $r_{a}^{j+1},$ the third-best $r_{a}^{j+2},$ $\cdots$ and the $k$th-best $r_{a}^{k}$ under $\left(r_{a},p_{-a}\right)$.
 If $a_{j}$ belongs $j$th group for $j=1,2,\cdots,k-1$, the preference orders of them are as follows:
@@
 By Lemma 2, there is $p_{-a}\in\mathcal{P}^{\left|\mathcal{A}\right|-1}$ such that $a$ strictly prefers $f\left(r_{a},p_{-a}\right)$ to $f\left(p_{a},p_{-a}\right)$.
-Therefore, in this case, $r_{a}$ is not strategically dominate by $p_{a}$.
-Second, suppose $k>\bar{k}\left(p_{a}\right)$.
-In this case, since $f$ satisfies ETE, $f\left(r_{a},p_{-a}\right)=f\left(p_{a},p_{-a}\right)$.
-Therefore, $p_{a}$ does not strategically dominate $r_{a}$.
+Therefore, in this case, $r_{a}$ is not dominate by $p_{a}$.
+Second, suppose $k>\bar{k}\left(p_{a}\right)$. In this case, since $f$ satisfies ETE, $f\left(r_{a},p_{-a}\right)=f\left(p_{a},p_{-a}\right)$. Therefore, $p_{a}$ does not dominate $r_{a}$.
 Proof of Theorem 1
@@
 Proof of Proposition 3
-We use the example introduced in the proof of Theorem 2.
-Suppose that there are $a\in\mathcal{A}$, $o,o^{\prime}\in\mathcal{O}$ satisfying (3).
-Let $d_{a}$ be such that $R\left(\varnothing,r_{a}\right)=R\left(o^{\prime},d_{a}\right)$ and $p_{a^{\prime}}=d_{a}$ for all $a^{\prime}\in\mathcal{A\setminus}\left\{a\right\}$.
-Then, since (3) is satisfied and $f^{U}$ satisfies ETE, there is $y\in\mathcal{Y}^{\ast}\left(d_{a},p_{-a}\right)$, $y_{ao^{\prime}}=1$ and $y_{a^{\prime}o^{\prime\prime}}=1$ for some $a^{\prime}\in\mathcal{A\setminus}\left\{a\right\}$ and $o^{\prime\prime}\in\mathcal{O}$ such that $R\left(o^{\prime},d_{a}\right)<R\left(o^{\prime\prime},d_{a}\right)$.
-Then, if $y$ is realized, $a$ refuses $o^{\prime}$ but $a^{\prime}$ prefers $o^{\prime}$ to $o^{\prime\prime}$.
-Therefore, $g\left(f^{U}\left(d_{a},p_{-a}\right),r\right)$ is wasteful.
+Let $r_{a}=\left(o_{1},\varnothing,\cdots\right)$ and $r_{a^{\prime}}=\left(o_{1},o_{2},\varnothing,\cdots\right)$ for all $a^{\prime}\in\mathcal{A\setminus}\left\{a\right\}$.
+Furthermore, let $d=\left(d_{\bar{a}}\right)_{\bar{a}\in\mathcal{A}}$ such that $d_{\bar{a}}=(o_{1},o_{2},o_{3},\cdots,\varnothing)$ for all $\bar{a}\in\mathcal{A}$; that is, $d$ is a profile in which every agent adopts an ODS.
+First, we show that no agent has an incentive to unilaterally deviate from $d$. It is sufficient to show that for all $\bar{a}\in\mathcal{A}$ and all $p_{\bar{a}}\in\mathcal{P}$, either
+$\left(f^{U}\left(d\right)\right)_{\bar{a}o_{1}}>\left(f^{U}\left(p_{\bar{a}},d_{{}_{-\bar{a}}}\right)\right)_{\bar{a}o_{1}}$ (18)
+or
+$\left(f^{U}\left(d\right)\right)_{\bar{a}\bar{o}}=\left(f^{U}\left(p_{\bar{a}},d_{{}_{-\bar{a}}}\right)\right)_{\bar{a}\bar{o}},\,\ \text{for all }\bar{o}\in\mathcal{O},$ (19)
+because $o$ is the best object type for all agents.
+Let $k$ be the smallest integer such that $p_{\bar{a}k}\neq d_{\bar{a}k}$.
+By Lemma 2, if $k>\bar{k}\left(p_{\bar{a}}\right)$, then (19) is satisfied.
+We assume $k\leq\bar{k}\left(p_{{}_{\bar{a}}}\right)$; that is,
+$\sum\limits_{\bar{o}:R(\bar{o},p_{\bar{a}})<k}q_{\bar{o}}\left(=\sum\limits_{\bar{o}:R(\bar{o},d_{\bar{a}})<k}q_{\bar{o}}\right)<\left|\mathcal{A}\right|$
+is satisfied for that $k$.
+Let $o^{\prime\prime}$ such that $R\left(o^{\prime\prime},p_{\bar{a}}\right)=k<R\left(o^{\prime\prime},d_{\bar{a}}\right)$.
+Let $y\in\mathcal{Y}^{\ast}\left(d\right)$ such that $y_{\bar{a}o_{k}}=1$.
+Moreover, we arbitrarily choose $y^{\prime}$ such that $y_{\bar{a}o_{1}}^{\prime}=1$.
+We consider two cases. First, we assume
+$\sum\limits_{\bar{o}:R(\bar{o},d_{\bar{a}})\leq k}q_{\bar{o}}=\sum\limits_{\bar{o}:R(\bar{o},d_{\bar{a}})<k}q_{\bar{o}}+q_{o_{k}}<\left|\mathcal{A}\right|.$
+In this case, some agent is assigned to $o_{k+1}$ in a rank-minimizing assignment under $d$.
+Let $y^{\prime\prime}$ be such that $y_{\bar{a}o^{\prime\prime}}^{\prime\prime}=1$, $y_{a^{\prime}o_{k}}^{\prime\prime}=1$ for $a^{\prime}$ with $y_{a^{\prime}o_{k+1}}=1,$ and $y_{a^{\prime\prime}\bar{o}}^{\prime\prime}=y_{a^{\prime\prime}\bar{o}}$ for all $a^{\prime\prime}\in\mathcal{A\setminus}\left\{\bar{a},a^{\prime}\right\}$ and $\bar{o}\in\mathcal{O}$.
+Then,
+$RV(y^{\prime\prime},\left(p_{\bar{a}},d_{-\bar{a}}\right))<RV(y,d)\leq RV(y^{\prime},d)\leq RV(y^{\prime},\left(p_{\bar{a}},d_{-\bar{a}}\right)).$
+Thus, $y^{\prime}\notin\mathcal{Y}^{\ast}\left(p_{\bar{a}},d_{-\bar{a}}\right)$ and (18) is satisfied in this case.
+Second, we assume
+$\sum\limits_{\bar{o}:R(\bar{o},d_{\bar{a}})\leq k}q_{\bar{o}}=\sum\limits_{\bar{o}:R(\bar{o},d_{\bar{a}})<k}q_{\bar{o}}+q_{o_{k}}\geq\left|\mathcal{A}\right|.$
+Then, since $q_{o_{1}}<\left|\mathcal{A}\right|$, $k>1$.
+Let $y^{\prime\prime}$ be such that $y_{\bar{a}o^{\prime\prime}}^{\prime\prime}=1$, and $y_{a^{\prime\prime}\bar{o}}^{\prime\prime}=y_{a^{\prime\prime}\bar{o}}$ for all $a^{\prime\prime}\in\mathcal{A\setminus}\left\{\bar{a}\right\}$ and $\bar{o}\in\mathcal{O}$.
+Moreover, let $y^{\prime\prime\prime}$ be an arbitrary assignment in $\mathcal{Y}^{\ast}\left(d\right)$ such that $y_{\bar{a}o_{k^{\prime}}}^{\prime\prime\prime}=1$ for some $k^{\prime}=1,\cdots,k-1$.
+$RV(y^{\prime\prime},\left(p_{\bar{a}},d_{-\bar{a}}\right))=RV(y,d)=RV(y^{\prime\prime\prime},d)=RV(y^{\prime\prime\prime},\left(p_{\bar{a}},d_{-\bar{a}}\right)).$
+Therefore, (19) is satisfied in this case.
+Finally, suppose $q_{o_{1}}+q_{o_{2}}<\left|\mathcal{A}\right|$.
+We show that $g\left(f^{U}\left(d\right),r\right)$ is wasteful.
+Since $q_{o_{1}}+q_{o_{2}}<\left|\mathcal{A}\right|$, there is $y\in\mathcal{Y}^{\ast}\left(d\right)$, $y_{ao_{2}}=1$ and $y_{a^{\prime}o_{1}}=y_{a^{\prime}o_{2}}=0$ for some $a^{\prime}\in\mathcal{A\setminus}\left\{a\right\}$.
+Thus, $\left(f^{U}\left(d\right)\right)_{ao_{2}}>0$ and $\left(g\left(f^{U}\left(d\right),r\right)\right)_{ao_{2}}=0$.
+Therefore, $g\left(f^{U}\left(d\right),r\right)$ is wasteful.
 Proof of Proposition 4
@@
 First, if $r_{a}^{1}=\varnothing,$ then $\left(f^{M}\left(r_{a},p_{-a}\right)\right)_{a\varnothing}=1$ for all $p_{-a}\in\mathcal{P}^{\left|\mathcal{A}\right|-1}$.
-Thus, in this case, there is no $p_{a}\in\mathcal{P}$ that strategically dominate $r_{a}$.
+Thus, in this case, there is no $p_{a}\in\mathcal{P}$ that dominate $r_{a}$.
 Thus, we assume $r_{a}^{1}\neq\varnothing$.
@@
 for all $p_{-a}\in\mathcal{P}^{\left|\mathcal{A}\right|-1}$.
-Therefore, for any $r_{a}\in\mathcal{P}$, there is no $p_{a}\in\mathcal{P}$ that strategically dominate $r_{a}$.
+Therefore, for any $r_{a}\in\mathcal{P}$, there is no $p_{a}\in\mathcal{P}$ that dominates $r_{a}$.
+Proof of Proposition 6
+Suppose not; that is, there exist $(r_{a},p_{-a})\in\mathcal{P}^{\left|\mathcal{A}\right|}$ such that $a$ strictly prefers $g\left(f^{U}\left(\pi_{a},p_{-a}\right),r\right)$ to $g\left(f^{U}\left(r_{a},p_{-a}\right),r\right)$ where $\pi_{a}$ is an OPS of $a$.
+By the definition of OPS, when $o$ and $o^{\prime}$ satisfy $R\left(o,r_{a}\right)<R\left(o^{\prime},r_{a}\right)<R\left(\varnothing,r_{a}\right)$ and $R\left(o^{\prime},\pi_{a}\right)>R\left(\varnothing,\pi_{a}\right),$
+$RV\left(y,\left(\pi_{a},p_{-a}\right)\right)$ $=$ $RV\left(y,\left(r_{a},p_{-a}\right)\right)\text{ if }y_{ao}=1,$ (20)
+$RV\left(y,\left(\pi_{a},p_{-a}\right)\right)$ $<$ $RV\left(y,\left(r_{a},p_{-a}\right)\right)\text{ if }y_{a\varnothing}=1,$ (21)
+$RV\left(y,\left(\pi_{a},p_{-a}\right)\right)$ $>$ $RV\left(y,\left(r_{a},p_{-a}\right)\right)\text{ if }y_{ao^{\prime}}=1.$ (22)
+First, we show that $g\left(f^{U}\left(r_{a},p_{-a}\right),r\right)_{a\varnothing}=0$.
+Suppose not; that is, $y\in\mathcal{Y}^{\ast}\left(r_{a},p_{-a}\right)$ such that $y_{ao}=1$ for some $o$ satisfying $R\left(\varnothing,r_{a}\right)\leq R\left(o,r_{a}\right)$.
+By Lemma 2, $y_{a\varnothing}=1$.
+Then, by (20), (21) and (22), for any $y^{\prime}\in\mathcal{Y}^{\ast}\left(\pi_{a},p_{-a}\right)$, $y_{a\varnothing}^{\prime}=1$.
+However, this contradicts that $a$ strictly prefers $g\left(f^{U}\left(\pi_{a},p_{-a}\right),r\right)$ to $g\left(f^{U}\left(r_{a},p_{-a}\right),r\right)$.
+Therefore, $g\left(f^{U}\left(r_{a},p_{-a}\right),r\right)_{a\varnothing}=0$; that is, for any $y\in\mathcal{Y}^{\ast}\left(r_{a},p_{-a}\right)$, $y_{a\varnothing}=0$.
+Then, for any $y\in\mathcal{Y}^{\ast}\left(r_{a},p_{-a}\right)$, $y_{ao}=1$ implies $R\left(\varnothing,r_{a}\right)>R\left(o,r_{a}\right)$.
+Since $a$ strictly prefers $g\left(f^{U}\left(\pi_{a},p_{-a}\right),r\right)$ to $g\left(f^{U}\left(r_{a},p_{-a}\right),r\right)$, for any $y^{\prime}\in\mathcal{Y}^{\ast}\left(\pi_{a},p_{-a}\right)$, $y_{a\varnothing}^{\prime}=0$ is also satisfied.
+Next, we show that for any $y\in\mathcal{Y}^{\ast}\left(r_{a},p_{-a}\right)$, $y_{ao}=1$ implies $R\left(o,\pi_{a}\right)<R\left(\varnothing,\pi_{a}\right)$.
+Suppose not; that is, $y_{ao}=1$ implies $R\left(o,\pi_{a}\right)>R\left(\varnothing,\pi_{a}\right)$.
+Then, let $y^{\prime}$ be such that $y_{a\varnothing}^{\prime}=1$ and $y_{\bar{a}\bar{o}}^{\prime}=y_{\bar{a}\bar{o}}$ for any other $\left(\bar{a},\bar{o}\right).$ By (20), (21) and (22),
+$RV\left(y^{\prime},\left(\pi_{a},p_{-a}\right)\right)$ $<$ $RV\left(y,\left(r_{a},p_{-a}\right)\right)\leq$
+$RV\left(y^{\prime\prime},\left(r_{a},p_{-a}\right)\right)$ $\leq$ $RV\left(y^{\prime\prime},\left(\pi_{a},p_{-a}\right)\right),$
+for any $y^{\prime\prime}$ such that $y_{a\varnothing}^{\prime\prime}=0$, because $y\in\mathcal{Y}^{\ast}\left(r_{a},p_{-a}\right)$.
+This contradicts that for any $y^{\prime}\in\mathcal{Y}^{\ast}\left(\pi_{a},p_{-a}\right)$, $y_{a\varnothing}^{\prime}=0$.
+Thus, for any $y\in\mathcal{Y}^{\ast}\left(r_{a},p_{-a}\right)$, $y_{ao}=1$ implies $R\left(o,\pi_{a}\right)<R\left(\varnothing,\pi_{a}\right)$.
+We arbitrarily choose $y\in\mathcal{Y}^{\ast}\left(r_{a},p_{-a}\right)$. By (20),
+$RV\left(y,\left(\pi_{a},p_{-a}\right)\right)$ $=$ $RV\left(y,\left(r_{a},p_{-a}\right)\right)\leq$
+$RV\left(y^{\prime\prime},\left(r_{a},p_{-a}\right)\right)$ $\leq$ $RV\left(y^{\prime\prime},\left(\pi_{a},p_{-a}\right)\right),$
+for any $y^{\prime\prime}$ such that $y_{a\varnothing}^{\prime\prime}=0$.
+Since $y_{a\varnothing}^{\prime}=0$ for any $y^{\prime}\in\mathcal{Y}^{\ast}\left(\pi_{a},p_{-a}\right)$, $\mathcal{Y}^{\ast}\left(r_{a},p_{-a}\right)=\mathcal{Y}^{\ast}\left(\pi_{a},p_{-a}\right)$ but this contradicts that $a$ strictly prefers $g\left(f^{U}\left(\pi_{a},p_{-a}\right),r\right)$ to $g\left(f^{U}\left(r_{a},p_{-a}\right),r\right)$.
+The proof above is satisfied even without the refusal option, because $g\left(f\left(\cdot,\cdot\right),r\right)=f\left(\cdot,\cdot\right)$ is satisfied for all $g\left(f\left(\cdot,\cdot\right),r\right)$ above.
+Proof of Proposition 7
+Let $f$ be a fair rank-minimizing rule such that, for any $r_{a}\in\mathcal{P}$, there is no $p_{a}\in\mathcal{P}$ that dominates $r_{a}$ with the refusal option.
+First, we consider the following specific example.
+Let $a_{1},a_{2},a_{3}\in\mathcal{A}$ and $o_{1},o_{2}\in\mathcal{O}$.
+Suppose $q_{o_{1}}=q_{o_{2}}=1$.
+Let $\delta=\left(o_{1},\varnothing,\cdots\right)$ and $\delta^{\prime}=\left(o_{1},o_{2},\cdots,\varnothing\right)$.
+First, we assume $r_{a_{1}}=p_{a_{2}}=\delta$, $R\left(o_{2},p_{a_{3}}\right)=1,$ and $R\left(\varnothing,p_{a}\right)$ for any $a\in\mathcal{A\setminus}\left\{a_{1},a_{2},a_{3}\right\}$ if there exists some.
+Then, $\delta^{\prime}$ is an ODS of $a,$ and
+$\left(f\left(\delta,p_{-a_{1}}\right)\right)_{a_{1}o_{1}}$ $=$ $\left(f\left(\delta,p_{-a_{1}}\right)\right)_{a_{1}\varnothing}=\frac{1}{2}\text{,}$
+$\left(f\left(\delta^{\prime},p_{-a_{1}}\right)\right)_{a_{1}o_{1}}$ $=$ $1$
+Therefore, $a_{1}$ strictly prefers $g\left(f\left(\delta^{\prime},p_{-a_{1}}\right),r\right)$ to $g\left(f\left(\delta,p_{-a_{1}}\right),r\right)$.
+Since $\delta^{\prime}$ does not dominate $r_{a_{1}}$ under $f$, there is some $p_{-a_{1}}^{\prime}\in\mathcal{P}^{\left|\mathcal{A}\right|-1}$ such that $a_{1}$ strictly prefers $g\left(f\left(\delta,p_{-a_{1}}^{\prime}\right),r\right)$ to $g\left(f\left(\delta^{\prime},p_{-a}^{\prime}\right),r\right)$.
+Next, we assume that $r_{a_{1}}=\delta^{\prime}$.
+Then, $\delta$ is an OPS of $a_{1}$.
+Therefore, $a_{1}$ strictly prefers $g\left(f\left(\pi_{a},p_{-a_{1}}^{\prime}\right),r\right)$ to $g\left(f\left(r_{a_{1}},p_{-a}^{\prime}\right),r\right)$, where $r_{a_{1}}=\delta^{\prime}$ and $\pi_{a}=\delta$.
+The proof above is satisfied even without the refusal option, because $g\left(f\left(\cdot,\cdot\right),r\right)=f\left(\cdot,\cdot\right)$ is satisfied for all $g\left(f\left(\cdot,\cdot\right),r\right)$ above.
+Declaration of generative AI and AI-assisted technologies in the writing process
+During the preparation of this work the author used ChatGPT (OpenAI) in order to improve readability and language.
+After using this, the author reviewed and edited the content as needed and takes full responsibility for the content of the publication.